A Complete Bibliography of Publications in Science in Context

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

02 January 2024 Version 1.32

Title word cross-reference

1 [947]. 1840s [1011]. 1860s [915, 779]. 1870s [696]. 1884-1894 [887]. 1910s [696, 1042]. 1920s [944, 501, 585]. 1930 [318]. 1930s [944, 779]. 1949 [820]. 1950s [1003, 898]. 1960s [518]. 1970s [841, 723]. 1977 [1062]. 1990s [1003]. 1992 [1063].

24.und [1062]. 24th [1062]. 25th [1062].

3 [723]. 32 [484].

400th [1063].

8 [524].

Aberration [562]. Abiding [703]. above [491]. Abraham [284]. Abstract [580]. Abstraction [491, 748, 498]. Abused [135]. Aby [362]. academia [982]. Academic [242, 761, 561]. Academy [670, 873, 64, 65, 965]. Accident [609]. Accidents [536]. Acclimatization [903]. Accommodating [763]. Accommodation [903]. Account [822, 15, 1023]. Accountability [159]. Accounting [168, 161, 164]. Accounts [159, 729]. acculturation [1016]. Accuracy [27, 931]. Acquiring [965]. across [729]. Act [642]. Activity [820]. Acts [242]. Actuaries [162]. Adaptability [569]. Addendum [976]. Adding [584]. Administrative [809, 614]. Adrian [838]. Advanced [181]. Advent [850]. adventu [596]. Adventures [439]. Advertising [619]. Advice [345]. Advising [675]. Advocating [882]. Aesthetic [544, 931]. Aestheticization [515]. Aesthetics [930, 193]. Affair [670, 691]. Affaire [388, 406]. after [266, 457, 984]. Afterword [194]. Against [300, 244, 720, 903, 566]. Age [517, 183, 764, 258, 940]. Ageing [679]. Agent [1045, 216, 948]. Agent-Based [1045]. Agents [626]. Ages [174, 222, 427, 596]. Agricultural [454]. AIDS [621]. Aimé [652]. Airy [1035]. Aksakov [575]. al [424]. al-Isfizari [424]. Albert [240, 689]. Alberto [605]. Albrecht [625]. alchemical [973]. Alcohol [821]. Aleksandr [496]. Alexandre [787, 200]. Alfred [739]. Algebra [909, 879, 990]. Algebras [420]. Algorithmics [215]. Alienism [631]. Alive [316]. Alpine [653]. Alps [654, 649, 653]. Alternative [225, 189, 663]. Altitude [656]. Am [792, 1062]. America [565, 536]. American [821, 706, 528, 676, 518, 370, 179, 88, 922, 104, 996, 331, 191, 675, 729, 317]. Americans [85]. Americas [993]. Among [279, 93, 162]. Analogies [720]. Analogy [488]. Analyses [563]. Analysis [716, 820, 502, 668, 1023]. Analytic [429]. Analytical [914]. Analyzing [107]. Anarchy [805]. Anatomy [1036, 640, 873]. Anchoring [829]. Ancient [490, 791, 525, 664, 491, 526, 635]. Andronov [504]. Andy [73]. Animal [576, 814, 874, 872, 875, 628]. Animals [705, 873]. Animism [1019]. anniversary [1025]. Annular [801]. anomaly [984]. Anthrax [616]. Anthropoid [456]. Anthropology [671, 252, 199, 958]. anthropometric [955]. Anti [617, 118]. Anti-Einstein [118]. Anti-Tuberculosis [617]. Antibiotic [810]. Antimicrobial [950]. Antiquity [910]. Antisociality [805]. Antitoxin [614]. Antonio [393, 411]. Anxiety [498, 874]. Any [204]. Anything [296]. Apes [456]. Apocryphal [394, 412]. Apodictic [47]. Apollonius [674, 488]. Apostolic [46]. Appendix [229, 397, 415, 593, 390, 408, 575]. Apperception [837]. Application [308, 485]. Applications [895, 922]. Applied [153]. Apprenticeship [963, 965]. Approach [661, 94, 601]. Approaches [674, 348, 663, 778, 812]. Apsidal [902]. ARA [955]. Arabic [427, 424, 419, 425, 418, 426]. Arbitage [167]. ARCAD [723]. archaeology [1003]. Architect [602]. Architecture [636]. Architectures [373]. Archive [242, 974]. Archytas [521]. Arctic [1000, 812]. Areas [461, 485, 989]. Argentina [623, 622, 532, 566, 533]. Argentine [642]. Argues [892]. Argument [722, 671, 680, 12]. arguments [994, 1028]. Aristotelian [424, 524, 523]. Aristotle [846, 947]. Arnold [716]. Arrest [724, 725]. Art [706, 518, 512, 73, 742, 383, 401, 259, 516, 134, 81].

Article [84]. Artificial [161, 704]. Artist [382, 400]. Artists [517]. Asajiro [683]. ash [1037]. Ashkenazi [280, 584]. Ashkenazic [281]. Aspects [308, 197, 428, 310]. Aspirations [922]. Assessing [627]. Astrology [759, 800, 596]. Astronom [1058]. Astronomer [1058, 802]. Astronomers [908, 26]. Astronomia [329]. Astronomical [273, 752, 771, 423, 540, 636]. Astronomy [759, 801, 902, 668]. Asúa [535]. Asylums [603]. Asymmetry [922]. Atlantic [812]. Atomic [941, 517]. Attachment [974, 874]. Attitude [280, 1026]. Attitudes [841, 518, 37]. Audience [898]. August [237, 765]. Augustin [562]. Augustus [450]. Australian [564]. Austrian [654]. Authenticity [931]. Author [208]. Authority [706, 604, 148, 927]. Authors [780]. authorship [963]. Autism [699]. Autobiographical [794]. Autograph [562]. Autonomy [621, 579, 533]. Avant [192]. Avant-Garde [192]. Avicenna [1019]. Avoid [314].

Back [10, 21, 32, 43, 59, 68, 78, 91, 99, 111, 129, 139, 150, 158, 170, 186, 196, 202, 210, 218, 231, 239, 248, 261, 268, 276, 286, 293, 303, 326, 338, 347, 354, 366, 376, 381, 399, 646, 658, 666, 673, 678, 685, 695, 701, 711, 719, 727, 735, 741, 751, 757, 768, 777, 786, 799, 807, 817, 826, 835, 845, 854, 863, 871, 878, 884, 890, 901, 907, 912, 919, 924, 933, 938, 946, 953, 961, 970, 978, 986, 992, 999, 1007, 1018]. Back [1030, 1040, 1047, 1056, 620, 633]. Bacteriology [369]. Bad [720, 470]. Bakhtinian [148]. Baldi [352]. Balloons [608]. Bank [829]. Baptiste [288, 246]. Barbarism [566]. barriers [974]. Barthez [626]. Based [1045]. Basis [360]. Battle [395, 413]. Battles [720]. BCG [617]. Be [181, 450, 477, 607, 477]. bearing [725]. Became [650, 193]. Becoming [474, 1014]. Before [577]. Beginnings [503]. Behavior [821, 716]. Behavioral [739]. behind [499]. Being [1009]. Belgian [620, 633]. Belgium [1041, 949, 1001]. Belief [447, 594]. Belsen [568]. Belsenitis [568]. ben [273, 284]. Benedetti [12]. Beneficence [335]. Benefiting [618]. Bentley [33]. Ber [590]. Bergson [921]. Berlin [503, 374]. Bernardino [352]. Bestehens [1062]. Between [178, 89, 335, 22, 259, 1016, 754, 253, 965, 848, 944, 545, 237, 869, 540, 859]. Beyond [664, 343, 526, 263, 183, 456, 980]. Biblical [317]. Bibliography [213, 8]. Big [455]. Billy [518]. Binary [544]. Bio [213, 8]. Bio-Bibliography [213, 8]. Biobricks [764]. Biochemistry [811]. Biographies [479]. Biography [508]. Biological [184, 499, 61, 335, 720]. Biologists [789]. Biology [305, 661, 547, 322, 204, 789, 206, 359, 579, 183, 324, 775, 317]. Biomathematics [133]. Biomechanical [738]. Biomedical [809]. Birds [850]. Birkenmajer [935]. Birth [838]. Birthday [1063]. bis [1063]. Bishopric [800]. Black [117, 513]. Blanc [656]. Blanks [513]. Blind [385, 403]. Blood [971]. bloodstain [1023]. Bloor [290]. Blue [887]. Bodies [814, 117, 638, 222, 859]. Body [204, 520]. Bohr [121]. Bois [988, 987]. Bois-Reymond [988, 987]. Bones [532, 641]. Book [206, 488]. Bookshelf [271]. Borderline [556]. Borders [563, 567]. Bordeu [627, 626]. Borel [725, 724]. Boris [1024]. Borkenau [8]. Borochov [590]. Boscovich [920]. Botanical [707]. Botanizing [881]. Botany [492]. Boulding [676]. Boundaries [348, 858, 612]. Boundary [312, 639, 823, 859]. Bounds [298]. Bourbaki [264, 263]. Bourbakism [153]. Boussinesq [867]. Box [513]. Boyle [447, 429, 1019, 143, 23]. Boys [470]. Brahe [759]. Brain [836, 442, 838, 842, 265, 443, 747, 1004]. Brains [334, 479]. Brazilian [530, 996, 531]. Breaking [902]. Breeding [456]. Bright [539]. Britain [780, 378, 34]. British [33, 805]. Brothers [754, 832]. Brouwer [301]. Bruhl [200]. Bruno [1002]. Bryson [1044]. Budgets [161]. Builders [24]. Building [814, 779, 957]. Burden [308]. Bureau [675]. Bush [577].

c [975, 1051, 972, 437]. C. [437]. ca [814, 705, 812, 811, 619, 176]. Caesar [753]. Calculating [162, 1059]. Calculation [418, 164]. Calculator [1061, 1057, 1059, 1058]. Calculators [682, 1061, 13, 1060]. Calculus [852, 899]. Called [564]. Calling [448]. came [1004]. Campaign [118, 955]. Camps [569]. Can [314, 629, 904]. Cancer [181, 1000]. Cantor [499]. Capitalism [683]. Capitalization [193]. Carapezza [605]. Care [813, 181, 859]. Carinthia [419]. Carl [126]. Cartographical [696]. Cartography [651]. CAS [65]. Case [610, 124, 467, 609, 282, 141, 244, 119, 177, 571, 631, 637, 133, 82, 942, 575, 216, 801, 818, 430, 680, 176, 297, 848, 982, 1022, 1041, 1001]. Cases [332]. Caspar [632]. Cassirer [356, 357, 227, 358, 363, 359, 360, 361]. Casti [318]. Castro [283]. Category [80]. Catholic [48, 52]. Catholicism [318]. Catoptrica [752]. Cats [875]. Causality [124]. Causation [691, 948]. cause [1005]. Caverni [391, 409]. Ceará [530]. ceiling [982]. Celebration [55]. Celebrity [478]. Cell [715, 579]. Cells [371]. Censorship [880, 929]. Center [492]. Central [30]. Centuries [729, 679, 423]. Century [453, 882, 564, 856, 566, 910, 778, 192, 964, 976, 988, 990, 1041, 1010, 4, 819, 703, 425, 836, 262, 681, 731, 233, 529, 474, 552, 61, 697, 513, 625, 698, 611, 498, 436, 479, 512, 571, 476, 917, 629, 603, 715, 602, 377, 608, 597, 216, 259, 451, 132, 532, 372, 280, 579, 190, 134, 374, 782, 873, 331, 76, 651, 675, 278, 47, 653, 607]. Cerebral [479]. Chain [575]. Chains [573, 574]. Challenging [135]. Chamalières [288]. Chance [536]. Change [763, 94, 197, 979, 1044]. Changes [233, 698]. Changing [1002, 928]. Chaos [644]. Character [539]. Charge [314]. Charlatan [1010, 1009]. charlatans [1012]. Charles [630, 478, 578, 921]. Chattel [335]. cheap [1037]. Chemical [233, 429, 151, 235, 234]. Chemist [291]. Chemist-Philosopher [291]. Chemistry [681, 237, 221, 429, 379, 134]. Chengdu [903]. Chicken [560]. Children [770, 967]. China [664, 262, 491, 526, 733, 618, 914, 538, 350, 903, 913]. Chinese [915, 990, 349, 520, 917, 916, 64, 65]. Choice [345]. Christi [596]. Christian [273, 543, 680]. Christopher [465, 463]. Chronology [273]. Church [387, 405, 45]. Científica [494]. Cigoli [386, 404]. Cinema [839, 744, 713, 553, 712]. Cinematic [715]. Cinematographs [921]. Circa [39]. Circle [1035]. Circles [117]. circulation [1013]. Circumpolar [812]. Citation [980]. Citizen [761]. City [897]. Civiale [652]. Civil [455, 804]. Clarity [83]. Clarke [948]. Class [755, 477]. Classic [609]. Classical [511, 592, 382, 400, 540]. Classification [846, 755, 640]. classroom [1052]. Clavius [463]. Clergymen [703]. Clever [359]. Clifford [578]. Climate [763, 644, 1044]. Clinical [610]. Coast [199]. Code [547]. Coefficients [909]. Cognition [934, 39]. Cognitive [94, 106, 199]. Coherence [425]. Cold [93, 637, 996]. Collaboration [881]. Collaborations [951]. Collaborators [505]. Collared [850]. Colleagues [115, 724]. Collecting [707, 1051, 537]. Collection [420, 1054]. collections [1036]. Collective [733, 200, 733]. collectives [1024]. Colleges [70]. Collin [581]. Collision [265]. Colombia [888]. Colonial [1054, 881, 567]. Colorimetry [241]. columns [956]. come [1009]. Comenius [137]. Comes [153]. Comisión [494]. Comments [66, 362]. Commercializing [618]. Commission [820]. Committee [241, 639]. Common [576, 917]. Communal [236]. Communicating [778]. Communication [538, 663, 782, 1015]. Communicative [783]. Communist [733]. Communities [860]. Community [308, 779]. Comparative [417, 728, 266, 527, 125, 640, 873, 548, 995]. Compasses [934]. Competition [93]. Competitors [779]. Complementarity [121]. complementary [974]. Complex [71]. Complexity [541, 215, 307]. Composite [576]. Compound [151]. Computer [828, 342]. Computing [644]. Comrade [1026]. Concentration [981]. Concepción [730]. Concept [626, 223, 29, 432, 151, 125, 1034]. Conception [272, 357]. conceptions [1002]. Concepts [627, 233, 234]. Conceptual [698]. Concern [950]. Concerning [574]. Concerns [651]. Conditionals [635]. Conditions [14]. Condorcet [581]. Cones [374]. Conference [820]. Configurations [501, 823]. Configuring [927]. Conflicting [841]. Confluence [264]. Conformist [508]. Confronting [984]. Congress [116, 693]. Conica [488]. Connected [573, 575]. Connecting [164]. Connection [574]. Connector [264]. connubii [318]. Consciousness [363, 987]. Considerations [902, 311]. Consistent [829]. Constantinople [636]. Constants [97]. Constitution [833, 532]. Constraints [141]. Construction [609, 833, 175, 241, 554]. Consumerism [608]. contact [1034]. Containers [858]. Contemporary [849, 748, 182, 320, 775, 1026, 979]. Contending [842]. Contest [706, 1012]. Context [424, 58, 59, 800, 803, 935, 236, 571, 73, 22, 318, 273, 434, 87, 721, 422, 382, 400, 53, 13, 367, 787]. Contexts [880, 528, 489, 1023]. Contextual [644]. Contextures [71]. Contingencies [329]. Continual [662]. Continuities [235]. Continuity [540]. Contra [123, 287]. contribute [967]. Contributions [935]. Control [370]. Controlling [616]. Controversial [1013]. Controversies [294]. Controversy [841, 621, 300, 177, 865, 390, 408, 298, 297, 972, 295]. Convention [243, 936]. Converge [356]. Conversion [119]. Converso [283]. Convexity [638]. Coordinations [749]. Copernican [395, 413].

Copernicanism [12]. Copernicus [800, 880, 422, 428, 12, 935]. Coral [764]. Corporate [721]. Correspondence [697, 56, 956]. Corrigendum [824, 633, 664]. Cortex [438]. Cosmology [800, 662, 540]. Costs [166]. Couch [378, 875]. Counseling [861, 314, 311]. Counter [387, 405]. Counter-Reformation [387, 405]. Counting [709, 652]. Counts [333]. Course [622, 623, 959]. Courses [852, 899]. Courtier [382, 400]. Courtroom [328]. Covariance [361]. Cover [9, 10, 20, 21, 31, 32, 42, 43, 58, 59, 67, 68, 77, 78, 90, 91, 98, 99, 110, 111, 128, 129, 138, 139, 149, 150, 157, 158, 169, 170, 185, 186, 195, 196, 201, 202, 209, 210, 217, 218, 230, 231, 238, 239, 247, 248, 260, 261, 267, 268, 275, 276, 285, 286, 292, 293, 302, 303, 325, 326, 337, 338, 346, 347, 353, 354, 365, 366, 375, 376, 380, 381, 398, 399, 645, 646, 657, 658, 665, 666, 672, 673, 677, 678]. Cover [684, 685, 694, 695, 700, 701, 710, 711, 718, 719, 726, 727, 734, 735, 740, 741, 750, 751, 756, 757, 767, 768, 776, 777, 785, 786, 798, 799, 806, 807, 816, 817, 825, 826, 834, 835, 844, 845, 853, 854, 862, 863, 870, 871, 877, 878, 883, 884, 889, 890, 900, 901, 906, 907, 911, 912, 918, 919, 923, 924, 932, 933, 937, 938, 945, 946, 952, 953, 960, 961, 969, 970, 977, 978, 985, 986, 991, 992, 998, 999, 1006, 1007, 1017, 1018]. Cover [1029, 1030, 1039, 1040, 1046, 1047, 1055, 1056]. Cowles [820]. Cox [637]. craft [963]. Crafting [537]. Created [75]. Creating [616]. Creative [543]. Credibility [721, 823]. Creed [765]. Creeping [715]. Cremona [697]. Crisis [148, 904, 732, 984, 1044]. Critical [168]. Criticism [250]. Critics [251]. Critique [558, 249, 256, 251, 257, 66, 177, 255]. Crocheted [764]. Cross [349, 350, 456]. Cross-Breeding [456]. Cross-Cultural [349, 350]. Crossing [348, 1053]. crusades [1053]. Crystallization [152]. Cuba [529]. Cubism [512]. Cultivated [947]. Cultivating [438, 377]. Cultural [264, 349, 563, 671, 350, 612, 782, 222, 188, 130, 344]. Culturally [178]. Culture [783, 453, 915, 189, 770, 11, 479, 82, 886, 530, 642, 183, 500, 193, 188, 1016, 968]. Cultures [191, 829]. Cumulative [208]. Cundurango [1013]. Cunha [531]. Cure [1013]. Curiosity [968]. curious [1034]. Current [935]. Customers [166]. Cybernetics [458]. Cyprus [569].

Dagognet [213]. Dance [915]. Danish [809]. D’Arcy [803]. Dark [117]. Darknes [145]. Darwin [478, 578, 730]. Darwinian [642, 766, 765]. Darwinism [133, 720, 324]. Data [981, 1048, 1054, 994, 1028, 1051, 1050]. David [282]. Davke [588]. Day [137]. Days [943]. d’Azyr [640]. Dead [1034, 908]. Deal [709]. death [1037, 1025]. deaths [1038]. Debate [699, 349, 867, 879, 561]. Debates [935]. Decade [690]. December [318]. Decision [827, 830]. Decision-Making [830]. decisions [1042]. Decline [904]. Deconstruction [144]. defective [972]. Defense [733, 55]. Deficit [931]. Defining [518, 284]. Definitional [671]. Definitions [832]. Deist [283]. Déjà [545]. del [494]. Demarcation [852, 899]. demise [1033]. Democracy [462]. Democrat [240]. Democratic [539, 1011]. demonstrate [648]. Denmark [973]. Department [37, 761]. Depths [837]. Descartes [51, 864]. Desecularization [451]. Desired [713]. Detainees [569]. Determination [864]. Development [800, 814, 125, 164, 993]. Developmental [506, 974]. Developments [504]. deviance [1052]. Devices [772, 532, 921]. d’Histoire [602]. Diagrams [774, 989]. Dialectic [297]. Dialogue [22, 796, 297]. Dictatorship [997]. Dictionary [832]. did [384, 402]. Didactic [429]. Diego [535]. Difference [256, 493]. Different [237, 638]. Diffraction [943]. Digby [576]. Dilemma [490, 448]. Dilemmas [699]. Dimensionless [97]. Dimensions [981]. Diminution [896]. Dingler [40]. Diphtheria [614, 612]. Diplomatic [880]. directions [984]. disadvantages [972]. Disaster [728]. Disciple [125]. Disciplinary [92]. Discipline [221, 52, 26, 847]. Disciplined [847]. Disciplines [93, 92]. Disciplining [921]. Discontinuities [235]. Discounting [869]. Discourse [759, 783, 100, 256, 49, 142, 177, 690, 428, 154, 102, 940, 176]. Discourses [928, 957]. Discover [384, 402]. Discovered [39]. Discovery [389, 407, 609, 106, 70, 943, 396, 414]. Discovery/Construction [609]. Discrimination [639]. Disease [914]. Diseases [308, 914]. Dispassionate [813]. Dispatches [764]. Dispersal [516]. Display [926]. Dispositif [782]. Disputes [296]. Dissecting [324]. Dissemination [581]. Dissolving [830]. Disturbing [789]. Diva [1014]. Divergences [949]. diverse [788]. Divine [682]. Division [722]. Divorce [332]. DNA [859]. Do [1032, 296, 793]. Doctor [95]. Doctrines [627]. Document [582]. Documentaries [927]. Documentary [928]. Documents [464, 465, 725]. Does [723, 724]. Dogmas [387, 405, 789]. domestic [1011]. Domesticated [253]. Don [690]. Done [181]. Door [705]. doorstep [1048]. doorsteps [1053]. Dove [850]. Draftsman [385, 403]. Dramatic [225]. Drawing [769, 775]. Dream [550, 553, 809]. Dreamer [586]. Dreamers [550]. Dreaming [550]. Dreams [552]. Dredging [837]. Drinking [715]. Driven [762, 797]. Drude [696]. drugs [1013]. Duality [865]. Duhem [15, 18, 19]. during [1043, 841, 709, 501, 240, 637, 939, 690, 1044, 804, 1042, 587]. Dust [856]. Dutch [467, 968, 1010, 344]. Duties [291]. Dyeing [1038]. Dyestuffs [62, 1038]. Dying [715, 642]. Dynamics [94, 698, 615, 133, 643]. Dynasty [818].

Early [504, 604, 446, 448, 418, 552, 512, 571, 520, 638, 553, 259, 943, 483, 519, 856, 579, 668, 47, 704, 317, 584, 967, 962, 994, 1028, 819, 120, 965]. Early-modern [584]. Earth [349, 422]. Earthquake [728, 732]. Earthquakes [733, 729, 605]. East [417, 351]. Eastern [954]. Ebba [819]. Eclipses [801]. Economic [299, 820, 345, 257, 690, 164, 587, 344, 984]. Economics [661, 343, 981, 869, 739, 720, 315, 153, 980, 982, 979, 983]. Economies [389, 407, 256]. Economist [620, 633]. Economists [639]. Economy [447, 565, 197, 161, 930, 811, 628]. Edgar [838]. Edith [720]. Edition [228, 818]. Editions [464]. Editor [482, 591, 512, 355, 368]. Editorial [1, 659, 818]. Editors [219, 509, 634, 44, 171, 549, 112]. Editorship [787]. Edizione [393, 411]. Edmund [688]. Educating [964, 976, 603]. Education [229, 714, 637, 967, 972]. Edward [477]. Effect [204, 922]. effects [1003]. Egede [973]. Egg [638]. Egg-Forms [638]. Egypt [637]. Egyptian [489]. Ehrlich [62]. Eighteenth [882, 703, 679, 529, 474, 625, 629, 608, 216, 372, 190, 134, 873, 607]. Eighteenth-Century [703, 529, 474, 625, 629, 216, 190, 134, 607]. Einfühlung [748]. Einstein [121, 124, 117, 122, 622, 118, 240, 114, 503, 115, 125, 689, 737, 361, 113, 572, 123, 287]. Electricity [439, 197]. Electroencephalography [441, 897]. Electron [943, 941]. Electrotechnological [439]. Elements [831, 429, 419, 484]. Elephant [384, 402]. Eliakim [284]. Elimination [841]. Elkana [758]. Ellen [610, 1053]. Elliptical [485]. Elusive [656]. Emanuel [692]. Embodiments [440]. Embrace [525]. Embryology [88]. Emergence [813, 763, 861, 449, 950, 631, 133, 941, 234]. Emergency [72, 732]. emigration [568]. Emil [988, 395, 413]. Emile [725, 724]. Emmy [539]. Emotion [742]. Emotions [442]. Empathy [744, 745, 743, 742, 746, 747]. Empire [914, 536, 530]. Empirical [343, 967]. Empiricks [681]. Empirico [833]. Empowering [447]. Enacting [612]. encounters [1054]. Encyclical [318]. End [145, 608, 532, 1032, 1031]. Ends [836, 1035]. Endurance [805]. Enduring [601]. Enemies [369]. Energy [762, 797, 120]. engagement [605]. Engineer [383, 401, 382, 400]. Engineer-Scientist [382, 400]. Engineering [944, 721, 320, 107]. Engineers [823]. England [964, 976, 597, 83, 450, 76, 965]. English [1043, 328, 377, 898]. Enharmonic [453]. Enlightenment [783, 86, 624, 177]. Enthusiasm [177]. Entrenching [243]. Environment [155, 859]. Environmental [828, 130]. Envisioning [371]. Envy [256]. Epic [225]. Epidemiological [957]. Epidemiologist [809]. Epigenetic [632]. Epilogue [843]. Epistemic [501, 858, 88, 927, 921, 847]. Epistemological [892, 541]. Epistemologies [643, 644]. Epistemologist [893]. Epistemology [212, 41, 595, 216, 36, 548, 341, 1010]. Equation [26]. Equations [886, 989]. Era [495, 887, 500, 198]. erfand [1057]. Erfinder [1060, 1058]. Erkenntnisproblem [227]. Erlanger [357]. Ernst [892, 893, 357, 363, 891, 359, 360]. Erratum [1028, 797, 899]. Error [644]. Essay [535, 172]. Essence [4]. Establishment [46, 271]. Eternal [263]. Ether [512, 1033]. Ethical [308, 72, 310]. Ethics [560]. Ethological [876, 173]. Ethology [266]. Ethos [503]. Etienne [236]. Etienne-Francois [236]. Euclid [421, 484]. Euclidean [424]. Euclides [531]. Eudoxan [486]. Eudoxean [487]. Eugene [574]. Eugenic [311, 313, 321]. Eugenics [305, 861, 314, 309, 179, 507, 318, 319, 315, 958, 180, 316, 317, 1042]. Europe [617, 954, 962, 446, 175, 573, 608, 778, 192]. European [174, 350, 466]. Eusapia [1014]. Evaluation [1045]. Eve [177]. even [533]. Event [950, 1004]. Events [573]. Ever [625]. Every [145]. Evidence [981, 1051]. Evolution [604, 815, 1043, 982]. Evolutionary [671, 561, 557]. Evolutionism [35]. Evolving [783]. exactness [648]. Examination [538]. Examinations [569]. Examining [321]. Example [886, 574]. Examples [861, 199]. Exceeding [485]. Excerpt [392, 410]. Exchange [869]. Exchanges [349]. Exclusion [244]. Exemplarity [882]. Exemplified [669]. Exerimental [181]. Exhibit [706]. Exhibit-Makers [706]. Exhibits [706]. Exile [113]. existence [1062]. Existing [920]. Exotic [705]. Expansion [212]. Expectations [904]. Experience [146]. Experiment [447, 29, 119, 22, 109, 975, 972]. Experimental [27, 391, 409, 392, 410, 105, 235, 152, 856, 144, 23, 875, 975]. Experimentalists [475]. Experimentation [762, 797]. Experimenting [340]. Experiments [762, 797, 456, 1015, 1052, 143]. Expert [328, 828, 761, 823]. Expertise [930, 107]. Explaining [1033, 430]. Explanation [255, 96]. Explanations [61]. Explorations [289]. Exploratory [762, 797]. Explosive [517]. Expressed [491]. Expression [173]. Extending [771]. Extent [305]. exterminated [450]. External [675, 577]. externalism [983]. Extinction [642]. Extroverted [300]. Eye [696]. Eyes [537, 537].

F [632]. Fable [543]. Fabri [606]. Fabrication [764]. facilitators [974]. Facing [298]. Factory [163, 73]. Facts [651]. faire [683]. Faith [595, 949]. Fall [565, 33, 460, 384, 402, 156]. Falling [485]. Fame [737]. Family [599]. Fascination [662]. Fascism [868]. fashioning [988]. fasting [1012]. Fatal [525]. Fate [495, 625]. Fatigue [649]. Faulhorn [650]. Fäulniss [434]. Favaro [393, 411]. Fear [256]. Features [38]. Feed [904]. Feel [916]. Feijó [530]. Female [101]. Feminist [256]. Fermentation [435]. Fernandez [493]. fertilizer [1037]. Few [252]. Fictions [28]. Fidelity [931]. Field [849, 123, 287, 708]. Fields [703]. Fifteenth [423]. Figure [131, 543]. Figures [664, 526]. Filling [134]. Film [716, 926, 927, 925, 554]. Films [928]. Filtration [811]. Financial [159]. Fingerprint [333, 847]. Fireworks [704]. First [225, 228, 462, 618, 690, 908, 765, 533, 116]. Fit [188]. Flattening [905]. Fleck [39]. Florence [1053]. Flourishing [492]. Food [1044, 721]. Force [601]. Forced [510]. Forces [920]. Forecasting [887, 643, 829, 644]. Foreword [219, 211, 304, 269]. Forgetting [552, 549]. Forging [457, 1044, 997]. Forgotten [390, 408]. Form [813, 803, 175, 152, 524, 632, 35, 1050]. Formation [14]. Formative [1054]. Forms [831, 356, 638, 360]. Fossil [537]. Foundations [674, 496, 7, 432, 83, 885, 739]. Four [420]. Fourteenth [423]. Fractured [819]. Fractures [508]. Fragmentation [516]. Framework [879]. France [264, 781, 868, 621, 698, 1011, 468, 614, 631, 724, 83]. Francis [625]. Franco [723, 784]. Franco-Soviet [723]. Francois [213, 236]. Francoist [760]. Franz [8]. Fraud [892]. Frederick [426]. Free [690, 867]. freedom [995]. French [550, 131, 429, 612, 904, 603, 898]. Frend [599]. Frend/De [599]. Fresnel [562]. Freud [549, 588, 550, 552, 300, 103, 874, 689, 557]. Freudenthal [367]. Front [9, 20, 31, 42, 58, 67, 77, 90, 98, 110, 128, 138, 149, 157, 169, 185, 195, 201, 209, 217, 230, 238, 247, 260, 267, 275, 285, 292, 302, 325, 337, 346, 353, 365, 375, 380, 398, 645, 657, 665, 672, 677, 684, 694, 700, 710, 718, 726, 734, 740, 750, 756, 767, 776, 785, 798, 806, 816, 825, 834, 844, 853, 862, 870, 877, 883, 889, 900, 906, 911, 918, 923, 932, 937, 945, 952, 960, 969, 977, 985, 991, 998, 1006, 1017]. Front [1029, 1039, 1046, 1055, 587]. Frontier [536]. Frontiers [951]. Function [173, 632]. Functional [626, 629]. Fundamental [943]. Fundamentalism [192]. Further [53]. Fusion [93]. Futurism [512].

Gährung [434]. Galen [525]. Galilean [394, 412, 127]. Galilée [388, 406]. Galileo [389, 407, 386, 404, 385, 403, 393, 411, 387, 405, 142, 383, 401, 428, 125, 382, 384, 400, 402, 395, 413, 47, 12, 297]. Galois [670]. Gans [282]. Gap [1045]. gardarensis} [815]. Garde [192]. Gardens [377]. Gassendi [388, 406]. Gatekeeper [866]. Gauge [502]. Gaze [565, 771]. Gazing [385, 403]. Geburtstag [1063]. Geer [819]. Geertz [578]. Gel [811]. Geminus [674]. Gender [474]. Genealogical [653]. Genealogy [859, 989]. General [491, 361, 493, 540]. Generality [664, 491, 526]. Generalization [902]. Generation [576, 632]. Genes [306]. Genesis [360, 92]. Genetic [905, 861, 547, 308, 314, 312, 851, 309, 507, 315, 311, 320, 310, 307, 316, 1022]. Genetics [861, 84]. Genius [34]. Genotype [307]. Genotype-Phenotype [307]. Geoffroy [236]. Geognosy [653]. Geograph [1058]. Geographer [1058]. Geography [1000, 696, 914]. Geological [730]. Geology [819]. Geometric [543]. Geometrical [664, 526, 420, 886, 879]. Geometry [432, 431]. Gerard [819, 545]. Gergonne [865]. German [547, 1054, 438, 1057, 89, 476, 984, 1059, 832, 600, 612, 1060, 88, 724, 782, 1058, 1062, 1063, 324]. Germany [439, 515, 469, 698, 118, 369, 614, 370, 850, 1051, 428, 1052]. Gerson [273]. Gertrude [749, 637]. Gesell [716]. Gesture [252]. Getting [580]. Gideon [367]. Gimignano [176]. Giordano [1002]. Giovanni [1012, 176]. Giving [956]. Glarus [653]. Glass [858, 982]. Glisson [625]. Global [763, 1013, 1001]. God [600, 753, 1020, 51, 1005]. God-Nature [600]. Godly [3]. Goethe [161, 544]. Goldstein [364, 363]. Good [115, 470]. Gor’kiy [504]. Gottfried [126]. Göttingen [500]. Governing [165]. Government [507]. gradually [1035]. Grammars [584]. Graphic [60]. Gravitation [117]. gravity [1005]. Great [455, 732]. Greek [424, 667, 674, 490, 427, 660]. Greenfield [1008]. Greenwich [1035]. Grey [441]. Grimaud [630]. Grimm [832]. Grossmann [8]. Grounding [934]. Growth [63]. Guevara [47]. Guillaume [630]. Gurevich [17]. Gustave [620, 633]. gynecological [1042]. Gynecology [182]. Gyorgy [518].

H1N1 [822]. ha-Penini [272]. Habitus [472]. Hagia [636]. Half [697]. Haller [627, 625]. Hand [167]. Hands [385, 403, 717, 962]. Hans [359, 973]. Hansen [763]. Happened [625]. Hard [533, 987]. Harmonious [476]. Harnessing [929]. Harris [506]. Hart [284]. Harvard [639]. Hate [458]. Hathaway [840]. Hayek [661]. Healing [179, 1011]. Health [822, 616, 954, 914, 514]. Healthy [564]. hearts [996]. Hebrew [688, 503, 271, 270, 584]. Hegemony [533]. Heiberg [421]. Heidegger [258]. Heideggerian [570]. Heinrich [669]. Heisenberg [481]. Hellenistic [860]. Heller [281]. Helmholtz [833, 433, 432, 434, 430, 857]. Henri [921]. Henrique [283]. Henryk [8]. Herder [543]. Hereditary [308]. Hermann [1021, 419, 833, 874, 857, 502]. Hermeneutics [2]. Hero [477]. Hessen [1024]. Hidden [522]. High [762, 797, 656, 852, 899]. High-Altitude [656]. High-Energy [762, 797]. Higher [229]. Hilbert [301, 263, 500]. Hill [887]. Hindu [418]. Hippocratism [860]. Historia [493]. Historiality [144]. Historical [610, 464, 465, 333, 528, 748, 105, 563, 803, 141, 935, 146, 80, 582, 974, 289]. Historicizing [100, 324, 995]. Histories [728, 471, 855, 556, 660, 607]. Historiographical [490, 428]. Historiography [549, 916]. History [634, 703, 348, 409, 410, 174, 497, 294, 281, 459, 670, 649, 692, 790, 934, 30, 328, 611, 131, 950, 789, 762, 797, 638, 452, 227, 742, 886, 527, 791, 926, 147, 583, 796, 52, 259, 483, 519, 525, 87, 372, 795, 916, 793, 847, 702, 81, 65, 1034, 975, 972, 1001, 706, 1062, 391, 392]. Hitherto [689]. Hohenstaufen [426]. Holding [902]. Holes [117]. Holism [630]. Holland [492]. Home [916]. Homeomeric [667]. Homo [815]. honor [1020]. Honoré [606]. Hooke [60, 577]. horse [1015, 606]. Hospitals [568]. Hotel [650]. Hours [629]. House [866]. Household [866]. Households [472]. Hugo [839, 744, 40]. Hui [664, 526, 989]. Hult [819]. Human [905, 821, 306, 446, 815, 335, 872, 107, 173, 548]. Human-Animal [872]. Humans [456]. Humidity [669]. Hungarian [1042]. Hunting [384, 402]. Hurtado [535]. Huxley [265]. Hydraulic [823]. Hyperbolical [485]. Hypnosis [551]. Hypotheses [738].

ibn [424, 664, 526]. Icon [588]. Iconoclastic [252]. Idea [97]. Ideal [86, 429]. Idealizations [28]. Ideals [603, 802]. Ideas [495, 167]. Identification [333, 221]. Identities [898, 761]. Identity [333, 823, 972]. Ideological [949]. Ideology [143, 689, 332]. Idiocy [603]. Idiosyncrasies [803]. Idiots [603]. ignorabimus [987]. ihres [1062]. II [426, 465, 266, 505, 512, 947, 850, 804, 644]. Illness [514, 1051]. Illumination [682]. illustrated [1010]. Il’ya [456]. im [1062]. Image [713, 748, 457, 743, 555, 222]. Images [229, 225, 334, 714, 722, 37, 34]. Imaging [856]. Immanent [254]. Immigrant [565, 566]. Immortality [264]. Immune [205]. Impact [559, 73]. Imperial [732, 783, 880, 369, 370, 881]. Impetus [12]. Implications [130]. Importance [357, 642]. Impossible [103]. Impulse [594]. Impure [216]. Imre [874]. Inbred [308]. Incarnation [503]. Including [689]. independence [971]. Independent [539]. Index [208, 416]. Indian [537, 37]. Indias [493]. indicator [995]. Individual [305, 307]. Individuals [1024]. Industrial [643, 189]. Industrializing [709, 536]. Industry [616, 904, 721, 191, 62]. Inertia [606, 96]. Infant [716]. Inference [635]. Infinitesimal [852, 899]. infinity [1002]. Influence [690, 818]. influenced [1042]. influential [995]. Information [1043, 838, 206]. Informational [604]. infrastructural [1023]. Infrastructure [809]. ingénieur [131]. Inherited [653]. Inhibition [102]. Inoculating [566]. Inoculation [882]. Inquiring [663]. inscription [1050]. Insect [370]. Insights [935, 849]. Inspection [565]. Inspire [106]. Instability [867]. Instead [41]. Institute [503, 903, 904, 613]. Instrument [289, 24, 193, 648]. Instrumental [168]. Instruments [349, 60, 893]. Integration [676, 955]. Intellect [101]. Intellectual [449, 916, 1043]. Intellectuals [291, 65]. Interaction [722]. Interactions [626, 572]. Interdisciplinary [869]. Interested [813]. Interests [351]. interior [1049]. International [229, 116, 814, 698, 637, 755, 761]. Internationalism [688]. Internment [569]. Interplay [119]. Interpretation [60, 255, 752, 818, 137]. Interpretations [122]. Interpreting [588]. Intersections [572]. Interspecific [560]. Intertextual [132]. Interventions [312]. Interview [1049]. Interviewer [709]. Interwar [781, 617, 954, 89, 939, 942, 815, 940, 957, 959, 780]. Intimate [881]. Introduction [327, 808, 249, 647, 228, 728, 497, 528, 686, 79, 471, 277, 44, 203, 855, 788, 512, 171, 444, 769, 893, 979, 939, 452, 355, 742, 140, 549, 368, 872, 542, 483, 220, 112, 187, 778, 913, 510, 702, 594, 925, 624, 1024]. Introductions [589]. Introductory [692, 5, 232, 390, 408, 16]. Introverted [300]. Intuition [259, 544]. intuitionism [1021]. Intuitions [578]. Inuit [1000]. Invariants [501]. invented [1057]. Inventing [815]. Invention [389, 407, 214, 114, 455]. Inventions [811]. Inventor [1058]. Inventors [114, 1060]. Inventory [840]. Invents [891]. Investigates [588]. Investigation [71, 574, 575]. Investigations [374]. Investigators [476]. Invisible [369, 373, 514]. Irenism [282]. Ironic [95]. Ironies [319]. Irrational [89, 909]. Irritability [627, 625]. Isaac [1020]. Isfizari [424]. Islamic [14, 352, 902, 320]. Israel [690, 689, 310]. Israeli [182, 690]. Isserles [279]. issue [9, 10, 20, 21, 31, 32, 42, 43, 58, 59, 67, 68, 77, 78, 90, 91, 98, 99, 110, 111, 128, 129, 138, 139, 149, 150, 157, 158, 169, 170, 185, 186, 195, 196, 201, 202, 209, 210, 217, 218, 230, 231, 238, 239, 247, 248, 260, 261, 267, 268, 275, 276, 285, 286, 292, 293, 302, 303, 325, 326, 337, 338, 346, 347, 353, 354, 365, 366, 375, 376, 380, 381, 398, 399, 645, 646, 657, 658, 665, 666, 672, 673, 677, 678]. issue [684, 685, 694, 695, 700, 701, 710, 711, 718, 719, 726, 727, 734, 735, 740, 741, 750, 751, 756, 757, 767, 768, 776, 777, 785, 786, 798, 799, 806, 807, 816, 817, 825, 826, 834, 835, 844, 845, 853, 854, 862, 863, 870, 871, 877, 878, 883, 884, 889, 890, 900, 901, 906, 907, 911, 912, 918, 919, 923, 924, 932, 933, 937, 938, 945, 946, 952, 953, 960, 961, 969, 970, 977, 978, 985, 986, 991, 992, 998, 999, 1006, 1007, 1017, 1018]. issue [1029, 1030, 1039, 1040, 1046, 1047, 1055, 1056]. Issues [601, 872]. Italian [278, 982]. Italy [391, 409, 392, 410, 474, 779, 605]. Ivan [495, 103]. Ivanov [456]. Ivory [199].

J. [480]. Jacob [586, 283]. Jacobi [936]. Jahr [1062]. James [763, 103]. Japan [941, 732, 319]. Japanese [934]. Jean [288, 246, 630]. Jean-Baptiste [288, 246]. Jean-Charles-Marguerite-Guillaume [630]. Jerusalem [503]. Jesuit [11, 46, 538, 463, 607]. Jesuits [47]. Jewish [588, 178, 446, 449, 692, 955, 908, 693, 284, 585, 278, 569, 959, 321]. Jews [322, 270]. João [530]. Job [849]. Johann [669]. John [119, 596, 577]. Joining [492]. Joint [634]. Journal [588]. Journalism [620, 633]. Journals [634]. Judith [506]. Juggling [761]. Julius [753]. June [1062, 1063, 820]. Jung [300]. Juni [1062, 1063]. Jurisprudence [936]. Just [115]. Justified [801].

Kansas [708]. Kant [298]. Kanto [732]. Karl [540]. Kate [1020]. Keeping [866]. Kelley [1053]. Kenelm [576]. Kenneth [676]. Kepes [518]. Kepler [668, 329]. Key [627, 794]. Khinchin [510]. Khrushchev [783]. Killed [1033]. Kishon [691]. Kit [669]. Klein [500]. Klüver [518]. Knot [501]. Knowing [887]. Knowledge [610, 760, 229, 48, 27, 887, 228, 63, 770, 11, 474, 888, 769, 538, 917, 290, 507, 377, 69, 75, 582, 37, 74, 675, 567, 222, 560, 708, 594, 965, 653, 962, 963, 1009, 968]. known’ [607]. Korea [571, 971, 881]. Koyré [200]. Kristensson [866]. Kristineberg [866]. Kurt [364, 108, 363]. Kuzari [274].

Lab [654]. Labor [722, 675]. Laboratories [551]. Laboratory [647, 839, 687, 814, 858, 743, 71, 163, 809]. Ladenness [762, 797]. Laissez [683]. Laissez-faire [683]. Lamarck [246]. Lambert [669, 1020]. land [1054]. Landau [688]. Language [233, 369, 179]. Lapland [492]. Large [981, 130, 830]. Large-Scale [981]. Last [225]. László [515]. Late [723, 898, 362, 222, 988, 1010, 732]. Latent [333]. Later [174, 596]. Latin [175, 425, 528, 427, 1002]. Law [336, 691, 755, 335, 109, 384, 402, 331, 1054, 96]. Laws [388, 406]. Lawyer [330]. Lay [447, 708, 702]. Learned [472, 270, 1016]. Learning [962, 175, 967, 350]. Leaves [367, 787]. Lecture [622, 623]. leden [788]. Legacy [926]. Legal [332, 329, 1023]. Legend [135]. Legitimacy [766, 765]. Legnano [596]. Leibniz [1060, 1060]. Leopoldo [534]. Lessing [543]. Lessons [699]. Lestschinsky [586]. Letter [318, 562, 922]. Letters [86, 689, 475, 364, 363, 351]. Level [307]. Levels [1015]. Levi [273]. Leviathan [857]. Lévy [200]. Lewin [108]. Lexical [356]. Li [818, 989]. Liberal [848]. Liberating [568]. Library [75]. Life [439, 714, 291, 206, 516, 764, 640, 1032, 1037]. Light [48, 117, 562]. Lightning [330]. Limitations [307]. Limits [27, 215]. line [1035]. Lineage [321]. linear [990]. Linearity [63]. Lines [667, 902, 708]. Linguistics [356]. Linking [721]. Linnaean [492]. Lipmann [281]. Literacy [770]. Literary [84]. Literature [453, 550, 542, 544, 548]. Liu [664, 526]. Lives [283, 352, 630]. Local [887, 1001]. Locating [611, 374]. Locke [577]. Łód’z [848]. Logic [495, 270, 372, 450]. Logical [440]. lonely [1036]. Longing [508]. Longitude [619]. Looking [773, 97]. Loose [749]. Lord [284]. Lorentz [115]. Lost [551]. Love [458, 441]. Love-Hate [458].

Love-Machine [441]. loved [607]. Lowering [308]. Lucien [200]. Ludwik [39]. Lugones [534]. Luigi [697]. Lunar [908, 486]. Luria [496]. Luso [530]. Luso-Brazilian [530].

Mach [892, 893, 891]. Machine [458, 441, 193, 940]. Machines [713, 25]. Macro [650]. Macro-Tool [650]. Made [651]. Madrid [729]. Magazin [298]. Magazines [780, 266]. Maghrib [423]. Magic [117, 182]. Magician [225]. Makers [706, 24]. Making [759, 821, 703, 769, 370, 755, 243, 160, 500, 477, 827, 598, 830]. Malariology [567]. Malthus [299]. Man [458]. Man-Machine [458]. Management [723, 643]. Managerial [154]. Mandatory [567]. Mannheim [290]. Mansion [1041]. Manufacture [7]. Manufacturing [154]. Maps [914, 966]. Marcel [545]. Marcello [605]. Margaret [754]. Marguerite [630]. Marian [1025]. Mark [629, 26]. Marker [307]. Market [699, 616, 690, 940, 829]. Marketing [721]. Marketplace [167]. Markets [191]. Markov [573]. Marple [506]. Marx [257]. Mason [602]. Masriera [784]. Mass [120]. Mass-Energy [120]. Material [772, 893, 641]. Materialism [629, 1027]. Mathematical [343, 262, 63, 501, 489, 638, 271, 580, 427, 535, 818, 258, 463, 153, 990]. Mathematicians [697, 804, 510, 352]. Mathematics [667, 674, 464, 264, 490, 491, 465, 263, 497, 944, 851, 698, 131, 498, 571, 14, 82, 503, 886, 601, 483, 132, 898, 83, 450, 500, 23, 660, 885, 1002, 503]. Matriarch [866]. Matter [950, 629, 898, 632, 9, 10, 20, 21, 31, 32, 42, 43, 58, 59, 67, 68, 77, 78, 90, 91, 98, 99, 110, 111, 128, 129, 138, 139, 149, 150, 157, 158, 169, 170, 185, 186, 195, 196, 201, 202, 209, 210, 217, 218, 230, 231, 238, 239, 247, 248, 260, 261, 267, 268, 275, 276, 285, 286, 292, 293, 302, 303, 325, 326, 337, 338, 346, 347, 353, 354, 365, 366, 375, 376, 380, 381, 398, 399, 645, 646, 657, 658, 665, 666, 672]. matter [673, 677, 678, 684, 685, 694, 695, 700, 701, 710, 711, 718, 719, 726, 727, 734, 735, 740, 741, 750, 751, 756, 757, 767, 768, 776, 777, 785, 786, 798, 799, 806, 807, 816, 817, 825, 826, 834, 835, 844, 845, 853, 854, 862, 863, 870, 871, 877, 878, 883, 884, 889, 890, 900, 901, 906, 907, 911, 912, 918, 919, 923, 924, 932, 933, 937, 938, 945, 946, 952, 953, 960, 961, 969, 970, 977, 978, 985, 986, 991, 992, 998, 999, 1006]. matter [1007, 1017, 1018, 1029, 1030, 1039, 1040, 1046, 1047, 1055, 1056]. may [1032]. McCarthyism [720]. McCulloch [440]. Meaning [214, 507, 102, 188, 1043]. Meanings [897]. Means [836, 243, 840]. Measure [638]. Measure-Bodies [638]. Measurement [669, 97, 167]. Measuring [569]. Mechanical [682, 61, 1027, 133, 3]. Mechanicism [35]. Mechanics [424, 516, 940, 47, 12, 936]. Mechanisms [244, 940]. Mechanistic [6, 7]. Media [951, 1050]. Mediated [967]. Mediating [25]. Medical [178, 861, 627, 563, 621, 565, 177, 179, 1013, 567, 1042, 180, 176, 956, 624]. Medicine [634, 824, 808, 813, 184, 860, 954, 692, 175, 618, 520, 72, 903, 525, 930, 183, 566, 640, 172, 630, 1011]. Medieval [801, 902, 417, 175, 271, 270].

Medium [439]. Melancholia [529]. Melo [802]. Memex [577]. Memoirs [146]. Memorial [479]. Memory [545, 146, 680, 554, 577, 1008]. Memory-Science [545]. Men [3]. Mendel [559]. Mendoza [535]. mental [1051]. Meridian [1035]. Merlatti [1012]. Merton [46, 290, 55, 56, 53, 54]. Mesopotamian [635]. Metaphor [343, 30]. Metaphors [48, 547, 458]. Metaphysics [624, 524]. Metapsychology [145]. Meteorology [760, 654]. Method [409, 410, 41, 851, 34, 1011, 391, 392]. Methodological [558, 38, 69, 736]. Methodology [661, 333]. Methods [461]. Métraux [787]. Metrology [910]. Mexico [905, 897]. Michael [291]. Michel [75]. Microcosm [650]. Micrographia [60]. Micronesia [1054]. Microscope [714]. Microscopic [715, 374]. Microscopy [372]. Mid [374]. Mid-Nineteenth-Century [374]. Middle [174, 898, 596, 222, 351, 427]. Midrashic [281]. Migration [804]. Migrations [639]. Miguel [535, 784]. Militarization [760]. Military [156]. Milk [813]. Mind [100, 836, 440, 842, 76, 437]. Mindful [436]. minds [962, 996]. Mineralogical [530]. Minerals [522]. Ministerial [242]. Minnesota [840]. Minorities [954]. Miracle [470]. Miracles [598]. Miriam [1008]. Miserable [564]. Mises [510, 508]. Misfortune [199]. Miss [506]. Missing [560]. Mission [739]. MMR [699]. Mobility [804]. Mobilization [494]. Mode [136, 708]. Model [821, 722, 24, 358, 931, 829, 680]. Modeling [774, 828, 829, 344]. Models [821, 28, 249, 343, 345, 61, 869, 580, 430, 339, 827, 341, 830]. Modern [490, 348, 228, 263, 497, 682, 356, 446, 858, 565, 448, 512, 357, 638, 662, 383, 401, 527, 234, 818, 916, 92, 965, 704, 35, 967, 962, 1027, 1036, 584]. Modernism [148]. Modernity [319, 508]. Modernization [928]. Modernizing [848]. Modes [828]. Moholy [515]. Moholy-Nagy [515]. Molecular [547, 604, 206, 84, 183]. Molecules [604]. Molinari [620, 633]. monde [514]. Monism [600]. Monitoring [733, 613]. Monopoly [389, 407]. Monroy [605]. Mont [656]. Mont-Blanc [656]. Montaigne [75]. Montgolfiers [117]. Montpellier [631, 628]. Moons [386, 404]. Moral [447, 312, 345, 226, 811]. Morality [34]. Morals [522]. Mores [321]. Morgan [450, 599]. Moshe [279]. Mother [813]. Mothers [470]. Motion [432, 902, 422, 388, 406]. Motives [499]. Motu [142]. Mount [650, 1037]. Mountain [652]. Mountains [647, 649, 651]. Mountainside [648]. Moved [546, 680]. Movement [743, 317]. Movements [417]. Movie [929]. Moving [748, 714, 546]. Multiphasic [840]. Multiplex [342]. Münsterberg [839, 744]. Murray [754]. Museum [706, 602]. Music [453, 476, 222]. Musical [193]. Mutual [881]. My [795]. Myograph [857]. Myopic [602]. Myriad [917]. Mysterious [523]. Myth [147, 747].

Nagy [515]. ‘Names [607]. Naming [876, 493]. Narration [547, 144]. Narrative [148, 141, 745, 929, 152, 147, 1049]. Narratives [387, 405, 837, 729, 1044]. Nation [697]. National [87, 85, 778]. Nationalism [89]. nationalist [958]. Nationality [531]. Natural [453, 703, 706, 681, 281, 282, 33, 161, 600, 926, 597, 2, 246, 97, 801, 3, 331, 596, 594, 704, 607, 531, 493, 1020, 1019, 972]. Naturalism [558]. Naturalist [703]. Naturalized [101]. Nature [48, 661, 647, 446, 920, 578, 51, 161, 476, 600, 580, 732, 183, 928, 278, 522, 435]. naturelle [602]. Naturforscher [476]. Navigation [619]. Nazi [428]. Nazionale [393, 411]. Needs [145]. Negative [312]. Neglected [131]. Nerval [545]. Nervous [216, 857]. Netherlands [467]. Network [708]. Neumann [126]. Neuroanatomy [438]. Neurology [363]. Neurons [440]. Neurophysics [216]. Neuropsychology [496]. Neuroscience [440]. Neurosciences [436]. Neurosis [875, 443, 975]. Neutron [120]. Never [680]. Newly [39]. Newspapers [784]. Newsprint [619]. Newton [1020, 117, 1005, 96, 601, 34]. Newtonian [33, 283, 127]. Newtonianism [284]. Next [705]. Nexus [824, 808]. Nicholas [581]. Nicht [1057]. Nicolas [264]. Nicolaus [831]. Nineteenth [856, 579, 910, 988, 1041, 453, 731, 697, 513, 611, 498, 479, 571, 476, 603, 602, 377, 564, 132, 532, 374, 331, 651, 653]. Nineteenth-Century [856, 910, 731, 611, 498, 571, 603, 602, 377, 132, 331, 651, 653]. No [2, 537, 1004]. Noether [539]. Non [772, 508, 643, 644]. Non-Conformist [508]. Non-Forecasting [643, 644]. Non-Material [772]. Nonlinear [504]. Nordic [810]. Normal [243]. North [821, 996, 812, 729]. North-American [821, 996]. Norway [459, 814, 460, 973]. Norwegian [703, 505]. nosology [1000]. nostalgia [1036]. Notations [736]. Note [509, 591, 5, 227, 390, 408, 183, 16, 857, 232]. Notebooks [773]. Nothing [181]. Notion [886]. Notions [576]. nova [329]. Novalis [82]. Nuclear [951, 93]. Number [688, 223]. Numbers [97]. Numeracy [770]. Numerical [668, 736]. nursing [959]. Nurture [578].

o [227]. obituaries [1033]. Object [813, 833, 628]. Objective [334]. Objectivity [481, 621, 1023]. Objects [357, 551, 339, 546, 1032, 1031, 1038]. Observation [656, 702, 704]. Observations [389, 407, 419, 705, 423]. Observatory [650, 887, 656]. Observed [237]. Observer [703]. Observers [546, 708]. Obstacles [494]. Occasionalism [948]. Occupation [724]. Odilon [514]. Odoyevskiy [453]. off [1038]. Office [113]. Oikos [858]. Oka [683]. Old [915, 913]. Old-New [913]. Olive [523]. One [630, 1020]. Onegin [574]. Ontological [805]. Ontology [864]. OP [176]. Opening [688, 622, 583, 623]. Operational [234]. Operationalism [27]. Operators [621]. Opinion [437]. Oppenheimer [480, 572]. Opposites [846]. Optical [636]. Optics [521, 431]. Oral [500]. Ordained [753]. Order [223, 29, 102]. Organic [626, 324]. Organical [61]. Organism [626, 859]. Organisms [821, 244]. Organisms-Mechanisms [244]. Organization [603, 37, 653, 822]. Organizational [154]. Oriental [910]. Orientalist [1058]. Origin [151, 502]. Original [935]. Origins [333, 263, 382, 400, 130, 344]. Orlov [495]. Ornithology [850]. Orthodox [178]. Oscar [696]. Oscillating [662]. Oscillation [744]. Other [589, 104, 123, 287, 54]. Other-Regard [104]. Otto [851]. Oulipo [264]. Our [443]. Outlining [775]. Overcoming [494]. overshadowed [1035]. Overview [512]. Oviedo [493]. Owen [265]. Own [713, 76, 1033]. Oxford [13].

P [739]. Pacific [564]. Pacífico [494]. pacifist [240]. Packaging [617]. Pain [174, 173]. Paintings [513]. Palaeontology [537]. Paleontology [532]. Palestine [687, 567]. Palliative [181]. Papal [880]. Paper [851, 772, 510, 957]. Papers [784]. Paperwork [771]. Parabolic [396, 414]. Paracelsianism [445]. Paracelsus [445]. Paradigmatic [491, 94, 243]. Paradigms [200]. Paradox [442]. Parallel [484]. Paraquaria [607]. Paris [670, 892, 602, 1012]. Parisian [873]. Parr [972]. Part [464, 465, 643, 644]. Participation [702]. Particle [762, 797, 342]. Particulars [519]. Partnership [819]. Party [539]. Pascal [1057, 1057]. Passions [631]. Past [507]. Pasteur [616, 687, 903, 613]. Pasteurian [514]. Pastorian [903]. Patent [755, 113]. Patents [165]. Paternity [332]. Path [145]. pathes [788]. Pathology [529, 363, 1000]. Patients [178]. Patinkin [690]. Patocka [137]. Patronage [637]. pattern [1023]. patterns [980]. Paul [773, 1041, 62]. Paving [913]. Pavlov [458, 103]. Peak [655]. Peculiarities [85]. Pedagogics [137]. Pedagogy [643]. Penini [272]. Penrose [720]. People [618]. Perception [173]. Perils [288]. Period [696, 939, 940, 959]. Periphery [492]. Perish [329]. persistence [979]. Person [184]. Persona [476]. Personae [471]. Personal [791, 26, 729]. Personality [306, 837, 840]. Perspective [417, 148, 570, 1021]. Perspectives [105, 541, 898]. persuasion [990, 972]. Persuasive [929]. Perverted [300]. Peter [270]. Petty [288]. Phantom [487]. Pharmacy [618]. Phenomena [801, 360]. Phenomenon [359]. Phenotype [307]. Phenotypic [305]. philhellenism [988]. Philology [592, 590]. Philosopher [291, 108]. Philosopher-Psychologist [108]. Philosophers [681, 3]. Philosophical [588, 496, 499, 680, 1021, 96]. Philosophisches [298]. Philosophize [510]. Philosophy [495, 228, 41, 356, 282, 168, 30, 7, 227, 363, 3, 278, 1020, 1024, 95]. Photograph [856]. Photographs [513, 892]. Photography [515, 896, 895, 652]. Phylloxera [370]. Physical [174, 499, 431, 12]. Physicians [693, 964, 976]. Physicist [36]. Physics [136, 951, 207, 944, 512, 221, 941, 762, 797, 939, 942, 601, 922, 342, 943, 516, 37, 856, 125, 1043, 1027, 116, 947]. Physiological [458, 431, 921]. Physiology [988, 458, 640]. Picture [6]. Picturing [513]. Pieces [989]. Pierre [15, 18, 19, 288]. Pietism [56]. Piezoelectricity [942]. Pine [904]. Pious [49]. Place [887, 70, 227, 69, 75, 898, 76]. Placelessness [656]. Plan [640]. Planetary [902]. Planning [690]. Plant [696]. Planting [454]. Plants [753, 947]. Plasticity [644]. Plata [607]. Plate [856]. Plato [664, 526]. Plausibility [172]. Please [1009]. Plücker [865]. Plural [898]. Pluralism [224]. Poetic [440, 546]. Poetry [534]. Poincaré [432, 936]. Point [226]. poison [1037]. Poland [957, 956, 959, 848]. Polanyi [291].

Polemics [865, 587]. Police [379]. Policy [345, 690, 566, 830]. Polish [279]. Political [763, 563, 621, 565, 639, 428, 993, 997, 996]. Politices [191]. Politics [781, 800, 493, 687, 954, 733, 691, 369, 441, 602, 190, 910, 253, 993]. Poncelet [865]. Poor [539]. Pope [270]. Popular [781, 780, 779, 782, 1016]. Popularization [784, 378, 430, 840, 802, 620, 633, 925]. Population [809, 133]. populations [1022]. Portal [715]. Portuguese [802]. Position [819]. Positive [312]. Possibilities [212, 134]. Possibility [635]. Post [850, 598]. Post-Tridentine [598]. Post-World [850]. postcolonial [971]. Postmodern [207]. Postmodernism [203, 204, 205]. Postmodernity [843]. Postulate [484]. Postwar [905]. Potential [527, 1034]. Power [753, 993, 779, 929]. Powerful [578]. Practical [603, 962]. Practice [1045, 100, 696, 182, 72, 235, 22, 234, 311, 23, 344]. Practices [760, 849, 990, 436, 638, 903, 104, 1042]. Practing [175]. Pragmatism [253]. Prague [282, 892]. Pre [348, 527, 689, 487]. pre-Eudoxean [487]. Pre-Modern [348, 527]. Pre-state [689]. Precedents [880]. precipitating [1004]. Precision [167, 648]. Predictive [305]. Preface [482]. prehistoric [1003]. Prehistory [641]. Prejudicial [334]. Preliminary [2]. Premodern [198, 879]. Prerequisites [892]. Present [905, 507, 137]. Present-Day [137]. Presentation [266, 832]. Presentations [1062]. Presentist [792]. press [956]. Pressure [815]. Prestige [815, 922]. Price [737]. Prime [1035]. Prince [800, 966, 453]. Princeton [501]. Principia [96, 601]. Principle [896]. Principles [127, 636]. Print [915]. Priority [296]. Probability [510, 1041]. Probable [582]. Problem [228, 546, 852, 899, 47, 987, 502]. Problematic [803]. Problems [541]. procedure [989]. Procedures [772, 773, 222, 1049]. Process [642]. Processes [698, 772, 830, 1003]. Producing [616, 810]. Production [136, 888, 73, 832, 288, 191, 708, 875]. Profession [162]. professionalization [1013, 959]. Professionals [655, 316]. Professions [842]. Professor [1057]. Professorial [473]. Program [425, 692, 739, 323, 628, 1011]. Programm [357]. Progressing [600]. Project [105, 440, 723, 103]. Projectile [396, 414]. Projecting [837]. Projection [744]. Projectionist [717]. Projects [723]. Proletarian [323]. promise [955]. Promises [604]. Proneness [609]. Proof [47]. Proofs [262]. Propaganda [914]. Propagation [857]. Proper [661]. Properties [629]. Proportion [487]. Proposals [693]. Propositions [517]. Prose [546]. Protestant [48, 317]. Proteus [480]. Prototype [358]. Proust [545]. Psychiatric [505]. Psychiatry [505, 805]. Psychic [439]. Psychoanalysis [145, 556, 549, 874, 378, 689, 548, 554]. Psychoanalytic [256, 551]. Psychobiology [89]. Psychodiagnosis [439]. Psychological [839, 743, 107, 1052, 569]. Psychologies [553]. Psychologist [108]. Psychology [609, 105, 833, 266, 106, 869, 671, 109, 535, 104, 921, 561, 506]. Psychometric [840]. Psychophysiology [433]. Psychosurgery [841]. Psychotechnics [839]. Ptolemy [902]. Public [783, 954, 621, 618, 950, 914, 166, 865, 198, 165, 931, 925, 619, 823]. Publication [880, 329]. Publish [329]. publishing [1003]. Pure [503, 153].

Puritanism [56]. purity [971]. Purloined [922]. Purpose [522]. Putrefaction [435]. Pythagorean [846].

Qing [818, 916, 913]. Quantification [882, 39]. Quantitative [160]. Quantum [122, 516]. Quarantine [564]. Quest [453, 246]. Question [579, 630, 1009]. Questions [499]. Quislingism [505]. Qurra [424, 664, 526].

R [496, 840]. Rabbi [281, 279]. Rabbis [279]. Racial [529, 958]. Radical [639, 958]. Radicalized [253]. radicals [450]. Raffaello [391, 409]. Random [461]. Randomness [70]. rapports [236]. Raspail [1011]. Rassenhygiene [323]. Ratio [464, 463]. rational [450]. Rationalism [281]. Rationality [168]. Re [578, 483, 75, 568]. Re-Created [75]. Re-emigration [568]. Re-reading [578]. Re-Writing [483]. React [517]. Reaction [626, 118]. Reactor [533]. Reader [550, 917, 584]. Readerships [780]. Reading [860, 578]. Realism [931]. Realities [178]. Reality [86, 23]. Reason [48, 682, 257, 51, 22, 732, 599]. Reasoning [96, 990]. Receding [874]. Reception [543]. Receptor [62]. Rechenmaschine [1057, 1059, 1058]. rechenmaschinen [1060]. Rechentechnik [1059]. Reconsidered [833]. Reconsidering [987]. Reconstruction [908]. recovering [994, 1028]. Red [585]. Rediscovering [803]. Redon [514]. Reductionism [133]. Reductionist [244]. Redux [843]. Reech [936]. Reference [132, 524]. References [397, 415]. Refiguring [472]. Reflected [697]. Reflections [336, 748, 794, 507, 886, 736]. Reflexive [104]. Reflexivity [63]. Reform [698, 640, 1053]. Reformation [387, 405]. Reformers [317]. refutation [990]. Regard [104]. Regarding [603, 486]. Regime [784]. Region [810, 955]. Regional [212, 533]. Registries [809]. Regulating [626, 614]. Regulation [721]. Reich [428]. Relation [419, 632]. Relations [237]. Relationship [307]. Relativism [224]. Relativistic [540]. relativitets [589]. relativitets-teorye [589]. relativity [1001, 623, 622, 118, 589, 737, 540, 546, 534]. Release [724, 725]. Relief [709, 955]. Religious [177, 176]. Relocating [903]. Remarkable [575]. Remarks [692, 895, 562, 434, 583, 484]. Remembering [552, 557]. Remote [881]. Renaissance [352, 909, 963]. Rendering [1000, 168]. Repair [859]. Replication [389, 407]. Reports [152]. Representation [610, 23]. Representations [352, 437, 222, 200, 350]. Representativeness [462]. Representing [1022]. Reprinted [623]. Reproduction [61]. Republic [351, 86, 475, 968, 1010]. Republican [802]. Research [821, 343, 308, 944, 501, 858, 692, 505, 769, 772, 612, 93, 37, 164, 738, 165, 811, 775, 587, 533, 1000, 984]. Resin [904]. Resistance [950]. Resource [809]. Response [634, 490, 728, 52]. Responses [281]. Restoration [3]. Resurrecting [204]. Rethinks [481]. Retina [856, 374]. Retreat [113]. Reverse [934]. Reversibility [867]. Revival [133]. Revolution [4, 454, 141, 514, 189, 587]. Revolutions [250]. Rewriting [497]. Reymond [988, 987]. Rhetoric [121, 490, 604, 143, 154, 641]. Rhetorical [763, 266].

Rhetorical/Political [763]. Rhythmic [222]. Ricardo [299]. Rich [506]. Richard [508]. Richards [1053]. Riddles [420]. Right [849, 788]. Rigor [83]. Ringelblum [692]. Río [607]. Rise [454, 565, 33, 197, 460, 608, 52, 873, 156, 1014]. Risk [822, 536]. Ritalin [470]. Ritual [362]. Road [595, 795, 540]. Robert [480, 447, 60, 429, 55, 23, 577]. Rod [330]. Rods [374]. Role [420, 784, 738]. Roles [830]. Roma [1022]. Roman [318]. Romanticist [832]. Roots [682, 290, 503]. Rotifer [372]. Royal [873]. Rudolfine [282]. Ruggiero [920]. Rui [818]. ruins [1036]. Russell [739]. Russia [783, 504, 459, 461, 190]. Russian [453, 454, 455, 452].

S [437]. S. [440]. S.J. [465]. Sabbath [116]. Sacred [178]. Sage [739]. Saint [598]. Saint-Making [598]. salmon [972]. Salvation [11, 113]. Samples [574, 575]. Sampling [459, 462, 460, 461]. Samuel [948, 477]. San [176]. Sanctioning [341]. Sarmento [283]. savant [131]. Scale [453, 981]. Scaliger [753]. Scaling [696]. Scandinavia [824, 808, 815]. Scandinavian [809, 812]. Scene [773]. Schalit [589]. Schemes [626, 619]. Schickard [1057, 1058, 1062, 1063, 1057, 1059, 1061, 1060, 1058, 1063]. Schickard-Symposion [1063]. Schickards [1059]. Scholar [495, 271]. Scholarly [949]. Scholars [472, 779]. Scholarship [585]. School [1043, 898, 504]. schoolchildren [1052]. Schools [852, 899]. Schwarzschild [540]. Science [783, 58, 59, 634, 824, 808, 100, 305, 225, 453, 495, 28, 336, 781, 763, 647, 467, 780, 348, 841, 836, 728, 528, 469, 504, 294, 281, 41, 814, 820, 866, 454, 545, 803, 733, 449, 448, 49, 595, 790, 279, 168, 203, 691, 518, 779, 131, 296, 570, 45, 512, 468, 538, 637, 357, 541, 457, 350, 917, 755, 929, 832, 455, 452, 904, 335, 274, 742, 166, 383, 401, 527, 791, 255, 39, 50, 927, 55, 583, 796, 52, 259, 451, 352, 801, 542, 519, 784, 663, 721, 159, 761, 280, 382, 400, 795, 134]. Science [164, 635, 258, 198, 430, 466, 778, 782, 426, 1062, 916, 76, 754, 802, 199, 793, 812, 544, 92, 62, 253, 620, 633, 847, 598, 931, 925, 135, 180, 712, 81, 65, 848, 531, 823, 993, 974, 1026, 971, 997, 56, 228, 367, 787]. Sciences [849, 676, 742, 530, 2, 85, 764, 910, 640, 331, 842, 463, 548, 65, 670, 873, 64]. Scientia [595]. Scientific [759, 783, 48, 4, 661, 515, 669, 706, 550, 349, 471, 954, 15, 866, 141, 142, 328, 779, 722, 46, 705, 538, 290, 530, 895, 93, 84, 930, 477, 508, 437, 546, 92, 567, 702, 200, 1032, 994, 1028, 971, 1023, 955, 1031, 1038, 1003]. Scientist [225, 481, 586, 474, 382, 400]. Scientists [550, 479, 996, 1014]. Scientometric [94]. Scotland [648]. Scott [921]. Screening [839, 553, 310]. Scripture [547]. Scroll [908]. Sea [599, 605, 908]. Sea-Something [599]. Search [661, 216, 599]. Seasons [290]. Second [223, 29, 697, 857, 1063]. Second-Order [223, 29]. Section [989]. Sections [464]. Sector [166, 165]. Secular [449, 450]. Secularization [52, 451, 445]. Seductive [449]. Seeds [454]. seem [477]. Seismology [731, 729]. Selection [568]. Self [744, 104, 988, 1050]. self-fashioning [988]. self-inscription [1050]. Self-Projection [744]. Self-Regard [104]. Selfhood [205]. Semantics [356]. Semiotic [722]. Sense [739, 951]. Senses [762, 797]. Sensibility [627, 174]. Sensitive [178]. sensorium [1005]. Senufo [199]. Separating [332]. Separation [759, 811]. Serial [896]. Seriously [926]. Serological [38]. sertões [531]. Serum [612]. Serve [288]. Service [428]. Services [315]. Set [499, 981]. Setting [551]. seventeenth [964, 976, 990, 1010, 4, 679, 681, 233, 61, 597, 216, 76, 47]. Seventeenth- [216]. seventeenth-century [964, 976, 990, 1010, 4, 61, 597, 76, 47]. Sexological [87]. Sexual [256, 321]. Sexuality [87]. Shape [349, 770, 396, 414]. Shaped [194]. Shaping [420, 942, 675, 859]. shared [994, 1028]. Sharing [930, 567, 994, 1028]. Sherrington [921, 437]. Shifting [930, 823]. Short [893, 485]. Showing [641]. Shrew [155]. Shtetl [587]. Shuang [664, 526]. Sibyls [596]. SIC [9, 10, 20, 21, 31, 32, 42, 43, 67, 68, 77, 78, 90, 91, 98, 99, 110, 111, 128, 129, 138, 139, 149, 150, 157, 158, 169, 170, 185, 186, 195, 196, 201, 202, 209, 210, 217, 218, 230, 231, 238, 239, 247, 248, 260, 261, 267, 268, 275, 276, 285, 286, 292, 293, 302, 303, 325, 326, 337, 338, 346, 347, 353, 354, 365, 366, 375, 376, 380, 381, 398, 399, 645, 646, 657, 658, 665, 666, 672, 673, 677, 678, 684, 685]. SIC [694, 695, 700, 701, 710, 711, 718, 719, 726, 727, 734, 735, 740, 741, 750, 751, 756, 757, 767, 768, 776, 777, 785, 786, 798, 799, 806, 807, 816, 817, 825, 826, 834, 835, 844, 845, 853, 854, 862, 863, 870, 871, 877, 878, 883, 884, 889, 890, 900, 901, 906, 907, 911, 912, 918, 919, 923, 924, 932, 933, 937, 938, 945, 946, 952, 953, 960, 961, 969, 970, 977, 978, 985, 986, 991, 992, 998, 999, 1006, 1007, 1017, 1018, 1029, 1030]. SIC [1039, 1040, 1046, 1047, 1055, 1056]. sick [956]. Sight [430]. Sigmund [103, 689]. Significance [358, 927, 137]. significant [1004]. Simas [802]. Similes [176]. Simplicius [486]. Simulation [745, 828, 342, 341]. Simulations [1045, 339, 643, 644]. Simultaneous [609]. since [821]. Sincience [165]. Singularity [650]. Site [73]. Siting [820]. Situated [821]. Situating [879]. Situation [801]. Siwalik [537]. Six [981, 420]. Sixteenth [280, 278]. Sixteenth-Century [278]. Size [535, 534]. sketch [983]. Skill [965]. Skills [707]. Skin [859]. Skulls [479]. Slavery [529]. Sloan [739]. Smallpox [903]. Smiles [477]. Smoluchowski [1025]. Social [1045, 105, 670, 586, 94, 676, 7, 656, 602, 720, 539, 37, 173, 847, 324, 180, 587, 747, 317, 848, 994, 1028, 1053, 995, 761]. Social-Economic [587]. Social-Epistemic [847]. Socialistic [848]. Societal [698]. Society [676, 934, 179, 897, 280]. Sociocultural [197]. Sociology [6, 290, 92, 949, 974]. Sofia [866]. Solitude [76]. Solvay [116]. Some [725, 38, 872, 319, 693, 689, 524, 660, 2]. Something [599]. sondern [1057]. Sooty [681]. Sophia [636]. Sorokin [56, 57]. Soul [4, 107]. Souls [3]. Sources [14, 358, 818]. Sovereignty [914]. Soviet [783, 504, 458, 457, 723, 996, 585]. Space [72, 723, 71, 134, 502, 535, 534]. Spain [760, 270]. Spanish [804]. Spatial [602]. Special [524]. Species [753, 245, 246, 456, 775]. Specific [38]. Specimen [95]. specimens [1036]. Spectacle [926, 754]. Spectacles [704]. Spectatorship [744]. Speculation [84]. Speechlessness [649]. Speed [857]. spetsyele [589]. Sphere [783, 760, 198, 619]. Spheres [332]. Sphygmograph [894, 891]. Spinoza [864]. Spirits [3]. Spirituality [46]. Spot [564]. Spots [385, 403]. Spreading [641]. St. [119]. Stability [92]. Stabilization [614]. Stabilizing [616, 676]. Stable [819, 613]. Stahl [626, 244]. Stake [262]. Stalin [457, 587]. Stance [284]. Standardization [858]. Standardizing [617]. Standards [707, 810]. Starke [840]. Start [356]. State [783, 824, 808, 579, 324, 957, 689, 566]. state-building [957]. States [189, 459, 462, 842, 581]. Station [866]. Stations [564]. Statistical [709, 574, 575]. Statisticians [162]. Statistics [637, 243, 536, 955, 675]. Status [706, 681]. Steeplechase [796]. Stefano [1012]. Stein [749]. Steps [252]. Stimulations [857]. Stimulus [626]. Stimulus-Reaction-Schemes [626]. Stones [641]. Stories [829, 967]. Storm [654]. Story [166, 560]. Storytelling [929]. Strange [850]. Strategies [266, 859]. Strategy [880]. Strength [746]. stress [975]. Strong [762, 797]. Structural [698, 357]. Structuralism [264]. Structure [70, 92]. Structures [502]. struggles [971]. Struggling [124]. Student [963]. Studies [528, 459, 872, 873, 1022]. Studiorum [464, 463]. Study [294, 604, 279, 38, 119, 637, 216, 801, 125, 585, 463, 641, 1010]. Style [927, 84, 80, 81, 644]. Styles [888, 88, 85, 299]. Subject [876]. Subjectivity [100]. Sublime [730]. Sublimity [649]. Submitted [693]. Substance [576]. Substances [237]. Succi [1012]. suicide [1004]. superiority [983]. Supply [197]. Surface [855]. Surfaces [856]. Surgical [841]. Surrealism [516]. Survey [69, 910, 316, 995]. Surveys [459, 462, 460, 1049]. Survival [184]. Susceptibility [810]. Suspicion [892, 548]. Swarming [714]. Sweden [707, 861]. Swedish [819, 950, 811]. Swiss [197, 653]. Switzerland [731, 651]. Symbol [362]. Symbolic [356, 846, 363, 360]. Symbolism [358]. Symposion [1062, 1063]. Symposium [1062, 1063]. syncretism [1016]. Synoptic [887]. Synthesis [544, 1044]. Synthetic [62]. System [626, 605, 395, 413]. Systems [144, 130].

Tūsī [422]. Table [237, 846, 236]. Tachistoscope [289]. Tactical [113]. Tactility [520]. Taking [926]. talents [1020]. Talking [193, 1015]. Tangible [716]. Task [430]. Tasks [592, 590]. Taste [539]. Taurellus [831]. Teaching [707, 851, 698]. Technics [939]. Technique [775, 1059]. Techniques [442, 769]. Technological [197, 71, 192, 130]. technologies [957, 1050]. Technology [634, 824, 808, 348, 841, 189, 518, 194, 939, 942, 190, 191, 193, 62, 192, 156, 188]. Telegraph [708]. teleology [1019]. Telescope [386, 404]. Tell [296, 724]. Temporal [869, 896]. Temporality [771]. ten [1020]. tenth [1025]. teorye [589]. Terms [491]. terrifying [1037]. Territories [653]. Territory [74]. Test [837, 430]. Testimony [328, 598, 486]. Testing [309, 810, 332]. Text [511, 592, 547, 39, 574, 575]. Textbooks [898]. Texts [418, 489, 427, 546]. textual [968]. Thabit [424, 664, 526]. Theater [225, 749, 680]. Their [887, 471, 489, 339, 852, 899, 830, 420]. Thematic [837]. Theology [33, 600, 597, 596, 278, 594]. Theoretical [774, 359]. Theories [626, 340, 106, 553, 487]. Theorizing [874]. Theory [1045, 576, 688, 299, 294, 345, 499, 122, 118, 119, 941, 762, 797, 638, 22, 573, 894, 562, 127, 908, 311, 579, 154, 510, 123, 287, 543, 766, 765, 632, 486, 1043, 623, 622, 589]. Theory-Driven [762, 797]. Theory-Ladenness [762, 797]. Therapeutic [611]. Therapy [178, 181, 689]. There [312, 2, 85]. Theses [2]. Thesis [46, 54]. Thin [859]. Thing [720]. Things [160, 607]. Thinking [229, 38, 749, 226, 450]. Third [262]. Third-Century [262]. Thirties [120]. Thomas [477]. Thomist [595]. Thompson [803]. Those [97]. Thought [320, 200, 1003]. Thoughts [524, 53]. Threat [634]. Three [459, 387, 405, 270]. Tier [229, 226]. Time [849, 433, 182, 1004, 259, 26]. Times [783]. Timiryazev [1026]. Tissue [716]. Titles [15]. Toeplitz [851]. Toledo [425]. Tomb [754]. Tongues [731]. Tool [650, 669]. Tools [106, 914]. Topical [71]. Topics [142]. Topinambes [492]. Topos [858]. Tortoise [441]. Touch [580]. Tov [281]. tracing [995]. Tractatus [270]. Trading [538]. Tradition [46, 429, 235, 50, 298, 973, 958]. Traditional [308]. Traditions [424, 87]. Trait [539]. Traits [306]. Trajectory [396, 414, 995]. Transcendence [599]. Transcendentalism [833]. Transference [683]. Transformation [417, 206, 881]. Transformative [950]. Transit [1035]. Transition [134, 997]. Translated [535]. Translation [417, 511, 592, 425, 743]. Translations [271, 270]. Translators [271]. Transmission [417, 419, 350, 573, 427]. Transnational [814]. Transposing [46, 81]. Trauma [555, 975, 1004]. Traumatic [975, 443]. travel [967]. Treasures [917]. treatise [990]. Treatises [818]. Treatment [546]. Treatments [621]. Trials [648]. Tricks [683]. Tridentine [598]. Triple [494]. Trojan [606]. Tropes [142]. True [1050]. Trust [349]. trustworthy [1009]. Truth [254, 263, 291, 172]. Tshuve} [956]. Tube [838]. Tuberculosis [617]. Tubingen [1063, 1062, 1057]. Tübinger [1057, 1063]. Turbulence [944]. Turn [917, 451, 566, 192, 1021]. Turn-of-the-Century [566, 192]. Tuvia [589]. Twelfth [425]. Twentieth [819, 836, 552, 698, 436, 512, 917, 715, 259, 451, 579, 778, 782, 675]. Twenty [621]. Two [674, 229, 364, 465, 299, 226, 191, 754, 729, 829, 630, 363]. Two-Tier [229, 226]. Tycho [759]. Typical [461]. Typography [1003]. Typologies [736]. typology [1003].

U.S. [675]. Ukrainian [958]. Ultra [178]. Ultra-Orthodox [178]. Unbound [480]. Uncertain [822]. Uncertainty [805, 154]. Unconscious [837]. undead [1034]. undercounting [1051]. Underpinnings [656]. Understanding [60, 931, 925]. Unending [796]. UNESCO [761]. Unexpected [652]. Unfallneigung [609]. Unfolding [342]. Unhealthy [564]. Union [585]. United [811, 189, 459, 462, 842, 581]. Unity [508]. Universal [519]. Universe [662]. Universität [1062]. Universities [331, 463]. University [503, 605, 1062, 852, 899, 848, 688, 963, 848]. Unpacking [905]. Unpublished [623, 622, 689, 363]. UNRRA [568]. untidy [1034]. Untold [560]. Unusual [272]. urchins [605]. Uruguay [997]. Use [774, 488, 1061, 943, 829, 974]. Useful [1012]. Uses [597, 975]. Using [1020, 307]. USSR [459]. Utopian [603].

v [332]. Vaccine [617, 699, 616]. Vaccine/Autism [699]. Vaccines [611, 614]. Vacuum [838]. Valéry’s [773]. Validation [653]. Value [887, 214]. Values [313]. Vannevar [577]. Variances [243]. Variation [905, 305]. Varieties [742]. Various [464]. Verbal [570]. vernacular [1002]. Version [419]. Versions [813]. Versus [255, 181, 301, 578, 762, 797, 200, 1001]. Vesuvius [1037]. Vicissitudes [689]. Vicq [640]. Victorian [162]. Victorians [33]. Vienna [501]. View [189, 791, 226, 173, 172]. Views [661, 714]. VII [488]. Violence [841]. Virgin [386, 404]. Virtual [680]. Virtually [828]. Visibility [370, 164]. Visible [70, 167, 1000]. Vision [515, 430]. Visual [770, 570, 968]. Visualizing [914]. Vitalism [631, 628, 624]. Vitalist [629, 630]. Vitalizing [699]. Voice [473, 584]. Void [272]. volcanoes [605]. Volga [955]. Volta [1034]. Volterra [868]. volume [9, 10, 20, 21, 31, 32, 42, 43, 58, 59, 67, 68, 77, 78, 90, 91, 98, 99, 110, 111, 128, 129, 138, 139, 149, 150, 157, 158, 169, 170, 185, 186, 195, 196, 201, 202, 209, 210, 217, 218, 230, 231, 238, 239, 247, 248, 260, 261, 267, 268, 275, 276, 285, 286, 292, 293, 302, 303, 325, 326, 337, 338, 346, 347, 353, 354, 365, 366, 375, 376, 380, 381, 398, 399, 645, 646, 657, 658, 665, 666, 672, 673, 677, 678]. volume [684, 685, 694, 695, 700, 701, 710, 711, 718, 719, 726, 727, 734, 735, 740, 741, 750, 751, 756, 757, 767, 768, 776, 777, 785, 786, 798, 799, 806, 807, 816, 817, 825, 826, 834, 835, 844, 845, 853, 854, 862, 863, 870, 871, 877, 878, 883, 884, 889, 890, 900, 901, 906, 907, 911, 912, 918, 919, 923, 924, 932, 933, 937, 938, 945, 946, 952, 953, 960, 961, 969, 970, 977, 978, 985, 986, 991, 992, 998, 999, 1006, 1007, 1017, 1018]. volume [1029, 1030, 1039, 1040, 1046, 1047, 1055, 1056]. Voluntarism [948]. Vorträge [1062]. vs [892, 829]. Vu [545]. Vulnerable [859, 1022].

Walter [441]. Wants [207]. War [266, 505, 240, 455, 850, 804, 637, 804, 996]. Warburg [362]. Warhol [73]. Warmian [880]. Warming [763]. Warren [440]. Wars [284]. was [1035]. Watching [705, 704]. Wave [562]. Waves [897]. Way [38, 782, 913, 533]. Ways [483]. Weak [779, 762, 797, 746]. Weather [887, 708]. Weimar [439, 515, 1052]. Weismann [765]. Well [316, 607]. Wentworth [803]. Were [720, 801]. Wertbestimmung [615]. Wesen [434]. West [417, 573, 610]. Western [914, 350, 573]. Weyl [1021, 502]. Where [356]. which [1032, 607]. White [384, 402]. Who [207, 206, 949]. Wildlife [926]. Wilhelm [1057, 1062, 1063, 1059, 1058, 1063]. Will [145, 867]. William [1062, 1057, 103]. Wish [553]. Wissenschaftsgeschichte [1062]. Witche [116]. Withering [264]. within [235, 879]. Without [251, 832, 624]. Witness [728]. Wohlwill [395, 413]. Wolff [244, 543, 632]. Wolfhound [495]. Women [467, 469, 779, 468, 466]. wonder [1010]. wonder-working [1010]. Wonderful [916]. Wonders [372]. Words [722]. Work [291, 73, 530, 273, 639, 818, 921, 540, 827, 823]. Work-Models [827]. Worker [761]. Working [893, 477, 1010]. Working-Class [477]. Works [393, 411, 281, 15, 360, 989, 1002]. Workshop [229, 235]. World [6, 713, 621, 715, 693, 916, 172, 316, 1016, 450, 1052, 822, 266, 505, 240, 850, 1044, 804]. World-Picture [6]. Worlds [265]. Worldview [445]. worthy [607]. would [450, 477]. Write [790, 793]. Writer [773]. Writing [610, 497, 141, 769, 772, 773, 483, 512]. Writings [394, 412, 430]. Written [465]. Wronski [885]. Wrote [206]. Würzburg [374].

Yang [989]. Ye [818, 989]. Years [621, 804, 1062]. Yeda’aya [272]. Yehuda [758]. Yiddish [590, 588, 592, 589, 583, 956]. Yiddishist [586]. Yom [281]. youth [968]. Youthful [962].

Z [524]. Zhao [664, 526]. Zionism [322]. Zionist [688, 503]. Zone [538]. Zoo [705]. Zoological [866]. Zum [1063]. Zweites [1063].

References

Anonymous:1987:E

[1] Anonymous. Editorial. Science in Context, 1(1):3, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Markus:1987:WTN

[2] Gyorgy Markus. Why is there no hermeneutics of natural sciences? Some preliminary theses. Science in Context, 1(1):5–51, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schaffer:1987:GMM

[3] Simon Schaffer. Godly men and mechanical philosophers: Souls and spirits in restoration natural philosophy. Science in Context, 1(1):53–85, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bechler:1987:ESS

[4] Zev Bechler. The essence and soul of Seventeenth-Century scientific revolution. Science in Context, 1(1):87–101, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Freudenthal:1987:IN

[5] Gideon Freudenthal. Introductory note. Science in Context, 1(1):103– 108, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Borkenau:1987:SMW

[6] Franz Borkenau. The sociology of the mechanistic world-picture. Science in Context, 1(1):109–127, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Grossmann:1987:SFM

[7] Henryk Grossmann. The social foundations of mechanistic philosophy and manufacture. Science in Context, 1(1):129–180, March 1987. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Russo:1987:HGF

[8] Valeria E. Russo. Henryk Grossmann and Franz Borkenau: a biobibliography. Science in Context, 1(1):181–191, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1987:SVIa

[9] Anonymous. SIC volume 1 issue 1 cover and front matter. Science in Context, 1(1):f1–f4, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1987:SVIb

[10] Anonymous. SIC volume 1 issue 1 cover and back matter. Science in Context, 1(1):b1–b3, March 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:1987:KSJ

[11] Rivka Feldhay. Knowledge and salvation in Jesuit culture. Science in Context, 1(2):195–213, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wolff:1987:IMP

[12] Michael Wolff. Impetus mechanics as a physical argument for Copernicanism: Copernicus, Benedetti, Galileo. Science in Context, 1(2):215– 256, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sylla:1987:OCC

[13] Edith Sylla. The Oxford calculators in context. Science in Context, 1(2): 257–279, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hoyrup:1987:FIM

[14] Jens Høyrup. The formation of ‘Islamic mathematics’: Sources and conditions. Science in Context, 1(2):281–329, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Duhem:1987:AST

[15] Pierre Duhem. An account of the scientific titles and works of Pierre Duhem. Science in Context, 1(2):333–348, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schaffer:1987:IN

[16] Simon Schaffer. Introductory note. Science in Context, 1(2):351–352, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gurevich:1987:G

[17] A. J. Gurevich. On Gurevich. Science in Context, 1(2):353–356, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gurevich:1987:PD

[18] A. J. Gurevich. On Pierre Duhem. Science in Context, 1(2):357–361, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Livesey:1987:PD

[19] Steven J. Livesey. On Pierre Duhem. Science in Context, 1(2):363–370, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1987:SVIc

[20] Anonymous. SIC volume 1 issue 2 cover and front matter. Science in Context, 1(2):f1–f4, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1987:SVId

[21] Anonymous. SIC volume 1 issue 2 cover and back matter. Science in Context, 1(2):b1–b4, September 1987. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lenoir:1988:PRC

[22] Timothy Lenoir. Practice, reason, context: The dialogue between theory and experiment. Science in Context, 2(1):3–22, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Shapin:1988:RBM

[23] Steven Shapin. Robert Boyle and mathematics: Reality, representation, and experimental practice. Science in Context, 2(1):23–58, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hughes:1988:MBI

[24] Thomas P. Hughes. Model builders and instrument makers. Science in Context, 2(1):59–75, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wise:1988:MM

[25] M. Norton Wise. Mediating machines. Science in Context, 2(1):77–113, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schaffer:1988:AMT

[26] Simon Schaffer. Astronomers mark time: Discipline and the personal equation. Science in Context, 2(1):115–145, Spring 1988. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Beller:1988:EAO

[27] Mara Beller. Experimental accuracy, operationalism, and limits of knowledge — 1925 to 1935. Science in Context, 2(1):147–162, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-Menahem:1988:MSF

[28] Yemima Ben-Menahem. Models of science: Fictions or idealizations? Science in Context, 2(1):163–175, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Elkana:1988:ESO

[29] Yehuda Elkana. Experiment as a second-order concept. Science in Context, 2(1):177–196, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Galison:1988:HPC

[30] Peter Galison. History, philosophy, and the central metaphor. Science in Context, 2(1):197–212, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1988:SVIa

[31] Anonymous. SIC volume 2 issue 1 cover and front matter. Science in Context, 2(1):f1–f4, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1988:SVIb

[32] Anonymous. SIC volume 2 issue 1 cover and back matter. Science in Context, 2(1):b1–b3, Spring 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gascoigne:1988:BVR

[33] John Gascoigne. From Bentley to the Victorians: The rise and fall of British Newtonian natural theology. Science in Context, 2(2):219–256, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Yeo:1988:GMM

[34] Richard Yeo. Genius, method, and morality: Images of Newton in Britain, 1760–1860. Science in Context, 2(2):257–284, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vonBorzeszkowski:1988:EMF

[35] Horst-Heino von Borzeszkowski and Renate Wahsner. Evolutionism as a modern form of mechanicism. Science in Context, 2(2):287–306, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stengers:1988:EP

[36] Isabelle Stengers. Epistemology and the physicist. Science in Context, 2 (2):307–316, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Raj:1988:IKS

[37] Kapil Raj. Images of knowledge, social organization, and attitudes to research in an Indian physics department. Science in Context, 2(2): 317–339, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fleck:1988:SSF

[38] Ludwik Fleck. Some specific features of the serological way of thinking: a methodological study. Science in Context, 2(2):343–344, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lowy:1988:QSC

[39] Ilana Löwy. Quantification in science and cognition circa 1937: a newly discovered text of Ludwik Fleck. Science in Context, 2(2):345–355, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wolters:1988:HD

[40] Gereon Wolters. Hugo Dingler. Science in Context, 2(2):359–367, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dingler:1988:MIE

[41] Hugo Dingler. Method instead of epistemology and philosophy of science. Science in Context, 2(2):369–408, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1988:SVIc

[42] Anonymous. SIC volume 2 issue 2 cover and front matter. Science in Context, 2(2):f1–f5, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1988:SVId

[43] Anonymous. SIC volume 2 issue 2 cover and back matter. Science in Context, 2(2):b1–b6, Autumn 1988. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:1989:EI

[44] Rivka Feldhay and Yehuda Elkana. Editors’ introduction. Science in Context, 3(1):3–8, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heilbron:1989:SC

[45] J. L. Heilbron. Science in the Church. Science in Context, 3(1):9–28, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Harris:1989:TMT

[46] Steven J. Harris. Transposing the Merton thesis: Apostolic spirituality and the establishment of the Jesuit scientific tradition. Science in Context, 3(1):29–65, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wallace:1989:PAP

[47] William A. Wallace. The problem of apodictic proof in early Seventeenth-Century mechanics. Galileo, Guevara, and the Jesuits. Science in Context, 3(1):67–87, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ashworth:1989:LRL

[48] William B. Ashworth. Light of reason, light of nature. Catholic and Protestant metaphors of scientific knowledge. Science in Context, 3(1): 89–107, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:1989:DPS

[49] Rivka Feldhay and Michael Heyd. The discourse of pious science. Science in Context, 3(1):109–142, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mali:1989:STS

[50] Joseph Mali. Science, tradition, and the science of tradition. Science in Context, 3(1):143–173, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hatfield:1989:RNG

[51] Gary Hatfield. Reason, nature, and God in Descartes. Science in Context, 3(1):175–201, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Motzkin:1989:CRS

[52] Gabriel Motzkin. The Catholic response to secularization and the rise of the history of science as a discipline. Science in Context, 3(1):203–226, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Struik:1989:FTM

[53] Dirk Struik. Further thoughts on Merton in context. Science in Context, 3(1):227–238, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zuckerman:1989:OMT

[54] Harriet Zuckerman. The other Merton thesis. Science in Context, 3 (1):239–267, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mendelsohn:1989:RKM

[55] Everett Mendelsohn. Robert K. Merton: The celebration and defense of science. Science in Context, 3(1):269–289, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Merton:1989:SMC

[56] Robert K. Merton. The Sorokin–Merton correspondence on ‘Puritanism, Pietism and Science,’ 1933–34. Science in Context, 3(1):291–298, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sorokin:1989:S

[57] P. A. Sorokin. On Sorokin. Science in Context, 3(1):299–302, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1989:SCVa

[58] Anonymous. Science in context volume 3 issue 1 cover and front matter. Science in Context, 3(1):f1–f4, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1989:SCVb

[59] Anonymous. Science in context volume 3 issue 1 cover and back matter. Science in Context, 3(1):b1–b6, Spring 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dennis:1989:GUI

[60] Michael Aaron Dennis. Graphic understanding: Instruments and interpretation in Robert Hooke’s Micrographia. Science in Context, 3(2):

309–364, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fouke:1989:MOM

[61] Daniel C. Fouke. Mechanical and ‘organical’ models in Seventeenth-Century explanations of biological reproduction. Science in Context, 3 (2):365–381, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Travis:1989:SRT

[62] Anthony S. Travis. Science as receptor of technology: Paul Ehrlich and the synthetic dyestuffs industry. Science in Context, 3(2):383–408, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:1989:LRG

[63] Leo Corry. Linearity and reflexivity in the growth of mathematical knowledge. Science in Context, 3(2):409–440, Autumn 1989. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sivin:1989:CAS

[64] Nathan Sivin. On the Chinese Academy of Sciences. Science in Context, 3 (2):443–445, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Yao:1989:CIS

[65] Shuping Yao. Chinese intellectuals and science: a history of the Chinese Academy of Sciences (CAS). Science in Context, 3(2):447–473, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heelan:1989:CC

[66] Patrick Heelan. Comments and critique. Science in Context, 3(2):477– 488, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1989:SVIa

[67] Anonymous. SIC volume 3 issue 2 cover and front matter. Science in Context, 3(2):f1–f4, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1989:SVIb

[68] Anonymous. SIC volume 3 issue 2 cover and back matter. Science in Context, 3(2):b1–b6, Autumn 1989. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ophir:1991:PKM

[69] Adi Ophir and Steven Shapin. The place of knowledge: a methodological survey. Science in Context, 4(1):3–22, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hillier:1991:VCS

[70] Bill Hillier and Alan Penn. Visible colleges: Structure and randomness in the place of discovery. Science in Context, 4(1):23–50, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lynch:1991:LST

[71] Michael Lynch. Laboratory space and the technological complex: An investigation of topical contextures. Science in Context, 4(1):51–78, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kelly:1991:SEP

[72] Michael Kelly and Ricardo Sanchez. The space of the ethical practice of emergency medicine. Science in Context, 4(1):79–100, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jones:1991:AWF

[73] Caroline A. Jones. Andy Warhol’s ‘Factory’: The production site, its context and its impact on the work of art. Science in Context, 4(1):101– 132, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Revel:1991:KT

[74] Jacques Revel. Knowledge of the territory. Science in Context, 4(1):133– 162, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ophir:1991:PKR

[75] Adi Ophir. A place of knowledge re-created: The library of Michel de Montaigne. Science in Context, 4(1):163–190, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Shapin:1991:MOP

[76] Steven Shapin. ‘The mind is its own place’: Science and solitude in Seventeenth-Century England. Science in Context, 4(1):191–218, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1991:SVIa

[77] Anonymous. SIC volume 4 issue 1 cover and front matter. Science in Context, 4(1):f1–f4, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1991:SVIb

[78] Anonymous. SIC volume 4 issue 1 cover and back matter. Science in Context, 4(1):b1–b3, Spring 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Daston:1991:I

[79] Lorraine Daston and Michael Otte. Introduction. Science in Context, 4 (2):223–232, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Otte:1991:SHC

[80] Michael Otte. Style as a historical category. Science in Context, 4(2): 233–264, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wessely:1991:TSH

[81] Anna Wessely. Transposing ‘style’ from the history of art to the history of science. Science in Context, 4(2):265–278, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jahnke:1991:MCC

[82] Hans Niels Jahnke. Mathematics and culture: The case of Novalis. Science in Context, 4(2):279–295, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Richards:1991:RCF

[83] Joan L. Richards. Rigor and clarity: Foundations of mathematics in France and England, 1800–1840. Science in Context, 4(2):297–319, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Myers:1991:SSL

[84] Greg Myers. Scientific speculation and literary style in a molecular genetics article. Science in Context, 4(2):321–346, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Reingold:1991:PAT

[85] Nathan Reingold. The peculiarities of the Americans or are there national styles in the sciences? Science in Context, 4(2):347–366, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Daston:1991:IRR

[86] Lorraine Daston. The ideal and reality of the Republic of letters in the enlightenment. Science in Context, 4(2):367–386, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Nye:1991:HSC

[87] Robert A. Nye. The history of sexuality in context: National sexological traditions. Science in Context, 4(2):387–406, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Maienschein:1991:ESG

[88] Jane Maienschein. Epistemic styles in German and American embryology. Science in Context, 4(2):407–427, Autumn 1991. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Harrington:1991:IGP

[89] Anne Harrington. Interwar ‘German’ psychobiology: Between nationalism and the irrational. Science in Context, 4(2):429–447, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1991:SVIc

[90] Anonymous. SIC volume 4 issue 2 cover and front matter. Science in Context, 4(2):f1–f4, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1991:SVId

[91] Anonymous. SIC volume 4 issue 2 cover and back matter. Science in Context, 4(2):b1–b7, Autumn 1991. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stichweh:1992:SSD

[92] Rudolf Stichweh. The sociology of scientific disciplines: On the Genesis and stability of the disciplinary structure of modern science. Science in Context, 5(1):3–15, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

McAllister:1992:CAS

[93] James W. McAllister. Competition among scientific disciplines in cold nuclear fusion research. Science in Context, 5(1):17–49, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fischer:1992:SCD

[94] Klaus Fischer. The social and cognitive dynamics of paradigmatic change: a scientometric approach. Science in Context, 5(1):51–96, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Clark:1992:ISD

[95] William Clark. On the ironic specimen of the Doctor of Philosophy. Science in Context, 5(1):97–137, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ludwig:1992:WNL

[96] Bernd Ludwig. What is Newton’s Law of Inertia about? Philosophical reasoning and explanation in Newton’s Principia. Science in Context, 5 (1):139–163, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mirowski:1992:LNN

[97] Philip Mirowski. Looking for those natural numbers: Dimensionless constants and the idea of natural measurement. Science in Context, 5(1): 165–188, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1992:SVIa

[98] Anonymous. SIC volume 5 issue 1 cover and front matter. Science in Context, 5(1):f1–f4, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1992:SVIb

[99] Anonymous. SIC volume 5 issue 1 cover and back matter. Science in Context, 5(1):b1–b11, Spring 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ash:1992:HMS

[100] Mitchell G. Ash. Historicizing mind science: Discourse, practice, subjectivity. Science in Context, 5(2):193–207, Autumn 1992. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Daston:1992:NFI

[101] Lorraine Daston. The naturalized female intellect. Science in Context, 5 (2):209–235, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Smith:1992:MID

[102] Roger Smith. The meaning of ‘inhibition’ and the discourse of order. Science in Context, 5(2):237–263, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Joravsky:1992:IPI

[103] David Joravsky. The impossible project of Ivan Pavlov (and William James and Sigmund Freud). Science in Context, 5(2):265–280, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Morawski:1992:SRO

[104] Jill G. Morawski. Self-regard and other-regard: Reflexive practices in American psychology, 1890–1940. Science in Context, 5(2):281–308, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Danzier:1992:PES

[105] Kurt Danzier. The project of an experimental social psychology: Historical perspectives. Science in Context, 5(2):309–328, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gigerenzer:1992:DCP

[106] Gerd Gigerenzer. Discovery in cognitive psychology: New tools inspire new theories. Science in Context, 5(2):329–350, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rose:1992:EHS

[107] Nikolas Rose. Engineering the human soul: Analyzing psychological expertise. Science in Context, 5(2):351–369, Autumn 1992. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:1992:KLP

[108] Alexandre Métraux. Kurt Lewin: Philosopher-psychologist. Science in Context, 5(2):372–384, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lewin:1992:LEP

[109] Kurt Lewin. Law and experiment in psychology. Science in Context, 5 (2):385–416, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1992:SVIc

[110] Anonymous. SIC volume 5 issue 2 cover and front matter. Science in Context, 5(2):f1–f4, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1992:SVId

[111] Anonymous. SIC volume 5 issue 2 cover and back matter. Science in Context, 5(2):b1–b4, Autumn 1992. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:1993:EI

[112] Jürgen Renn and Robert S. Cohen. Editors’ introduction. Science in Context, 6(1):3–13, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schulmann:1993:EPO

[113] Robert Schulmann. Einstein at the Patent Office: Exile, salvation, or tactical retreat? Science in Context, 6(1):17–24, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hughes:1993:EII

[114] Thomas P. Hughes. Einstein, inventors, and invention. Science in Context, 6(1):25–42, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kox:1993:ELM

[115] Anne J. Kox. Einstein and Lorentz: More than just good colleagues. Science in Context, 6(1):43–56, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Barkan:1993:WSF

[116] Diana Kormos Barkan. The witches’ sabbath: The First International Solvay Congress in Physics. Science in Context, 6(1):59–82, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Eisenstaedt:1993:DBB

[117] Jean Eisenstaedt. Dark bodies and black holes, magic circles and Montgolfiers: Light and gravitation from Newton to Einstein. Science in Context, 6(1):83–106, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goenner:1993:RRT

[118] Hubert Goenner. The reaction to Relativity theory I: The anti-Einstein campaign in Germany in 1920. Science in Context, 6(1):107–133, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hentschel:1993:CSJ

[119] Klaus Hentschel. The conversion of St. John: a case study on the interplay of theory and experiment. Science in Context, 6(1):137–194, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stuewer:1993:MEN

[120] Roger H. Stuewer. Mass-energy and the neutron in the Early Thirties. Science in Context, 6(1):195–238, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Beller:1993:EBR

[121] Mara Beller. Einstein and Bohr’s rhetoric of complementarity. Science in Context, 6(1):241–255, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fine:1993:EIQ

[122] Arthur Fine. Einstein’s interpretations of the quantum theory. Science in Context, 6(1):257–273, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stachel:1993:OEE

[123] John Stachel. The other Einstein: Einstein contra field theory. Science in Context, 6(1):275–290, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-Menahem:1993:SCE

[124] Yemima Ben-Menahem. Struggling with causality: Einstein’s case. Science in Context, 6(1):291–310, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:1993:EDG

[125] Jürgen Renn. Einstein as a disciple of Galileo: a comparative study of concept development in physics. Science in Context, 6(1):311–341, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Disalle:1993:CGN

[126] Robert Disalle. Carl Gottfried Neumann. Science in Context, 6(1):345– 353, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Neumann:1993:PGN

[127] Carl Neumann. On the principles of the Galilean–Newtonian theory. Science in Context, 6(1):355–368, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1993:SVIa

[128] Anonymous. SIC volume 6 issue 1 cover and front matter. Science in Context, 6(1):f1–f4, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1993:SVIb

[129] Anonymous. SIC volume 6 issue 1 cover and back matter. Science in Context, 6(1):b1–b6, Spring 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Williams:1993:COE

[130] Rosalind Williams. Cultural origins and environmental implications of large technological systems. Science in Context, 6(2):377–403, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Grattan-Guinness:1993:ISN

[131] I. Grattan-Guinness. The ingénieur savant, 1800–1830: a neglected figure in the history of French mathematics and science. Science in Context, 6 (2):405–433, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

ONeill:1993:IRN

[132] John O’Neill. Intertextual reference in Nineteenth-Century mathematics. Science in Context, 6(2):435–468, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Israel:1993:EBC

[133] Giorgio Israel. The emergence of biomathematics and the case of population dynamics: a revival of mechanical reductionism and Darwinism. Science in Context, 6(2):469–509, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Roberts:1993:FSP

[134] Lissa Roberts. Filling the space of possibilities: Eighteenth-Century Chemistry’s transition from art to science. Science in Context, 6(2): 511–553, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weingart:1993:SAC

[135] Peter Weingart. Science abused? Challenging a legend. Science in Context, 6(2):555–567, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Baltas:1993:PMP

[136] Aristides Baltas. Physics as a mode of production. Science in Context, 6 (2):569–616, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schaller:1993:PIC

[137] Klaus Schaller. Patočka’s interpretation of Comenius and its significance for present-day pedagogics. Science in Context, 6(2):617–631, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1993:SVIc

[138] Anonymous. SIC volume 6 issue 2 cover and front matter. Science in Context, 6(2):f1–f4, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1993:SVId

[139] Anonymous. SIC volume 6 issue 2 cover and back matter. Science in Context, 6(2):b1–b12, Autumn 1993. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

M:1994:I

[140] J. M. Introduction. Science in Context, 7(1):3–6, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:1994:NCH

[141] Rivka Feldhay. Narrative constraints on historical writing: The case of the scientific revolution. Science in Context, 7(1):7–24, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gal:1994:TTS

[142] Ofer Gal. Tropes and topics in scientific discourse: Galileo’s De Motu. Science in Context, 7(1):25–52, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Krips:1994:IRB

[143] Henry Krips. Ideology, rhetoric, and Boyle’s New Experiments. Science in Context, 7(1):53–64, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Reinberger:1994:ESH

[144] Hans-Jörg Reinberger. Experimental systems: Historiality, narration, and deconstruction. Science in Context, 7(1):65–81, Spring 1994. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brunner:1994:EPW

[145] José Brunner. ‘Every path will end in darkness’ or: Why psychoanalysis needs metapsychology. Science in Context, 7(1):83–101, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Motzkin:1994:MMH

[146] Gabriel Motzkin. Memoirs, memory, and historical experience. Science in Context, 7(1):103–119, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mali:1994:NMH

[147] Joseph Mali. Narrative, myth, and history. Science in Context, 7(1):121– 142, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Erdibast-Vulcan:1994:NMC

[148] Daphna Erdibast-Vulcan. Narrative, modernism, and the crisis of authority: a Bakhtinian perspective. Science in Context, 7(1):143–158, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVIa

[149] Anonymous. SIC volume 7 issue 1 cover and front matter. Science in Context, 7(1):f1–f4, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVIb

[150] Anonymous. SIC volume 7 issue 1 cover and back matter. Science in Context, 7(1):b1–b5, Spring 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Klein:1994:OCC

[151] Ursula Klein. Origin of the concept chemical compound. Science in Context, 7(2):163–204, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Licoppe:1994:CNN

[152] Christian Licoppe. The crystallization of a new narrative form in experimental reports (1660–1690). Science in Context, 7(2):205–244, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weintraub:1994:PAB

[153] E. Roy Weintraub and Philip Mirowski. The pure and the applied: Bourbakism comes to mathematical economics. Science in Context, 7(2): 245–272, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Shenhav:1994:MUU

[154] Yehouda Shenhav. Manufacturing uncertainty and uncertainty in manufacturing: Managerial discourse and the rhetoric of organizational theory. Science in Context, 7(2):275–305, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Portugali:1994:SE

[155] Juval Portugali. The shrew environment. Science in Context, 7(2):307– 326, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

VanCreveld:1994:RFM

[156] Martin Van Creveld. The rise and fall of military technology. Science in Context, 7(2):327–351, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVIc

[157] Anonymous. SIC volume 7 issue 2 cover and front matter. Science in Context, 7(2):f1–f4, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVId

[158] Anonymous. SIC volume 7 issue 2 cover and back matter. Science in Context, 7(2):b1–b7, Summer 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Power:1994:SAF

[159] Michael Power. From the science of accounts to the financial accountability of science. Science in Context, 7(3):355–387, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Porter:1994:MTQ

[160] Theodore M. Porter. Making things quantitative. Science in Context, 7 (3):389–407, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jackson:1994:NAB

[161] Myles W. Jackson. Natural and artificial budgets: Accounting for Goethe’s economy of nature. Science in Context, 7(3):409–431, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Alborn:1994:CPV

[162] Timothy L. Alborn. A calculating profession: Victorian actuaries among the statisticians. Science in Context, 7(3):433–468, Autumn 1994. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Miller:1994:FL

[163] Peter Miller and Ted O’Leary. The factory as laboratory. Science in Context, 7(3):469–496, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Robson:1994:CSE

[164] Keith Robson. Connecting science to the economic: Accounting calculation and the visibility of research and development. Science in Context, 7 (3):497–514, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sherman:1994:GSP

[165] Brad Sherman. Governing sincience: Patents and public sector research. Science in Context, 7(3):515–538, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Law:1994:CCS

[166] John Law and Madeleine Akrich. On customers and costs: a story from public sector science. Science in Context, 7(3):539–561, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mirowski:1994:VHM

[167] Philip Mirowski. A visible hand in the marketplace of ideas: Precision measurement as arbitage. Science in Context, 7(3):563–589, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fuller:1994:TPS

[168] Steve Fuller. Toward a philosophy of science accounting: a critical rendering of instrumental rationality. Science in Context, 7(3):591–621, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVIe

[169] Anonymous. SIC volume 7 issue 3 cover and front matter. Science in Context, 7(3):f1–f4, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1994:SVIf

[170] Anonymous. SIC volume 7 issue 3 cover and back matter. Science in Context, 7(3):b1–b6, Autumn 1994. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heyd:1995:EI

[171] Michael Heyd and Hans-Jörg Rheinberger. Editors’ introduction. Science in Context, 8(1):3–8, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Unschuld:1995:PTE

[172] Paul U. Unschuld. Plausibility or truth? An essay on medicine and world view. Science in Context, 8(1):9–30, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schiefenhovel:1995:PES

[173] Wulf Schiefenhövel. Perception, expression, and social function of pain: a human ethological view. Science in Context, 8(1):31–46, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cohen:1995:THE

[174] Esther Cohen. Towards a history of European physical sensibility: Pain in the later middle ages. Science in Context, 8(1):47–74, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Garcia-Ballester:1995:CNF

[175] Luis García-Ballester. The construction of a new form of learning and practing medicine in Medieval Latin Europe. Science in Context, 8(1): 75–102, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ziegler:1995:MSR

[176] Joseph Ziegler. Medical similes in religious discourse: The case of Giovanni di San Gimignano OP (ca. 1260–ca. 1333). Science in Context, 8 (1):103–131, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heyd:1995:MDR

[177] Michael Heyd. Medical discourse in religious controversy: The case of the critique of ‘enthusiasm’ on the eve of the Enlightenment. Science in Context, 8(1):133–157, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bilu:1995:BSM

[178] Yoram Bilu and Eliezer Witztum. Between sacred and medical realities: Culturally sensitive therapy with Jewish Ultra-Orthodox patients. Science in Context, 8(1):159–173, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kamrat-Lang:1995:HSM

[179] Debora Kamrat-Lang. Healing society: Medical language in American eugenics. Science in Context, 8(1):175–196, Spring 1995. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weingart:1995:EMS

[180] Peter Weingart. Eugenics — medical or social science? Science in Context, 8(1):197–207, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lowy:1995:NMD

[181] Ilana Löwy. ‘Nothing more to be done’: Palliative care versus exerimental therapy in advanced cancer. Science in Context, 8(1):209–229, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jacoby:1995:TMG

[182] Miriam Jacoby. Time, magic, and gynecology contemporary Israeli practice. Science in Context, 8(1):231–248, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rheinberger:1995:BNC

[183] Hans-Jörg Rheinberger. Beyond nature and culture: a note on medicine in the age of molecular biology. Science in Context, 8(1):249–263, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Atlan:1995:BMS

[184] Henri Atlan. Biological medicine and the survival of the person. Science in Context, 8(1):265–277, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIa

[185] Anonymous. SIC volume 8 issue 1 cover and front matter. Science in Context, 8(1):f1–f4, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIb

[186] Anonymous. SIC volume 8 issue 1 cover and back matter. Science in Context, 8(1):b1–b5, Spring 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rieber:1995:I

[187] Alfred J. Rieber and Marsha Siefert. Introduction. Science in Context, 8 (2):281–292, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wallace:1995:TCM

[188] Anthony F. C. Wallace. Technology in culture: The meaning of cultural fit. Science in Context, 8(2):293–324, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cochran:1995:CTA

[189] Thomas C. Cochran. The culture of technology: An alternative view of the Industrial Revolution in the United States. Science in Context, 8(2): 325–339, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rieber:1995:PTE

[190] Alfred J. Rieber. Politics and technology in Eighteenth-Century Russia. Science in Context, 8(2):341–368, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Scranton:1995:PPT

[191] Philip Scranton. The politices of production: Technology, markets, and the two cultures of American industry. Science in Context, 8(2):369–395, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Trommler:1995:AGT

[192] Frank Trommler. The avant-garde and technology: Toward technological fundamentalism in turn-of-the-century Europe. Science in Context, 8(2): 397–416, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Siefert:1995:ATC

[193] Marsha Siefert. Aesthetics, technology, and the capitalization of culture: How the talking machine became a musical instrument. Science in Context, 8(2):417–449, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hughes:1995:STA

[194] Thomas Hughes. Shaped technology: An afterword. Science in Context, 8(2):451–455, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIc

[195] Anonymous. SIC volume 8 issue 2 cover and front matter. Science in Context, 8(2):f1–f4, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVId

[196] Anonymous. SIC volume 8 issue 2 cover and back matter. Science in Context, 8(2):b1–b4, Summer 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gugerli:1995:SAT

[197] David Gugerli. Sociocultural aspects of technological change: The rise of the Swiss electricity supply economy. Science in Context, 8(3):459–486, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rupp:1995:NSP

[198] Jan C. C. Rupp. The new science and the public sphere in the premodern era. Science in Context, 8(3):487–507, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sindzingre:1995:AMC

[199] Nicole Alice Sindzingre. The anthropology of misfortune and cognitive science. Examples from the Ivory Coast Senufo. Science in Context, 8 (3):509–529, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zambelli:1995:AKV

[200] Paola Zambelli. Alexandre Koyré versus Lucien Lévy–Bruhl: From collective representations to paradigms of scientific thought. Science in Context, 8(3):531–555, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIe

[201] Anonymous. SIC volume 8 issue 3 cover and front matter. Science in Context, 8(3):f1–f4, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIf

[202] Anonymous. SIC volume 8 issue 3 cover and back matter. Science in Context, 8(3):b1–b3, Autumn 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gilbert:1995:IPS

[203] Scott F. Gilbert. Introduction postmodernism and science. Science in Context, 8(4):559–561, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gilbert:1995:RBP

[204] Scott F. Gilbert. Resurrecting the body: Has postmodernism had any effect on biology? Science in Context, 8(4):563–577, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Tauber:1995:PIS

[205] Alfred I. Tauber. Postmodernism and immune selfhood. Science in Context, 8(4):579–607, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kay:1995:WWB

[206] Lily E. Kay. Who wrote the book of life? Information and the transformation of molecular biology, 1945–55. Science in Context, 8(4):609–634, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carson:1995:WWP

[207] Cathryn Carson. Who wants a postmodern physics? Science in Context, 8(4):635–655, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:CAI

[208] Anonymous. Cumulative author index. Science in Context, 8(4):657–664, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIg

[209] Anonymous. SIC volume 8 issue 4 cover and front matter. Science in Context, 8(4):f1–f3, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1995:SVIh

[210] Anonymous. SIC volume 8 issue 4 cover and back matter. Science in Context, 8(4):b1–b5, Winter 1995. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:F

[211] Anonymous. Foreword. Science in Context, 9(1):3, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dagognet:1996:REP

[212] François Dagognet. A regional epistemology with possibilities for expansion. Science in Context, 9(1):5–14, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:FDB

[213] Anonymous. François Dagognet: Bio-bibliography. Science in Context, 9(1):15–16, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Guery:1996:MVI

[214] François Guéry. The meaning and value of invention. Science in Context, 9(1):17–38, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Parrochia:1996:ALC

[215] Daniel Parrochia. Algorithmics and the limits of complexity. Science in Context, 9(1):39–56, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:1996:IES

[216] Alexandre Métraux. Impure epistemology and the search for the nervous agent: a case study in Seventeenth- and Eighteenth-Century neurophysics. Science in Context, 9(1):57–78, Spring 1996. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIa

[217] Anonymous. SIC volume 9 issue 1 cover and front matter. Science in Context, 9(1):f1–f4, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIb

[218] Anonymous. SIC volume 9 issue 1 cover and back matter. Science in Context, 9(1):b1–b4, Spring 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:EF

[219] Anonymous. Editors’ foreword. Science in Context, 9(2):83–84, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Nowotny:1996:I

[220] Helga Nowotny. Introduction. Science in Context, 9(2):85–91, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hiebert:1996:DIC

[221] Erwin N. Hiebert. Discipline identification in chemistry and physics. Science in Context, 9(2):93–119, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Tanay:1996:IMB

[222] Dorit Tanay. The image of music and the bodies of knowledge in the late middle ages: Rhythmic procedures as cultural representations. Science in Context, 9(2):121–136, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Damerow:1996:NSO

[223] Peter Damerow. Number as a second-order concept. Science in Context, 9 (2):139–149, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Freudenthal:1996:PR

[224] Gideon Freudenthal. Pluralism or relativism? Science in Context, 9(2): 151–162, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Barkan:1996:LSF

[225] Diana L. Kormos Barkan. The last scientist, the first magician: Dramatic and Epic theater as alternative images of science. Science in Context, 9 (2):163–175, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ophir:1996:TTT

[226] Adi Ophir. Two-tier thinking: a moral point of view. Science in Context, 9(2):177–188, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Krois:1996:NAP

[227] John Michael Krois. A note about philosophy and history: The place of o Cassirer’s Erkenntnisproblem. Science in Context, 9(2):191–194, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cassirer:1996:IFE

[228] Ernst Cassirer. From the introduction to the first edition of The Problem of Knowledge in Modern Philosophy and Science. Science in Context, 9 (2):195–215, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:AIW

[229] Anonymous. Appendix: International workshop on images of knowledge, two-tier thinking and higher education. Science in Context, 9(2):217–218, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIc

[230] Anonymous. SIC volume 9 issue 2 cover and front matter. Science in Context, 9(2):f1–f5, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVId

[231] Anonymous. SIC volume 9 issue 2 cover and back matter. Science in Context, 9(2):b1–b8, Summer 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lefevre:1996:IN

[232] Wolfgang Lefevre. Introductory note. Science in Context, 9(3):223, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Crosland:1996:CCC

[233] Maurice Crosland. Changes in chemical concepts and language in the Seventeenth Century. Science in Context, 9(3):225–240, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Multhauf:1996:OPE

[234] Robert P. Multhauf. Operational practice and the emergence of modern chemical concepts. Science in Context, 9(3):241–249, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Klein:1996:CWT

[235] Ursula Klein. The chemical workshop tradition and the experimental practice: Discontinuities within continuities. Science in Context, 9(3): 251–287, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Holmes:1996:CCE

[236] Frederic L. Holmes. The communal context for Etienne-François Geoffroy’s ‘Table des rapports’. Science in Context, 9(3):289–311, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Geoffroy:1996:TDR

[237] Etienne-François Geoffroy. Table of the different relations observed in chemistry between different substances 27 August 1718. Science in Context, 9(3):313–320, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIe

[238] Anonymous. SIC volume 9 issue 3 cover and front matter. Science in Context, 9(3):f1–f4, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIf

[239] Anonymous. SIC volume 9 issue 3 cover and back matter. Science in Context, 9(3):b1–b6, Autumn 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goenner:1996:AEP

[240] Hubert Goenner and Giuseppe Castagnetti. Albert Einstein as pacifist and Democrat during World War I. Science in Context, 9(4):325–386, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Johnston:1996:CCC

[241] Sean F. Johnston. The construction of colorimetry by committee. Science in Context, 9(4):387–420, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Clark:1996:MAA

[242] William Clark. On the ministerial archive of academic acts. Science in Context, 9(4):421–486, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Krieger:1996:MPC

[243] Martin H. Krieger. Making a paradigmatic convention normal: Entrenching means and variances as statistics. Science in Context, 9(4):487–509, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gierer:1996:OMS

[244] Alfred Gierer. Organisms-mechanisms: Stahl, Wolff, and the case against reductionist exclusion. Science in Context, 9(4):511–528, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lamarck:1996:S

[245] Jean-Baptiste Lamarck. Species. Science in Context, 9(4):531–539, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:1996:JBL

[246] Alexandre Métraux. Jean-Baptiste Lamarck’s quest for natural species. Science in Context, 9(4):541–553, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIg

[247] Anonymous. SIC volume 9 issue 4 cover and front matter. Science in Context, 9(4):f1–f4, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1996:SVIh

[248] Anonymous. SIC volume 9 issue 4 cover and back matter. Science in Context, 9(4):b1–b7, Winter 1996. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-menahem:1997:MCI

[249] Yemima Ben-menahem and Adi Ophir. Models of critique: Introduction. Science in Context, 10(1):3–12, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Beller:1997:CR

[250] Mara Beller. Criticism and revolutions. Science in Context, 10(1):13–37, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dascal:1997:CC

[251] Marcelo Dascal. Critique without critics? Science in Context, 10(1):39– 62, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Latour:1997:FST

[252] Bruno Latour. A few steps toward an anthropology of the iconoclastic gesture. Science in Context, 10(1):63–83, Spring 1997. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Unger:1997:SPB

[253] Roberto Mangabeira Unger. Science and politics between domesticated and radicalized pragmatism. Science in Context, 10(1):85–95, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Alcoff:1997:IT

[254] Linda Martín Alcoff. Immanent truth. Science in Context, 10(1):97–112, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Longino:1997:IVE

[255] Helen E. Longino. Interpretation versus explanation in the critique of science. Science in Context, 10(1):113–128, Spring 1997. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brunner:1997:FES

[256] José Brunner. Fear and envy: Sexual difference and the economies of feminist critique in psychoanalytic discourse. Science in Context, 10 (1):129–170, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Freudenthal:1997:MCE

[257] Gideon Freudenthal. Marx’s critique of economic reason. Science in Context, 10(1):171–198, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Roubach:1997:HSM

[258] Michael Roubach. Heidegger, science, and the mathematical age. Science in Context, 10(1):199–206, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Motzkin:1997:ITB

[259] Gabriel Motzkin. The intuition of time between science and art history in the early Twentieth Century. Science in Context, 10(1):207–220, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIa

[260] Anonymous. SIC volume 10 issue 1 cover and front matter. Science in Context, 10(1):f1–f4, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIb

[261] Anonymous. SIC volume 10 issue 1 cover and back matter. Science in Context, 10(1):b1–b9, Spring 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Chemla:1997:WSM

[262] Karine Chemla. What is at stake in mathematical proofs from Third-Century China? Science in Context, 10(2):227–251, Summer 1997. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:1997:OET

[263] Leo Corry. The origins of eternal truth in modern mathematics: Hilbert to Bourbaki and beyond. Science in Context, 10(2):253–296, Summer 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Aubin:1997:WIN

[264] David Aubin. The withering immortality of Nicolas Bourbaki: a cultural connector at the confluence of mathematics, structuralism, and the Oulipo in France. Science in Context, 10(2):297–342, Summer 1997. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Smith:1997:WCO

[265] C. U. M. Smith. Worlds in collision: Owen and Huxley on the brain. Science in Context, 10(2):343–365, Summer 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dewsbury:1997:RSP

[266] Donald A. Dewsbury. Rhetorical strategies in the presentation of ethology and comparative psychology in magazines after World War II. Science in Context, 10(2):367–386, Summer 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIc

[267] Anonymous. SIC volume 10 issue 2 cover and front matter. Science in Context, 10(2):f1–f4, Summer 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVId

[268] Anonymous. SIC volume 10 issue 2 cover and back matter. Science in Context, 10(2):b1–b4, Summer 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Langermann:1997:F

[269] Y. Tzvi Langermann. Foreword. Science in Context, 10(3):391–392, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Manekin:1997:WJL

[270] Charles H. Manekin. When the Jews learned logic from the Pope: Three Medieval Hebrew translations of the Tractatus of Peter of Spain. Science in Context, 10(3):395–430, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Levy:1997:EMB

[271] Tony Lévy. The establishment of the mathematical bookshelf of the Medieval Hebrew scholar: Translations and translators. Science in Context, 10(3):431–451, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Glasner:1997:YPU

[272] Ruth Glasner. Yeda’aya ha-Penini’s unusual conception of void. Science in Context, 10(3):453–470, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mancha:1997:LBG

[273] J. L. Mancha. Levi ben Gerson’s astronomical work: Chronology and Christian context. Science in Context, 10(3):471–493, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Langermann:1997:SB

[274] Y. Tzvi Langermann. Science and the Kuzari. Science in Context, 10(3): 495–522, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIe

[275] Anonymous. SIC volume 10 issue 3 cover and front matter. Science in Context, 10(3):f1–f4, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIf

[276] Anonymous. SIC volume 10 issue 3 cover and back matter. Science in Context, 10(3):b1–b4, Autumn 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Efron:1997:I

[277] Noah Efron. Introduction. Science in Context, 10(4):527–528, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Tirosh-Samuelson:1997:TNS

[278] Hava Tirosh-Samuelson. Theology of nature in Sixteenth-Century Italian Jewish philosophy. Science in Context, 10(4):529–570, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fishman:1997:RMI

[279] David E. Fishman. Rabbi Moshe Isserles and the study of science among Polish rabbis. Science in Context, 10(4):571–588, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Reiner:1997:AAS

[280] Elchanan Reiner. The attitude of Ashkenazi society to the new science in the Sixteenth Century. Science in Context, 10(4):589–603, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Davis:1997:ARM

[281] Joseph Davis. Ashkenazic rationalism and Midrashic natural history: Responses to the new science in the works of Rabbi Yom Tov Lipmann Heller (1578–1654). Science in Context, 10(4):605–626, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Efron:1997:INP

[282] Noah J. Efron. Irenism and natural philosophy in Rudolfine Prague: The case of David Gans. Science in Context, 10(4):627–649, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goldish:1997:NCD

[283] Matt Goldish. Newtonian, Converso, and Deist: The lives of Jacob (Henrique) de Castro Sarmento. Science in Context, 10(4):651–675, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ruderman:1997:DJS

[284] David Ruderman. On defining a Jewish stance toward Newtonianism: Eliakim ben Abraham Hart’s Wars of the Lord. Science in Context, 10 (4):677–691, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIg

[285] Anonymous. SIC volume 10 issue 4 cover and front matter. Science in Context, 10(4):f1–f4, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1997:SVIh

[286] Anonymous. SIC volume 10 issue 4 cover and back matter. Science in Context, 10(4):b1–b3, Winter 1997. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stachel:1998:OEE

[287] John Stachel. The other Einstein: Einstein contra field theory. Science in Context, 6(1):275–290, March 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rosenband:1998:PPP

[288] Leonard N. Rosenband. The perils of petty production: Pierre and Jean-Baptiste Serve of Chamalières. Science in Context, 11(1):3–21, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Benschop:1998:WTH

[289] Ruth Benschop. What is a tachistoscope? Historical explorations of an instrument. Science in Context, 11(1):23–50, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kaiser:1998:MAS

[290] David Kaiser. A Mannheim for all seasons: Bloor, Merton, and the roots of the sociology of scientific knowledge. Science in Context, 11(1):51–87, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jha:1998:DIT

[291] S. R. Jha. On the duties of intellectuals to truth: The life and work of chemist-philosopher Michael Polanyi. Science in Context, 11(1):89–141, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVIa

[292] Anonymous. SIC volume 11 issue 1 cover and front matter. Science in Context, 11(1):f1–f4, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVIb

[293] Anonymous. SIC volume 11 issue 1 cover and back matter. Science in Context, 11(1):b1–b4, Spring 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dascal:1998:SCT

[294] Marcelo Dascal. The study of controversies and the theory and history of science. Science in Context, 11(2):147–154, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Freudenthal:1998:C

[295] Gideon Freudenthal. ‘Controversy’. Science in Context, 11(2):155–160, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gross:1998:DDP

[296] Alan G. Gross. Do disputes over priority tell us anything about science? Science in Context, 11(2):161–179, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zuber:1998:DDC

[297] Marta Spranzi Zuber. Dialectic, dialogue, and controversy: The case of Galileo. Science in Context, 11(2):181–203, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Senderowicz:1998:FBT

[298] Yaron Senderowicz. Facing the bounds of tradition: Kant’s controversy with the philosophisches magazin. Science in Context, 11(2):205–228, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cremaschi:1998:MRT

[299] Sergio Cremaschi and Marcelo Dascal. Malthus and Ricardo: Two styles for economic theory. Science in Context, 11(2):229–254, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gruengard:1998:IEP

[300] Ora Gruengard. Introverted, extroverted, and perverted controversy: Jung against Freud. Science in Context, 11(2):255–290, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carl:1998:BVH

[301] J. Posy Carl. Brouwer versus Hilbert: 1907–1928. Science in Context, 11 (2):291–325, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVIc

[302] Anonymous. SIC volume 11 issue 2 cover and front matter. Science in Context, 11(2):f1–f4, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVId

[303] Anonymous. SIC volume 11 issue 2 cover and back matter. Science in Context, 11(2):b1–b4, Summer 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Falk:1998:F

[304] Raphael Falk, Diane B. Paul, and Garland Allen. Foreword. Science in Context, 11(3–4):329–330, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Balaban:1998:EIP

[305] Evan Balaban. Eugenics and individual phenotypic variation: To what extent is biology a predictive science? Science in Context, 11(3–4): 331–356, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Benjamin:1998:GHP

[306] Jonathan Benjamin. Genes for human personality traits. Science in Context, 11(3–4):357–372, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Templeton:1998:CGP

[307] Alan R. Templeton. The complexity of the genotype-phenotype relationship and the limitations of using genetic ‘markers’ at the individual level. Science in Context, 11(3–4):373–389, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carmi:1998:LBH

[308] Rivka Carmi, Khalil Elbedour, Dahlia Wietzman, Val Sheffield, and Ilana Shoham-Vardi. Lowering the burden of hereditary diseases in a traditional, inbred community: Ethical aspects of genetic research and its application. Science in Context, 11(3–4):391–395, January 1998. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Holtzman:1998:EGT

[309] Neil A. Holtzman. Eugenics and genetic testing. Science in Context, 11(3–4):397–417, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sagi:1998:EAG

[310] Michal Sagi. Ethical aspects of genetic screening in Israel. Science in Context, 11(3–4):419–429, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Resta:1998:ECT

[311] Robert G. Resta. Eugenic considerations in the theory and practice of genetic counseling. Science in Context, 11(3–4):431–438, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Daniels:1998:NPG

[312] Norman Daniels. Negative and positive genetic interventions: Is there a moral boundary? Science in Context, 11(3–4):439–453, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wilker:1998:EV

[313] Daniel Wilker. Eugenic values. Science in Context, 11(3–4):455–470, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Chadwick:1998:CGC

[314] Ruth Chadwick. Can genetic counseling avoid the charge of eugenics? Science in Context, 11(3–4):471–480, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Paul:1998:GSE

[315] Diane B. Paul. Genetic services, economics, and eugenics. Science in Context, 11(3–4):481–491, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wertz:1998:EAW

[316] Dorothy C. Wertz. Eugenics is alive and well: a survey of genetic professionals around the world. Science in Context, 11(3–4):493–510, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zenderland:1998:BBA

[317] Leila Zenderland. Biblical biology: American Protestant social reformers and the early eugenics movement. Science in Context, 11(3–4):511–525, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lepicard:1998:ERC

[318] Etienne Lepicard. Eugenics and Roman Catholicism: An encyclical letter in context: Casti connubii, December 31, 1930. Science in Context, 11(3– 4):527–544, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Otsubo:1998:EJS

[319] Sumiko Otsubo and James R. Bartholomew. Eugenics in Japan: Some ironies of modernity, 1883–1945. Science in Context, 11(3–4):545–565, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rispler-chaim:1998:GEC

[320] Vardit Rispler-chaim. Genetic engineering in contemporary Islamic thought. Science in Context, 11(3–4):567–573, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zohar:1998:LSM

[321] Noam J. Zohar. From lineage to sexual mores: Examining ‘Jewish eugenics’. Science in Context, 11(3–4):575–585, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Falk:1998:ZBJ

[322] Raphael Falk. Zionism and the biology of the Jews. Science in Context, 11(3–4):587–607, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mocek:1998:PPR

[323] Reinhard Mocek. The program of proletarian Rassenhygiene. Science in Context, 11(3–4):609–617, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weindling:1998:DGS

[324] Paul Weindling. Dissecting German social Darwinism: Historicizing the biology of the organic state. Science in Context, 11(3–4):619–637, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVIe

[325] Anonymous. SIC volume 11 issue 3–4 cover and front matter. Science in Context, 11(3–4):f1–f4, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1998:SVIf

[326] Anonymous. SIC volume 11 issue 3–4 cover and back matter. Science in Context, 11(3–4):b1–b3, January 1998. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:I

[327] Anonymous. Introduction. Science in Context, 12(1):3–6, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Golan:1999:HSE

[328] Tal Golan. The history of scientific expert testimony in the English courtroom. Science in Context, 12(1):7–32, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Voelkel:1999:PPL

[329] James R. Voelkel. Publish or perish: Legal contingencies and the publication of Kepler’s Astronomia nova. Science in Context, 12(1):33–59, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Riskin:1999:LLR

[330] Jessica Riskin. The lawyer and the lightning rod. Science in Context, 12(1):61–99, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schweber:1999:LNS

[331] Howard Schweber. Law and the natural sciences in Nineteenth-Century American universities. Science in Context, 12(1):101–121, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rudavsky:1999:SSL

[332] Shari Rudavsky. Separating spheres: Legal ideology v. paternity testing in divorce cases. Science in Context, 12(1):123–138, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cole:1999:WCI

[333] Simon Cole. What counts for identity? The historical origins of the methodology of latent fingerprint identification. Science in Context, 12 (1):139–172, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dumit:1999:OBP

[334] Joseph Dumit. Objective brains, prejudicial images. Science in Context, 12(1):173–201, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Landecker:1999:BBC

[335] Hannah Landecker. Between beneficence and chattel: The human biological in law and science. Science in Context, 12(1):203–225, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-Menahem:1999:LSR

[336] Hanina Ben-Menahem and Yemima Ben-Menahem. Law and science — reflections. Science in Context, 12(1):227–243, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIa

[337] Anonymous. SIC volume 12 issue 1 cover and front matter. Science in Context, 12(1):f1–f4, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIb

[338] Anonymous. SIC volume 12 issue 1 cover and back matter. Science in Context, 12(1):b1–b3, Spring 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sismondo:1999:MST

[339] Sergio Sismondo. Models, simulations, and their objects. Science in Context, 12(2):247–260, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dowling:1999:ET

[340] Deborah Dowling. Experimenting on theories. Science in Context, 12(2): 261–273, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Winsberg:1999:SME

[341] Eric Winsberg. Sanctioning models: The epistemology of simulation. Science in Context, 12(2):275–292, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Merz:1999:MUC

[342] Martina Merz. Multiplex and unfolding: Computer simulation in particle physics. Science in Context, 12(2):293–316, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Breslau:1999:BMM

[343] Daniel Breslau and Yuval Yonay. Beyond metaphor: Mathematical models in economics as empirical research. Science in Context, 12(2):317–332, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vandenBogaard:1999:COD

[344] Adrienne van den Bogaard. The cultural origins of the Dutch economic modeling practice. Science in Context, 12(2):333–350, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Evans:1999:EMP

[345] Robert Evans. Economic models and policy advice: Theory choice or moral choice? Science in Context, 12(2):351–376, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIc

[346] Anonymous. SIC volume 12 issue 2 cover and front matter. Science in Context, 12(2):f1–f4, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVId

[347] Anonymous. SIC volume 12 issue 2 cover and back matter. Science in Context, 12(2):b1–b6, Summer 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brentjes:1999:CBN

[348] Sonja Brentjes. Crossing boundaries: New approaches to the history of ‘pre-modern’ science and technology. Science in Context, 12(3):381–384, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Chu:1999:TIC

[349] Pingyi Chu. Trust, instruments, and cross-cultural scientific exchanges: Chinese debate over the shape of the Earth, 1600–1800. Science in Context, 12(3):385–412, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jami:1999:ESC

[350] Catherine Jami. ‘European Science in China’ or ‘Western Learning’? Representations of cross-cultural transmission, 1600–1800. Science in Context, 12(3):413–434, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brentjes:1999:IRL

[351] Sonja Brentjes. The interests of the Republic of Letters in the Middle East, 1550–1700. Science in Context, 12(3):435–468, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Moyer:1999:RRI

[352] Ann Moyer. Renaissance representations of Islamic science: Bernardino Baldi and His Lives of Mathematicians. Science in Context, 12(3):469– 484, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIe

[353] Anonymous. SIC volume 12 issue 3 cover and front matter. Science in Context, 12(3):f1–f4, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIf

[354] Anonymous. SIC volume 12 issue 3 cover and back matter. Science in Context, 12(3):b1–b3, Autumn 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Krois:1999:EI

[355] John Michael Krois. Editor’s introduction. Science in Context, 12(4): 489–491, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dor:1999:SFL

[356] Daniel Dor. From symbolic forms to lexical semantics: Where modern linguistics and Cassirer’s philosophy start to converge. Science in Context, 12(4):493–511, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ihmig:1999:ECS

[357] Karol-Nobert Ihmig. Ernst Cassirer and the structural conception of objects in modern science: The importance of the ‘Erlanger Programm’. Science in Context, 12(4):513–529, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Krois:1999:CPM

[358] John Michael Krois. Cassirer’s ‘prototype and model’ of symbolism: Its sources and significance. Science in Context, 12(4):531–547, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Moynahan:1999:ECT

[359] Gregory B. Moynahan. Ernst Cassirer, theoretical biology, and the Clever Hans phenomenon. Science in Context, 12(4):549–574, Winter 1999. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Naumann:1999:GSF

[360] Barbara Naumann. The genesis of symbolic forms: Basis phenomena in Ernst Cassirer’s works. Science in Context, 12(4):575–584, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ryckman:1999:ECG

[361] T. A. Ryckman. Einstein, Cassirer, and general covariance — then and now. Science in Context, 12(4):585–619, Winter 1999. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schoell-Glass:1999:AWL

[362] Charlotte Schoell-Glass. Aby Warburg’s late comments on symbol and ritual. Science in Context, 12(4):621–642, Winter 1999. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:1999:PNP

[363] Alexandre Mètraux. Philosophy, neurology, and the pathology of symbolic consciousness: On two unpublished letters from Ernst Cassirer to Kurt Goldstein. Science in Context, 12(4):645–660, Winter 1999. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cassirer:1999:TLK

[364] Ernst Cassirer. Two letters to Kurt Goldstein. Science in Context, 12 (4):661–667, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIg

[365] Anonymous. SIC volume 12 issue 4 cover and front matter. Science in Context, 12(4):f1–f4, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:1999:SVIh

[366] Anonymous. SIC volume 12 issue 4 cover and back matter. Science in Context, 12(4):b1–b3, Winter 1999. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:2000:GFL

[367] Leo Corry, Yehuda Elkana, Snait Gissis, Alexandre Métraux, and Jürgen Renn. Gideon Freudenthal leaves Science in Context. Science in Context, 13(1):3–4, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:2000:EI

[368] Alexandre Métraux. Editor’s introduction. Science in Context, 13(1):5– 7, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gradmann:2000:IEB

[369] Christoph Gradmann. Invisible enemies: Bacteriology and the language of politics in Imperial Germany. Science in Context, 13(1):9–30, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jansen:2000:AII

[370] Sarah Jansen. An American insect in Imperial Germany: Visibility and control in making the Phylloxera in Germany, 1870–1914. Science in Context, 13(1):31–70, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Parnes:2000:EC

[371] Ohad Parnes. The envisioning of cells. Science in Context, 13(1):71–92, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ratcliff:2000:WLM

[372] M. J. Ratcliff. Wonders, logic, and microscopy in the Eighteenth Century: a history of the rotifer. Science in Context, 13(1):93–119, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rheinberger:2000:IA

[373] Hans-Jörg Rheinberger. Invisible architectures. Science in Context, 13 (1):121–136, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schickore:2000:LRC

[374] Jutta Schickore. Locating rods and cones: Microscopic investigations of the retina in mid-Nineteenth-Century Berlin and Würzburg. Science in Context, 13(1):137–152, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVIa

[375] Anonymous. SIC volume 13 issue 1 cover and front matter. Science in Context, 13(1):f1–f4, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVIb

[376] Anonymous. SIC volume 13 issue 1 cover and back matter. Science in Context, 13(1):b1–b2, Spring 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lustig:2000:CKN

[377] A. J. Lustig. Cultivating knowledge in Nineteenth-Century English gardens. Science in Context, 13(2):155–181, Summer 2000. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Richards:2000:BCP

[378] Graham Richards. Britain on the couch: The popularization of psychoanalysis in Britain 1918–1940. Science in Context, 13(2):183–230, Summer 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wakefield:2000:PC

[379] R. Andre Wakefield. Police chemistry. Science in Context, 13(2):231–267, Summer 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVIc

[380] Anonymous. SIC volume 13 issue 2 cover and front matter. Science in Context, 13(2):f1–f4, Summer 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVId

[381] Anonymous. SIC volume 13 issue 2 cover and back matter. Science in Context, 13(2):b1–b3, Summer 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:2000:GCE

[382] Jürgen Renn. Galileo in context: An engineer-scientist, artist, and courtier at the origins of classical science. Science in Context, 13(3– 4):271–278, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lefevre:2000:GEA

[383] Wolfgang Lefèvre. Galileo engineer: Art and modern science. Science in Context, 13(3–4):281–297, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:2000:HWE

[384] Jürgen Renn, Peter Damerow, Simone Rieger, and Domenico Giulini. Hunting the white elephant: When and how did Galileo discover the law of fall? Science in Context, 13(3–4):299–419, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bredekamp:2000:GHB

[385] Horst Bredekamp. Gazing hands and blind spots: Galileo as draftsman. Science in Context, 13(3–4):423–462, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Booth:2000:VTM

[386] Sara Elizabeth Booth and Albert van Helden. The virgin and the telescope: The moons of Cigoli and Galileo. Science in Context, 13(3–4): 463–486, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:2000:RNG

[387] Rivka Feldhay. Recent narratives on Galileo and the church: or the three dogmas of the counter-reformation. Science in Context, 13(3–4):489–507, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Galluzzi:2000:GLG

[388] Paolo Galluzzi. Gassendi and l’Affaire Galilée of the Laws of Motion. Science in Context, 13(3–4):509–545, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Biagioli:2000:RME

[389] Mario Biagioli. Replication or monopoly? The economies of invention and discovery in Galileo’s observations of 1610. Science in Context, 13(3– 4):547–590, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:2000:FCI

[390] Jürgen Renn. A forgotten controversy introductory note to the appendix. Science in Context, 13(3–4):593–595, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Castagnetti:2000:RCH

[391] Giuseppe Castagnetti and Michele Camerota. Raffaello Caverni and his History of the Experimental Method in Italy. Science in Context, 13(3– 4):597–609, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Caverni:2000:EBE

[392] Raffaello Caverni. An excerpt from History of the Experimental Method in Italy. Science in Context, 13(3–4):611–625, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Castagnetti:2000:AFE

[393] Giuseppe Castagnetti and Michele Camerota. Antonio Favaro and the Edizione Nazionale of Galileo’s works. Science in Context, 13(3–4):627– 631, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Favaro:2000:AGW

[394] Antonio Favaro. Apocryphal Galilean writings. Science in Context, 13(3– 4):633–639, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schutt:2000:EWG

[395] Hans-Werner Schütt. Emil Wohlwill, Galileo and his battle for the Copernican system. Science in Context, 13(3–4):641–643, January 2000. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wohlwill:2000:DPS

[396] Emil Wohlwill. The discovery of the parabolic shape of the projectile trajectory. Science in Context, 13(3–4):645–680, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:RA

[397] Anonymous. References for appendix. Science in Context, 13(3–4):681– 691, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVIe

[398] Anonymous. SIC volume 13 issue 3–4 cover and front matter. Science in Context, 13(3–4):f1–f4, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2000:SVIf

[399] Anonymous. SIC volume 13 issue 3–4 cover and back matter. Science in Context, 13(3–4):b1–b3, January 2000. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Renn:2001:GCE

[400] Jürgen Renn. Galileo in context: An engineer-scientist, artist, and courtier at the origins of classical science. Science in Context, 14(S1): 1–8, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Lefevre:2001:GEA

[401] Wolfgang Lefèvre. Galileo engineer: Art and modern science. Science in Context, 14(S1):11–27, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Renn:2001:HWE

[402] Jürgen Renn, Peter Damerow, Simone Rieger, and Domenico Giulini. Hunting the white elephant: When and how did Galileo discover the law of fall? Science in Context, 14(S1):29–149, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Bredekamp:2001:GHB

[403] Horst Bredekamp. Gazing hands and blind spots: Galileo as draftsman. Science in Context, 14(S1):153–192, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Booth:2001:VTM

[404] Sara Elizabeth Booth and Albert Van Helden. The virgin and the telescope: The moons of Cigoli and Galileo. Science in Context, 14(S1):

193–216, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Feldhay:2001:RNG

[405] Rivka Feldhay. Recent narratives on Galileo and the church: or the three dogmas of the counter-reformation. Science in Context, 14(S1): 219–237, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Galluzzi:2001:GLG

[406] Paolo Galluzzi. Gassendi and l’Affaire Galilée of the Laws of Motion. Science in Context, 14(S1):239–275, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Biagioli:2001:RME

[407] Mario Biagioli. Replication or monopoly? The economies of invention and discovery in Galileo’s observations of 1610. Science in Context, 14 (S1):277–320, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Renn:2001:FCI

[408] Jürgen Renn. A forgotten controversy introductory note to the appendix. Science in Context, 14(S1):321–325, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Castagnetti:2001:RCH

[409] Giuseppe Castagnetti and Michele Camerota. Raffaello Caverni and his history of the experimental method in Italy. Science in Context, 14(S1): 327–339, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Caverni:2001:EHE

[410] Raffaello Caverni. An excerpt from history of the experimental method in Italy. Science in Context, 14(S1):341–355, June 2001. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Castagnetti:2001:AFE

[411] Giuseppe Castagnetti and Michele Camerota. Antonio Favaro and the Edizione Nazionale of Galileo’s works. Science in Context, 14(S1):357– 361, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Favaro:2001:AGW

[412] Antonio Favaro. Apocryphal Galilean writings. Science in Context, 14 (S1):363–369, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Schutt:2001:EWG

[413] Hans-Werner Schütt. Emil Wohlwill, Galileo and his battle for the Copernican system. Science in Context, 14(S1):371–373, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Wohlwill:2001:DPS

[414] Emil Wohlwill. The discovery of the parabolic shape of the projectile trajectory. Science in Context, 14(S1):375–410, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Anonymous:2001:RA

[415] Anonymous. References for appendix. Science in Context, 14(S1):411– 421, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Anonymous:2001:I

[416] Anonymous. Index. Science in Context, 14(S1):423–431, June 2001. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Galileo in Context.

Abattouy:2001:TTT

[417] Mohammed Abattouy, Jürgen Renn, and Paul Weinig. Transmission as transformation: The translation movements in the Medieval East and West in a comparative perspective. Science in Context, 14(1–2):1–12, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Folkerts:2001:ETH

[418] Menso Folkerts. Early texts on Hindu–Arabic calculation. Science in Context, 14(1–2):13–38, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brentjes:2001:OHC

[419] Sonja Brentjes. Observations on Hermann of Carinthia’s version of the Elements and its relation to the Arabic transmission. Science in Context, 14(1–2):39–84, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hoyrup:2001:CGR

[420] Jens Høyrup. On a collection of geometrical riddles and their role in the shaping of four to six “algebras”. Science in Context, 14(1–2):85– 131, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Knorr:2001:HE

[421] Wilbur R. Knorr. On Heiberg’s Euclid. Science in Context, 14(1–2):133– 143, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ragep:2001:TCE

[422] F. Jamil Ragep. Tūsī and Copernicus: The Earth’s motion in context. Science in Context, 14(1–2):145–163, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Samso:2001:AOM

[423] Julio Samsó. Astronomical observations in the Maghrib in the Fourteenth and Fifteenth Centuries. Science in Context, 14(1–2):165–178, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Abattouy:2001:GMA

[424] Mohammed Abattouy. Greek mechanics in Arabic context: Thabit ibn Qurra, al-Isfizari and the Arabic traditions of Aristotelian and Euclidean mechanics. Science in Context, 14(1–2):179–247, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Burnett:2001:CAL

[425] Charles Burnett. The coherence of the Arabic–Latin translation program in Toledo in the Twelfth Century. Science in Context, 14(1–2):249– 288, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schramm:2001:FIH

[426] Matthias Schramm. Frederick II of Hohenstaufen and Arabic science. Science in Context, 14(1–2):289–312, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lorch:2001:GAL

[427] Richard Lorch. Greek–Arabic–Latin: The transmission of mathematical texts in the Middle Ages. Science in Context, 14(1–2):313–331, June 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Remmert:2001:SRA

[428] Volker R. Remmert. In the service of the Reich: Aspects of Copernicus and Galileo in Nazi Germany’s historiographical and political discourse. Science in Context, 14(3):333–359, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kim:2001:AIC

[429] Mi Gyung Kim. The analytic ideal of chemical elements: Robert Boyle and the French didactic tradition of chemistry. Science in Context, 14(3): 361–395, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schickore:2001:TES

[430] Jutta Schickore. The task of explaining sight — Helmholtz’s writings on vision as a test case for models of science popularization. Science in Context, 14(3):397–417, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hyder:2001:POP

[431] David Jalal Hyder. Physiological optics and physical geometry. Science in Context, 14(3):419–456, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heinzmann:2001:FGC

[432] Gerhard Heinzmann. The foundations of geometry and the concept of motion: Helmholtz and Poincaré. Science in Context, 14(3):457–470, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Debru:2001:HPT

[433] Claude Debru. Helmholtz and the psychophysiology of time. Science in Context, 14(3):471–492, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

McDonald:2001:RCH

[434] Patrick McDonald. Remarks on the context of Helmholtz’s “Über das Wesen der Fäulniss und Gährung”. Science in Context, 14(3):493–498, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vonHelmholtz:2001:NPF

[435] Hermann von Helmholtz. On the nature of putrefaction and fermentation. Science in Context, 14(3):499–504, September 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hagner:2001:MPN

[436] Michael Hagner and Cornelius Borck. Mindful practices: On the neurosciences in the Twentieth Century. Science in Context, 14(4):507–510, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Smith:2001:RMC

[437] Roger Smith. Representations of mind: C. S. Sherrington and scientific opinion, c. 1930–1950. Science in Context, 14(4):511–539, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hagner:2001:CCG

[438] Michael Hagner. Cultivating the cortex in German neuroanatomy. Science in Context, 14(4):541–563, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Borck:2001:EMP

[439] Cornelius Borck. Electricity as a medium of psychic life: Electrotechnological adventures into psychodiagnosis in Weimar Germany. Science in Context, 14(4):565–590, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kay:2001:LNP

[440] Lily E. Kay. From logical neurons to poetic embodiments of mind: Warren S. McCulloch’s project in neuroscience. Science in Context, 14(4): 591–614, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hayward:2001:TLM

[441] Rhodri Hayward. The tortoise and the love-machine: Grey Walter and the politics of electroencephalography. Science in Context, 14(4):615– 641, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dror:2001:TBP

[442] Otniel E. Dror. Techniques of the brain and the paradox of emotions, 1880–1930. Science in Context, 14(4):643–660, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Young:2001:OTN

[443] Allan Young. Our traumatic neurosis and its brain. Science in Context, 14(4):661–683, December 2001. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heyd:2002:I

[444] Michael Heyd. Introduction. Science in Context, 15(1):1–8, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Webster:2002:PPS

[445] Charles Webster. Paracelsus, paracelsianism, and the secularization of the worldview. Science in Context, 15(1):9–27, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Efron:2002:NHN

[446] Noah J. Efron. Nature, human nature, and Jewish nature in early modern Europe. Science in Context, 15(1):29–49, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ben-Chaim:2002:ELB

[447] Michael Ben-Chaim. Empowering lay belief: Robert Boyle and the moral economy of experiment. Science in Context, 15(1):51–77, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feingold:2002:SCE

[448] Mordechai Feingold. Science as a calling? The early modern dilemma. Science in Context, 15(1):79–119, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feiner:2002:SSE

[449] Shmuel Feiner. Seductive science and the emergence of the secular Jewish intellectual. Science in Context, 15(1):121–135, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Richards:2002:RWA

[450] Joan L. Richards. “In a rational world all radicals would be exterminated”: Mathematics, logic and secular thinking in Augustus De Morgan’s England. Science in Context, 15(1):137–164, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Motzkin:2002:SSD

[451] Gabriel Motzkin. Science, secularization, and desecularization at the turn of the Twentieth Century. Science in Context, 15(1):165–175, March 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kojevnikov:2002:INH

[452] Alexei Kojevnikov. Introduction: a new history of Russian science. Science in Context, 15(2):177–182, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bayuk:2002:LMS

[453] Dimitri Bayuk. Literature, music, and science in Nineteenth Century Russian culture: Prince Odoyevskiy’s quest for a natural enharmonic scale. Science in Context, 15(2):183–207, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Elina:2002:PSR

[454] Olga Elina. Planting seeds for the revolution: The rise of Russian agricultural science, 1860–1920. Science in Context, 15(2):209–237, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kojevnikov:2002:GWR

[455] Alexei Kojevnikov. The great war, the Russian civil war, and the invention of big science. Science in Context, 15(2):239–275, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rossiianov:2002:BSI

[456] Kirill Rossiianov. Beyond species: Il’ya Ivanov and his experiments on cross-breeding humans with anthropoid apes. Science in Context, 15 (2):277–316, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ivanov:2002:SAS

[457] Konstantin Ivanov. Science after Stalin: Forging a new image of Soviet science. Science in Context, 15(2):317–338, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gerovitch:2002:LHM

[458] Slava Gerovitch. Love-hate for man-machine metaphors in Soviet physiology: From Pavlov to “physiological cybernetics”. Science in Context, 15(2):339–374, June 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Desrosieres:2002:TSH

[459] Alain Desrosières. Three studies on the history of sampling surveys: Norway, Russia–USSR, United States. Science in Context, 15(3):377– 383, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lie:2002:RFS

[460] Einar Lie. The rise and fall of sampling surveys in Norway, 1875–1906. Science in Context, 15(3):385–409, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mespoulet:2002:TAR

[461] Martine Mespoulet. From typical areas to random sampling: Sampling methods in Russia from 1875 to 1930. Science in Context, 15(3):411–425, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Didier:2002:SDR

[462] Emmanuel Didier. Sampling and democracy: Representativeness in the first United States surveys. Science in Context, 15(3):427–445, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Smolarski:2002:JBS

[463] Dennis C. Smolarski. The Jesuit Ratio Studiorum, Christopher Clavius, and the study of mathematical sciences in universities. Science in Context, 15(3):447–457, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2002:HDP

[464] Anonymous. Historical documents, Part I: Sections on mathematics from the various editions of the Ratio Studiorum. Science in Context, 15(3): 459–464, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Clavius:2002:HDP

[465] Christopher Clavius. Historical documents, Part II: Two documents on mathematics written by Christopher Clavius, S.J. Science in Context, 15(3):465–470, September 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schiebinger:2002:EWS

[466] Londa Schiebinger. European women in science. Science in Context, 15(4):473–481, December 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bosch:2002:WSN

[467] Mineke Bosch. Women and science in The Netherlands: a Dutch case? Science in Context, 15(4):483–527, December 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hermann:2002:WSF

[468] Claudine Hermann and Françoise Cyrot-Lackmann. Women in science in France. Science in Context, 15(4):529–556, December 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Costas:2002:WSG

[469] Ilse Costas. Women in science in Germany. Science in Context, 15(4): 557–576, December 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Singh:2002:BBG

[470] Ilina Singh. Bad boys, good mothers, and the “miracle” of ritalin. Science in Context, 15(4):577–603, December 2002. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Daston:2003:ISP

[471] Lorraine Daston and H. Otto Sibum. Introduction: Scientific personae and their histories. Science in Context, 16(1–2):1–8, March 2003. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Algazi:2003:SHR

[472] Gadi Algazi. Scholars in households: Refiguring the learned habitus, 1480–1550. Science in Context, 16(1–2):9–42, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Clark:2003:PV

[473] William Clark. On the professorial voice. Science in Context, 16(1–2):43– 57, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Findlen:2003:BSG

[474] Paula Findlen. Becoming a scientist: Gender and knowledge in Eighteenth-Century Italy. Science in Context, 16(1–2):59–87, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sibum:2003:ERL

[475] H. Otto Sibum. Experimentalists in the Republic of letters. Science in Context, 16(1–2):89–120, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jackson:2003:HIN

[476] Myles Jackson. Harmonious investigators of nature: Music and the persona of the German Naturforscher in the Nineteenth Century. Science in Context, 16(1–2):121–145, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Secord:2003:WYW

[477] Anne Secord. “Be what you would seem to be”: Samuel Smiles, Thomas Edward, and the making of a working-class scientific hero. Science in Context, 16(1–2):147–173, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Browne:2003:CDC

[478] Janet Browne. Charles Darwin as a celebrity. Science in Context, 16(1– 2):175–194, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hagner:2003:SBM

[479] Michael Hagner. Skulls, brains, and memorial culture: On cerebral biographies of scientists in the Nineteenth Century. Science in Context, 16 (1–2):195–218, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schweber:2003:JRO

[480] Silvan S. Schweber. J. Robert Oppenheimer: Proteus unbound. Science in Context, 16(1–2):219–242, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carson:2003:OSH

[481] Cathryn Carson. Objectivity and the scientist: Heisenberg rethinks. Science in Context, 16(1–2):243–269, March 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2003:EP

[482] Anonymous. Editor’s preface. Science in Context, 16(3):273–274, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Netz:2003:IHE

[483] Reviel Netz. Introduction: The history of early mathematics — ways of re-writing. Science in Context, 16(3):275–286, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mueller:2003:REB

[484] Ian Mueller. Remarks on Euclid’s Elements I,32 and the parallel postulate. Science in Context, 16(3):287–297, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Taisbak:2003:EFS

[485] Christian Marinus Taisbak. Exceeding and falling short: Elliptical and hyperbolical application of areas. Science in Context, 16(3):299–318, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Yavetz:2003:STR

[486] Ido Yavetz. On Simplicius’ testimony regarding Eudoxan lunar theory. Science in Context, 16(3):319–329, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Saito:2003:PTP

[487] Ken Saito. Phantom theories of pre-Eudoxean proportion. Science in Context, 16(3):331–347, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fried:2003:UAB

[488] Michael Fried. The use of analogy in Book VII of Apollonius’ Conica. Science in Context, 16(3):349–365, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Imhausen:2003:EMT

[489] Annette Imhausen. Egyptian mathematical texts and their contexts. Science in Context, 16(3):367–389, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bernard:2003:ARG

[490] Alain Bernard. Ancient rhetoric and Greek mathematics: a response to a modern historiographical dilemma. Science in Context, 16(3):391–412, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Chemla:2003:GAA

[491] Karine Chemla. Generality above abstraction: The general expressed in terms of the paradigmatic in mathematics in Ancient China. Science in Context, 16(3):413–458, September 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Muller-Wille:2003:JLT

[492] Staffan Müller-Wille. Joining Lapland and the Topinambes in flourishing Holland: Center and periphery in Linnaean botany. Science in Context, 16(4):461–488, December 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Castillo:2003:NDP

[493] Jesús Carrillo Castillo. Naming difference: The politics of naming in Fernandez de Oviedo’s Historia general y natural de las Indias. Science in Context, 16(4):489–504, December 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lopez-Ocon:2003:OOT

[494] Leoncio López-Ocón and Sara Badía. Overcoming obstacles: The triple mobilization of the Comisión Científica del Pacífico. Science in Context, 16(4):505–534, December 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bazhanov:2003:SWE

[495] Valentin A. Bazhanov. The scholar and the “wolfhound era”: The fate of Ivan E. Orlov’s ideas in logic, philosophy, and science. Science in Context, 16(4):535–550, December 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Eilam:2003:PFA

[496] Gavriela Eilam. The philosophical foundations of Aleksandr R. Luria’s neuropsychology. Science in Context, 16(4):551–577, December 2003. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:2004:IHM

[497] Leo Corry. Introduction: The history of modern mathematics — writing and rewriting. Science in Context, 17(1–2):1–21, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gray:2004:AAN

[498] Jeremy J. Gray. Anxiety and abstraction in Nineteenth-Century mathematics. Science in Context, 17(1–2):23–47, June 2004. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ferreiros:2004:MBC

[499] José Ferreirós. The motives behind Cantor’s set theory — physical, biological, and philosophical questions. Science in Context, 17(1–2):49–83, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rowe:2004:MMO

[500] David E. Rowe. Making mathematics in an oral culture: Göttingen in the era of Klein and Hilbert. Science in Context, 17(1–2):85–129, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Epple:2004:KIV

[501] Moritz Epple. Knot invariants in Vienna and Princeton during the 1920s: Epistemic configurations of mathematical research. Science in Context, 17(1–2):131–164, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Scholz:2004:HWA

[502] Erhard Scholz. Hermann Weyl’s analysis of the “Problem of Space” and the origin of gauge structures. Science in Context, 17(1–2):165– 197, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Katz:2004:BRZ

[503] Shaul Katz. Berlin roots — Zionist incarnation: The ethos of pure mathematics and the beginnings of the Einstein Institute of Mathematics at the Hebrew University of Jerusalem. Science in Context, 17(1–2):199– 234, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dalmedico:2004:EDN

[504] Amy Dahan Dalmedico. Early developments of nonlinear science in Soviet Russia: The Andronov School at Gor’kiy. Science in Context, 17 (1–2):235–265, June 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Giaever:2004:PQN

[505] Øyvind Giæver. The psychiatry of Quislingism: Norwegian psychiatric research on the collaborators of World War II. Science in Context, 17(3): 267–292, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wesseling:2004:JRH

[506] Elisabeth Wesseling. Judith Rich Harris: The Miss Marple of developmental psychology. Science in Context, 17(3):293–314, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Koch:2004:MER

[507] Lene Koch. The meaning of eugenics: Reflections on the government of genetic knowledge in the past and the present. Science in Context, 17(3): 315–331, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Siegmund-Schultze:2004:NCL

[508] Reinhard Siegmund-Schultze. A non-conformist longing for unity in the fractures of modernity: Towards a scientific biography of Richard von Mises (1883–1953). Science in Context, 17(3):333–370, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2004:EN

[509] Anonymous. Editors’ note. Science in Context, 17(3):371, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Siegmund-Schultze:2004:MFP

[510] Reinhard Siegmund-Schultze. Mathematicians forced to philosophize: An introduction to Khinchin’s paper on von Mises’ theory of probability. Science in Context, 17(3):373–390, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2004:CTT

[511] Anonymous. Classical text in translation. Science in Context, 17(3): 391–422, September 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Henderson:2004:EIW

[512] Linda Dalrymple Henderson. Editor’s introduction: I. Writing modern art and science — an overview; II. Cubism, futurism, and ether physics in the early Twentieth Century. Science in Context, 17(4):423–466, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Geimer:2004:PBB

[513] Peter Geimer. Picturing the black box: On blanks in Nineteenth Century paintings and photographs. Science in Context, 17(4):467–501, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Larson:2004:ORP

[514] Barbara Larson. Odilon Redon and the Pasteurian revolution: Health, illness, and le monde invisible. Science in Context, 17(4):503–524, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Botar:2004:LMN

[515] Oliver A. I. Botar. László Moholy-Nagy’s new vision and the aestheticization of scientific photography in Weimar Germany. Science in Context, 17(4):525–556, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Parkinson:2004:SQM

[516] Gavin Parkinson. Surrealism and quantum mechanics: Dispersal and fragmentation in art, life, and physics. Science in Context, 17(4):557–577, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Petersen:2004:EPA

[517] Stephen Petersen. Explosive propositions: Artists react to the atomic age. Science in Context, 17(4):579–609, December 2004. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goodyear:2004:GKB

[518] Anne Collins Goodyear. Gyorgy Kepes, Billy Klüver, and American art of the 1960s: Defining attitudes toward science and technology. Science in Context, 17(4):611–635, December 2004. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Netz:2005:ESU

[519] Reviel Netz and Serafina Cuomo. Early science: a universal history of particulars. Science in Context, 18(1):1–6, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hsu:2005:TBE

[520] Elisabeth Hsu. Tactility and the body in early Chinese medicine. Science in Context, 18(1):7–34, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Burnyeat:2005:AO

[521] M. F. Burnyeat. Archytas and optics. Science in Context, 18(1):35–53, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Vallance:2005:MMH

[522] John Vallance. Minerals, morals, and the hidden purpose of nature. Science in Context, 18(1):55–68, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wardy:2005:MAO

[523] Robert Wardy. The mysterious Aristotelian olive. Science in Context, 18(1):69–91, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sharples:2005:STA

[524] R. W. Sharples. Some thoughts on Aristotelian form: With special reference to Metaphysics Z 8. Science in Context, 18(1):93–109, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Nutton:2005:FEG

[525] Vivian Nutton. The fatal embrace: Galen and the history of ancient medicine. Science in Context, 18(1):111–121, March 2005. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Chemla:2005:GFG

[526] Karine Chemla. Geometrical figures and generality in Ancient China and beyond: Liu Hui and Zhao Shuang, Plato and Thabit ibn Qurra. Science in Context, 18(1):123–166, March 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See corrigendum [664].

Lloyd:2005:PCH

[527] Geoffrey Lloyd. The potential of the comparative history of pre-modern science. Science in Context, 18(1):167–172, March 2005. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:2005:ISL

[528] Leo Corry. Introduction: Science in Latin–American contexts — historical studies. Science in Context, 18(2):173–178, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Denis:2005:MSR

[529] Adrián López Denis. Melancholia, slavery, and racial pathology in Eighteenth-Century Cuba. Science in Context, 18(2):179–199, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lopes:2005:SCM

[530] Maria Margaret Lopes, Clarete Paranhos da Silva, Silvia Fernanda de M. Figueirôa, and Rachel Pinheiro. Scientific culture and mineralogical sciences in the Luso-Brazilian Empire: The work of João da Silva Feijó (1760–1824) in Ceará. Science in Context, 18(2):201–224, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

deSantana:2005:NSB

[531] José Carlos Barreto de Santana. Natural science and Brazilian nationality: Os sertões by Euclides da Cunha. Science in Context, 18(2):225–247, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Podgorny:2005:BDC

[532] Irina Podgorny. Bones and devices in the constitution of paleontology in Argentina at the end of the Nineteenth Century. Science in Context, 18(2):249–283, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

deMendoza:2005:AER

[533] Diego Hurtado de Mendoza. Autonomy, even regional hegemony: Argentina and the “hard way” toward its first research reactor (1945–1958). Science in Context, 18(2):285–308, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

deMendoza:2005:PRL

[534] Diego Hurtado de Mendoza and Miguel de Asúa. The poetry of Relativity: Leopoldo Lugones’ The Size of Space. Science in Context, 18 (2):309–315, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lugones:2005:BSS

[535] Leopoldo Lugones. The Size of Space (an essay on mathematical psychology) translated by Miguel de Asúa and Diego Hurtado de Mendoza. Science in Context, 18(2):317–336, June 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mohun:2005:FBE

[536] Arwen Mohun. On the frontier of The Empire of Chance: Statistics, accidents, and risk in industrializing America. Science in Context, 18(3): 337–357, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Nair:2005:ENE

[537] Savithri Preetha Nair. “Eyes and no eyes”: Siwalik fossil collecting and the crafting of Indian palaeontology (1830–1847). Science in Context, 18(3):359–392, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Huang:2005:TZC

[538] Xiang Huang. The trading zone communication of scientific knowledge: An examination of Jesuit science in China (1582–1773). Science in Context, 18(3):393–427, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

McLarty:2005:PTB

[539] Colin McLarty. Poor taste as a bright character trait: Emmy Noether and the Independent Social Democratic Party. Science in Context, 18(3): 429–450, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schemmel:2005:ARG

[540] Matthias Schemmel. An astronomical road to General Relativity: The continuity between classical and relativistic cosmology in the work of Karl Schwarzschild. Science in Context, 18(3):451–478, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Israel:2005:SCE

[541] Giorgio Israel. The science of complexity: Epistemological problems and perspectives. Science in Context, 18(3):479–509, September 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Naumann:2005:ISL

[542] Barbara Naumann. Introduction: Science and literature. Science in Context, 18(4):511–523, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Torra-Mattenklott:2005:FFC

[543] Caroline Torra-Mattenklott. The fable as figure: Christian Wolff’s geometric fable theory and its creative reception by Lessing and Herder. Science in Context, 18(4):525–552, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stephenson:2005:BSG

[544] Roger H. Stephenson. “Binary synthesis”: Goethe’s aesthetic intuition in literature and science. Science in Context, 18(4):553–581, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ender:2005:DVM

[545] Evelyne Ender. “Déjà vu” or memory-science between Gérard de Nerval and Marcel Proust. Science in Context, 18(4):583–606, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stadler:2005:MOM

[546] Ulrich Stadler. Moving objects, moved observers: On the treatment of the problem of Relativity in poetic texts and scientific prose. Science in Context, 18(4):607–627, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Brandt:2005:GCT

[547] Christina Brandt. Genetic code, text, and scripture: Metaphors and narration in German molecular biology. Science in Context, 18(4):629– 648, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Strowick:2005:CES

[548] Elisabeth Strowick. Comparative epistemology of suspicion: Psychoanalysis, literature, and the human sciences. Science in Context, 18(4): 649–669, December 2005. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Marinelli:2006:EIF

[549] Lydia Marinelli and Andreas Mayer. Editors’ introduction: Forgetting freud? For a new historiography of psychoanalysis. Science in Context, 19(1):1–13, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carroy:2006:DSS

[550] Jacqueline Carroy. Dreaming scientists and scientific dreamers: Freud as a reader of French dream literature. Science in Context, 19(1):15–35, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mayer:2006:LOL

[551] Andreas Mayer. Lost objects: From the laboratories of hypnosis to the psychoanalytic setting. Science in Context, 19(1):37–64, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Forrester:2006:RFF

[552] John Forrester. Remembering and forgetting Freud in early Twentieth– Century dreams. Science in Context, 19(1):65–85, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Marinelli:2006:SWT

[553] Lydia Marinelli. Screening wish theories: Dream psychologies and early cinema. Science in Context, 19(1):87–110, March 2006. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Winter:2006:FCM

[554] Alison Winter. Film and the construction of memory in psychoanalysis, 1940–1960. Science in Context, 19(1):111–136, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Leys:2006:IT

[555] Ruth Leys. Image and trauma. Science in Context, 19(1):137–149, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lunbeck:2006:BHP

[556] Elizabeth Lunbeck. Borderline histories: Psychoanalysis inside and out. Science in Context, 19(1):151–173, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Young:2006:REF

[557] Allan Young. Remembering the evolutionary Freud. Science in Context, 19(1):175–189, March 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Abimbola:2006:CMN

[558] Kólá Abímbólá. A critique of methodological naturalism. Science in Context, 19(2):191–213, June 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Falk:2006:MI

[559] Raphael Falk. Mendel’s impact. Science in Context, 19(2):215–236, June 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Tsovel:2006:USC

[560] Ariel Tsovel. The untold story of a chicken and the missing knowledge in interspecific ethics. Science in Context, 19(2):237–267, June 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Suplizio:2006:EPA

[561] Jean Suplizio. Evolutionary psychology: The academic debate. Science in Context, 19(2):269–293, June 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Magalhaes:2006:RNA

[562] Gildo Magalhães. Remarks on a new autograph letter from Augustin Fresnel: Light aberration and wave theory. Science in Context, 19(2): 295–307, June 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Davidovitch:2006:MBH

[563] Nadav Davidovitch and Rakefet Zalashik. Medical borders: Historical, political, and cultural analyses. Science in Context, 19(3):309–316, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Maglen:2006:MSC

[564] Krista Maglen. “In This Miserable Spot Called Quarantine”: The healthy and unhealthy in Nineteenth Century Australian and Pacific quarantine stations. Science in Context, 19(3):317–336, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fairchild:2006:RFM

[565] Amy L. Fairchild. The rise and fall of the medical gaze: The political economy of immigrant medical inspection in modern America. Science in Context, 19(3):337–356, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rodriguez:2006:IAB

[566] Julia Rodriguez. Inoculating against barbarism? State medicine and immigrant policy in turn-of-the-century Argentina. Science in Context, 19(3):357–380, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sufian:2006:CMM

[567] Sandy Sufian. Colonial malariology, medical borders, and sharing scientific knowledge in mandatory Palestine. Science in Context, 19(3): 381–400, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weindling:2006:BLB

[568] Paul Weindling. “Belsenitis”: Liberating Belsen, its hospitals, UNRRA, and selection for re-emigration, 1945–1948. Science in Context, 19(3): 401–418, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zalashik:2006:MAP

[569] Rakefet Zalashik and Nadav Davidovitch. Measuring adaptability: Psychological examinations of Jewish detainees in Cyprus internment camps. Science in Context, 19(3):419–441, September 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gross:2006:VVS

[570] Alan G. Gross. The verbal and the visual in science: a Heideggerian perspective. Science in Context, 19(4):443–474, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hoon:2006:MCC

[571] Jun Yong Hoon. Mathematics in context: a case in early Nineteenth-Century Korea. Science in Context, 19(4):475–512, December 2006. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schweber:2006:EOI

[572] Silvan S. Schweber. Einstein and Oppenheimer: Interactions and intersections. Science in Context, 19(4):513–559, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Link:2006:CWM

[573] David Link. Chains to the West: Markov’s theory of connected events and its transmission to Western Europe. Science in Context, 19(4):561– 589, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Markov:2006:ESI

[574] A. A. Markov. An example of statistical investigation of the text Eugene Onegin concerning the connection of samples in chains. Science in Context, 19(4):591–600, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Markov:2006:RCS

[575] A. A. Markov. On a remarkable case of samples connected in a chain. Appendix on the statistical investigation of a text by Aksakov. Science in Context, 19(4):601–604, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Blank:2007:CSC

[576] Andreas Blank. Composite substance, common notions, and Kenelm Digby’s theory of animal generation. Science in Context, 20(1):1–20, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Yeo:2007:BMR

[577] Richard Yeo. Before Memex: Robert Hooke, John Locke, and Vannevar Bush on external memory. Science in Context, 20(1):21–47, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Harman:2007:PIR

[578] Oren Solomon Harman. Powerful intuitions: Re-reading nature versus nurture with Charles Darwin and Clifford Geertz. Science in Context, 20(1):49–70, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Reynolds:2007:TCS

[579] Andrew Reynolds. The theory of the Cell state and the question of Cell autonomy in nineteenth and early Twentieth–Century biology. Science in Context, 20(1):71–95, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Loettgers:2007:GAM

[580] Andrea Loettgers. Getting abstract mathematical models in touch with nature. Science in Context, 20(1):97–124, March 2007. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Urken:2007:NCD

[581] Arnold B. Urken and Iain McLean. Nicholas Collin and the dissemination of Condorcet in the United States. Science in Context, 20(1):125–133, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Prosperin:2007:HDP

[582] Erik Prosperin and Nicolaus Collin. Historical document: On probable knowledge. Science in Context, 20(1):135–143, March 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:2007:ORH

[583] Alexandre Métraux. Opening remarks on the history of science in Yiddish. Science in Context, 20(2):145–162, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zwiep:2007:ARV

[584] Irene E. Zwiep. Adding the Reader’s voice: Early-modern Ashkenazi grammars of Hebrew. Science in Context, 20(2):163–195, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Shneer:2007:SRJ

[585] David Shneer. A study in red: Jewish scholarship in the 1920s Soviet union. Science in Context, 20(2):197–213, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Estraikh:2007:JLY

[586] Gennady Estraikh. Jacob Lestschinsky: a Yiddishist dreamer and social scientist. Science in Context, 20(2):215–237, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Yalen:2007:SEF

[587] Deborah Yalen. “On the social-economic front”: The polemics of shtetl research during the Stalin Revolution. Science in Context, 20(2):239– 301, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Berger:2007:IFY

[588] Shlomo Berger. Interpreting Freud: The Yiddish philosophical journal Davke investigates a Jewish icon. Science in Context, 20(2):303–316, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gruschka:2007:TSB

[589] Roland Gruschka. Tuvia Schalit’s Di spetsyele relativitets-teorye of 1927 and other introductions to the Theory of Relativity in Yiddish. Science in Context, 20(2):317–339, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Trachtenberg:2007:BBB

[590] Barry Trachtenberg. Ber Borochov’s “The Tasks of Yiddish Philology”. Science in Context, 20(2):341–352, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2007:EN

[591] Anonymous. Editor’s note. Science in Context, 20(2):353, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Borochov:2007:CTT

[592] Ber Borochov. Classical text in translation: The tasks of Yiddish philology. Science in Context, 20(2):355–373, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2007:A

[593] Anonymous. Appendix. Science in Context, 20(2):375–379, June 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Vidal:2007:IKB

[594] Fernando Vidal and Bernhard Kleeberg. Introduction: Knowledge, belief, and the impulse to natural theology. Science in Context, 20(3): 381–400, September 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Feldhay:2007:TEF

[595] Rivka Feldhay. Thomist epistemology of faith: The road from “scientia” to science. Science in Context, 20(3):401–421, September 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Smoller:2007:ASJ

[596] Laura Ackerman Smoller. Astrology and the Sibyls: John of Legnano’s De adventu Christi and the natural theology of the Later Middle Ages. Science in Context, 20(3):423–450, September 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mandelbrote:2007:UNT

[597] Scott Mandelbrote. The uses of natural theology in Seventeenth-Century England. Science in Context, 20(3):451–480, September 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Vidal:2007:MST

[598] Fernando Vidal. Miracles, science, and testimony in post-Tridentine saint-making. Science in Context, 20(3):481–508, September 2007. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Richards:2007:SSS

[599] Joan L. Richards. In search of the “sea-something”: Reason and transcendence in the Frend/De Morgan family. Science in Context, 20(3): 509–536, September 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kleeberg:2007:GNP

[600] Bernhard Kleeberg. God-nature progressing: Natural theology in German monism. Science in Context, 20(3):537–569, September 2007. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Maglo:2007:FMP

[601] Koffi Maglo. Force, mathematics, and physics in Newton’s Principia: a new approach to enduring issues. Science in Context, 20(4):571–600, December 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lee:2007:SAM

[602] Paula Young Lee. The social architect and the myopic mason: The spatial politics of the Muséum d’Histoire naturelle in Nineteenth-Century Paris. Science in Context, 20(4):601–625, December 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lachapelle:2007:EIU

[603] Sofie Lachapelle. Educating idiots: Utopian ideals and practical organization regarding idiocy inside Nineteenth-Century French asylums. Science in Context, 20(4):627–648, December 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Diaz:2007:RIM

[604] Edna Suárez Díaz. The rhetoric of informational molecules: Authority and promises in the early study of molecular evolution. Science in Context, 20(4):649–677, December 2007. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Pagliaro:2007:SUV

[605] Mario Pagliaro. “Of sea urchins, volcanoes, earthquakes … and engagement”: Marcello Carapezza, Alberto Monroy, and Italy’s university system. Science in Context, 20(4):679–691, December 2007. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Elazar:2008:HFT

[606] Michael Elazar. Honoré Fabri and the Trojan Horse of inertia. Science in Context, 21(1):1–38, March 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

deAsua:2008:NWL

[607] Miguel de Asúa. ‘Names which he loved, and things well worthy to be known’: Eighteenth-Century Jesuit natural histories of Paraquaria and Río de la Plata. Science in Context, 21(1):39–72, March 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lynn:2008:CRB

[608] Michael R. Lynn. Consumerism and the rise of balloons in Europe at the end of the Eighteenth Century. Science in Context, 21(1):73–98, March 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Burnham:2008:APU

[609] John C. Burnham. Accident proneness (Unfallneigung): a classic case of simultaneous discovery/construction in psychology. Science in Context, 21(1):99–118, March 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Akavia:2008:WBC

[610] Naamah Akavia. Writing ‘The Case of Ellen West’: Clinical knowledge and historical representation. Science in Context, 21(1):119–144, March 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gradmann:2008:LTV

[611] Christoph Gradmann. Locating therapeutic vaccines in Nineteenth-Century history. Science in Context, 21(2):145–160, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kloppel:2008:ECB

[612] Ulrike Klöppel. Enacting cultural boundaries in French and German diphtheria serum research. Science in Context, 21(2):161–180, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Simon:2008:MSP

[613] Jonathan Simon. Monitoring the stable at the Pasteur Institute. Science in Context, 21(2):181–200, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hess:2008:ASV

[614] Volker Hess. The administrative stabilization of vaccines: Regulating the diphtheria antitoxin in France and Germany, 1894–1900. Science in Context, 21(2):201–227, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Huntelmann:2008:DW

[615] Axel C. Hüntelmann. The dynamics of Wertbestimmung. Science in Context, 21(2):229–252, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cassier:2008:PCS

[616] Maurice Cassier. Producing, controlling, and stabilizing Pasteur’s anthrax vaccine: Creating a new industry and a health market. Science in Context, 21(2):253–278, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bonah:2008:PBS

[617] Christian Bonah. Packaging BCG: Standardizing an anti-tuberculosis vaccine in interwar Europe. Science in Context, 21(2):279–310, June 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Goldschmidt:2008:CMB

[618] Asaf Goldschmidt. Commercializing medicine or benefiting the people — the first public pharmacy in China. Science in Context, 21(3):311–350, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wigelsworth:2008:NNA

[619] Jeffrey R. Wigelsworth. Navigation and newsprint: Advertising longitude schemes in the public sphere ca. 1715. Science in Context, 21(3):351–376, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

VanDijck:2008:SPBa

[620] Maarten Van Dijck. From science to popularization, and back — the science and journalism of the Belgian economist Gustave de Molinari. Science in Context, 21(3):377–402, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See corrigendum [633].

Dodier:2008:AOP

[621] Nicolas Dodier and Janine Barbot. Autonomy and objectivity as political operators in the medical world: Twenty years of public controversy about AIDS treatments in France. Science in Context, 21(3):403–434, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gangui:2008:EUO

[622] Alejandro Gangui and Eduardo L. Ortiz. Einstein’s unpublished opening lecture for his course on Relativity Theory in Argentina, 1925. Science in Context, 21(3):435–450, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Einstein:2008:UOL

[623] Albert Einstein. Unpublished opening lecture for the course on the Theory of Relativity in Argentina, 1925 (reprinted). Science in Context, 21(3):451–459, September 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wolfe:2008:IVM

[624] Charles T. Wolfe. Introduction: Vitalism without metaphysics? Medical vitalism in the enlightenment. Science in Context, 21(4):461–463, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Giglioni:2008:WEH

[625] Guido Giglioni. What ever happened to Francis Glisson? Albrecht Haller and the fate of Eighteenth-Century irritability. Science in Context, 21(4): 465–493, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Cheung:2008:RAF

[626] Tobias Cheung. Regulating agents, functional interactions, and stimulusreaction-schemes: The concept of ‘organism’ in the organic system theories of Stahl, Bordeu, and Barthez. Science in Context, 21(4):495–519, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Boury:2008:ISK

[627] Dominique Boury. Irritability and sensibility: Key concepts in assessing the medical doctrines of Haller and Bordeu. Science in Context, 21(4): 521–535, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wolfe:2008:AEO

[628] Charles T. Wolfe and Motoichi Terada. The animal economy as object and program in Montpellier vitalism. Science in Context, 21(4):537–579, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kaitaro:2008:CMM

[629] Timo Kaitaro. Can matter mark the hours? Eighteenth-Century vitalist materialism and functional properties. Science in Context, 21(4):581– 592, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Williams:2008:TLO

[630] Elizabeth A. Williams. Of two lives one? Jean-Charles-Marguerite-Guillaume Grimaud and the question of holism in vitalist medicine. Science in Context, 21(4):593–613, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Huneman:2008:MVE

[631] Philippe Huneman. Montpellier vitalism and the emergence of alienism in France (1750–1800): The case of the passions. Science in Context, 21(4):615–647, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Witt:2008:FMG

[632] Elke Witt. Form — a matter of generation: The relation of generation, form, and function in the epigenetic theory of Caspar F. Wolff. Science in Context, 21(4):649–664, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

VanDijck:2008:SPBb

[633] Maarten Van Dijck. From science to popularization, and back — the science and journalism of the Belgian economist Gustave de Molinari — corrigendum. Science in Context, 21(4):665, December 2008. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See [620].

Anonymous:2009:JUT

[634] Anonymous. Journals under threat: a joint response from history of science, technology, and medicine Editors. Science in Context, 22(1):1– 4, March 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rochberg:2009:CIP

[635] Francesca Rochberg. Conditionals, inference, and possibility in Ancient Mesopotamian science. Science in Context, 22(1):5–25, March 2009. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schibille:2009:AOP

[636] Nadine Schibille. Astronomical and optical principles in the architecture of Hagia Sophia in Constantinople. Science in Context, 22(1):27–46, March 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hunter:2009:GCE

[637] Patti W. Hunter. Gertrude Cox in Egypt: a case study in science patronage and international statistics education during the Cold War. Science in Context, 22(1):47–83, March 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kjeldsen:2009:EFM

[638] Tinne Hoff Kjeldsen. Egg-forms and measure-bodies: Different mathematical practices in the early history of the modern theory of convexity. Science in Context, 22(1):85–113, March 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mata:2009:MBW

[639] Tiago Mata. Migrations and boundary work: Harvard, radical economists, and the committee on political discrimination. Science in Context, 22(1):115–143, March 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schmitt:2009:PCC

[640] Stéphane Schmitt. From physiology to classification: Comparative anatomy and Vicq d’Azyr’s plan of reform for life sciences and medicine (1774–1794). Science in Context, 22(2):145–193, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

VanReybrouck:2009:MRS

[641] David Van Reybrouck, Raf de Bont, and Jan Rock. Material rhetoric: Spreading stones and showing bones in the study of prehistory. Science in Context, 22(2):195–216, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Novoa:2009:APD

[642] Adriana Novoa. The act or process of dying out: The importance of Darwinian extinction in Argentine culture. Science in Context, 22(2): 217–244, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Thomas:2009:ENF

[643] William Thomas and Lambert Williams. The epistemologies of nonforecasting simulations, Part I: Industrial dynamics and management pedagogy at MIT. Science in Context, 22(2):245–270, June 2009. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Williams:2009:ENF

[644] Lambert Williams and William Thomas. The epistemologies of nonforecasting simulations, Part II: Climate, chaos, computing style, and the contextual plasticity of error. Science in Context, 22(2):271–310, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:SVIa

[645] Anonymous. SIC volume 22 issue 2 cover and front matter. Science in Context, 22(2):f1–f3, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:SVIb

[646] Anonymous. SIC volume 22 issue 2 cover and back matter. Science in Context, 22(2):b1–b10, June 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bigg:2009:ILN

[647] Charlotte Bigg, David Aubin, and Philipp Felsch. Introduction: The laboratory of nature — science in the mountains. Science in Context, 22(3):311–321, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Reeves:2009:DEI

[648] Nicky Reeves. ‘To demonstrate the exactness of the instrument’: Mountainside trials of precision in Scotland, 1774. Science in Context, 22(3): 323–340, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Felsch:2009:MSM

[649] Philipp Felsch. Mountains of sublimity, mountains of fatigue: Towards a history of speechlessness in the Alps. Science in Context, 22(3):341–364, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Aubin:2009:HBO

[650] David Aubin. The hotel that became an observatory: Mount Faulhorn as singularity, microcosm, and macro-tool. Science in Context, 22(3): 365–386, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Speich:2009:MMS

[651] Daniel Speich. Mountains made in Switzerland: Facts and concerns in Nineteenth-Century cartography. Science in Context, 22(3):387–408, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vonBrevern:2009:CUA

[652] Jan von Brevern. Counting on the unexpected: Aimé Civiale’s mountain photography. Science in Context, 22(3):409–437, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Westermann:2009:ITG

[653] Andrea Westermann. Inherited territories: The Glarus Alps, knowledge validation, and the genealogical organization of Nineteenth-Century Swiss alpine geognosy. Science in Context, 22(3):439–461, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Coen:2009:SLM

[654] Deborah R. Coen. The storm lab: Meteorology in the Austrian Alps. Science in Context, 22(3):463–486, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Becker:2009:PP

[655] Catherine Nisbett Becker. Professionals on the peak. Science in Context, 22(3):487–507, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

LeGars:2009:EPM

[656] Stéphane Le Gars and David Aubin. The elusive placelessness of the Mont-Blanc Observatory (1893–1909): The social underpinnings of highaltitude observation. Science in Context, 22(3):509–531, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:SVIc

[657] Anonymous. SIC volume 22 issue 3 cover and front matter. Science in Context, 22(3):f1–f3, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:SVId

[658] Anonymous. SIC volume 22 issue 3 cover and back matter. Science in Context, 22(3):b1–b6, September 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:E

[659] Anonymous. Editorial. Science in Context, 22(4):533, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wagner:2009:SHG

[660] Roy Wagner. For some histories of Greek mathematics. Science in Context, 22(4):535–565, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Beck:2009:SPS

[661] Naomi Beck. In search of the proper scientific approach: Hayek’s views on biology, methodology, and the nature of economics. Science in Context, 22(4):567–585, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kragh:2009:CFO

[662] Helge Kragh. Continual fascination: The oscillating universe in modern cosmology. Science in Context, 22(4):587–612, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Oleinik:2009:ICS

[663] Anton Oleinik. Inquiring into communication in science: Alternative approaches. Science in Context, 22(4):613–646, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:GFG

[664] Anonymous. Geometrical figures and generality in Ancient China and beyond: Liu Hui and Zhao Shuang, Plato and Thabit ibn Qurra — Corrigendum. Science in Context, 22(4):647–650, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See [526].

Anonymous:2009:SVIe

[665] Anonymous. SIC volume 22 issue 4 cover and front matter. Science in Context, 22(4):f1–f3, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2009:SVIf

[666] Anonymous. SIC volume 22 issue 4 cover and back matter. Science in Context, 22(4):b1–b10, December 2009. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Acerbi:2010:HLG

[667] Fabio Acerbi. Homeomeric lines in Greek mathematics. Science in Context, 23(1):1–37, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Thorvaldsen:2010:ENA

[668] Steinar Thorvaldsen. Early numerical analysis in Kepler’s new astronomy. Science in Context, 23(1):39–63, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bullynck:2010:JHL

[669] Maarten Bullynck. Johann Heinrich Lambert’s scientific tool kit, exemplified by his measurement of humidity, 1769–1772. Science in Context, 23(1):65–89, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ehrhardt:2010:SHG

[670] Caroline Ehrhardt. A social history of the ‘Galois Affair’ at the Paris Academy of Sciences (1831). Science in Context, 23(1):91–119, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Jackson:2010:DAE

[671] John P. Jackson. Definitional argument in evolutionary psychology and cultural anthropology. Science in Context, 23(1):121–150, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVIa

[672] Anonymous. SIC volume 23 issue 1 cover and front matter. Science in Context, 23(1):f1–f3, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVIb

[673] Anonymous. SIC volume 23 issue 1 cover and back matter. Science in Context, 23(1):b1–b10, March 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Acerbi:2010:TAF

[674] Fabio Acerbi. Two approaches to foundations in Greek mathematics: Apollonius and Geminus. Science in Context, 23(2):151–186, June 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Stapleford:2010:SKA

[675] Thomas A. Stapleford. Shaping knowledge about American labor: External advising at the U.S. Bureau of Labor Statistics in the Twentieth Century. Science in Context, 23(2):187–220, June 2010. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fontaine:2010:SAS

[676] Philippe Fontaine. Stabilizing American society: Kenneth Boulding and the integration of the social sciences, 1943–1980. Science in Context, 23 (2):221–265, June 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVIc

[677] Anonymous. SIC volume 23 issue 2 cover and front matter. Science in Context, 23(2):f1–f3, June 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVId

[678] Anonymous. SIC volume 23 issue 2 cover and back matter. Science in Context, 23(2):b1–b4, June 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carvallo:2010:ASE

[679] Sarah Carvallo. Ageing in the Seventeenth and Eighteenth Centuries. Science in Context, 23(3):267–288, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Whitmer:2010:MNM

[680] Kelly J. Whitmer. The model that never moved: The case of a virtual memory theater and its Christian philosophical argument, 1700–1732. Science in Context, 23(3):289–327, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Clericuzio:2010:SEN

[681] Antonio Clericuzio. ‘Sooty Empiricks’ and natural philosophers: The status of chemistry in the Seventeenth Century. Science in Context, 23(3):329–350, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Dear:2010:DIM

[682] Peter Dear. Divine illumination, mechanical calculators, and the roots of modern reason. Science in Context, 23(3):351–366, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sullivan:2010:TTO

[683] Gregory Sullivan. Tricks of transference: Oka Asajiro (1868–1944) on laissez-faire capitalism. Science in Context, 23(3):367–391, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVIe

[684] Anonymous. SIC volume 23 issue 3 cover and front matter. Science in Context, 23(3):f1–f3, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2010:SVIf

[685] Anonymous. SIC volume 23 issue 3 cover and back matter. Science in Context, 23(3):b1–b10, September 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:2010:I

[686] Leo Corry and Tal Golan. Introduction. Science in Context, 23(4): 393–399, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Davidovitch:2010:PPP

[687] Nadav Davidovitch and Rakefet Zalashik. Pasteur in Palestine: The politics of the laboratory. Science in Context, 23(4):401–425, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Corry:2010:ZIT

[688] Leo Corry and Norbert Schappacher. Zionist internationalism through number theory: Edmund Landau at the opening of the Hebrew University in 1925. Science in Context, 23(4):427–471, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Rolnik:2010:TIP

[689] Eran J. Rolnik. Therapy and ideology: Psychoanalysis and its vicissitudes in pre-state Israel (including some hitherto unpublished letters by Sigmund Freud and Albert Einstein). Science in Context, 23(4):473–506, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Krampf:2010:EPF

[690] Arie Krampf. Economic planning of the free market in Israel during the first decade: The influence of Don Patinkin on Israeli policy discourse. Science in Context, 23(4):507–534, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Golan:2010:KAS

[691] Tal Golan. The Kishon Affair: Science, law, and the politics of causation. Science in Context, 23(4):535–569, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Finkelstein:2010:ERN

[692] Guy Finkelstein and Alexandre Métraux. On emanuel Ringelblum’s new research program for the history of Jewish medicine: Introductory remarks. Science in Context, 23(4):571–580, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Ringelblum:2010:SPS

[693] Emanuel Ringelblum. Some proposals submitted to the world Congress of Jewish physicians. Science in Context, 23(4):581–586, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Anonymous:2010:SVIg

[694] Anonymous. SIC volume 23 issue 4 cover and front matter. Science in Context, 23(4):f1–f3, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Anonymous:2010:SVIh

[695] Anonymous. SIC volume 23 issue 4 cover and back matter. Science in Context, 23(4):b1–b6, December 2010. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science in an Israeli Context: Case Studies.

Guttler:2011:SPE

[696] Nils Robert Güttler. Scaling the period eye: Oscar Drude and the cartographical practice of plant geography, 1870s–1910s. Science in Context, 24(1):1–41, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gasca:2011:MNS

[697] Ana Millán Gasca. Mathematicians and the nation in the second half of the Nineteenth Century as reflected in the Luigi Cremona correspondence. Science in Context, 24(1):43–72, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gispert:2011:SSC

[698] Hélène Gispert and Gert Schubring. Societal, structural, and conceptual changes in mathematics teaching: Reform processes in France and Germany over the Twentieth Century and the international dynamics. Science in Context, 24(1):73–106, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bragesjo:2011:DVV

[699] Fredrik Bragesjö and Margareta Hallberg. Dilemmas of a vitalizing vaccine market: Lessons from the MMR vaccine/autism debate. Science in Context, 24(1):107–125, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2011:SVIa

[700] Anonymous. SIC volume 24 issue 1 cover and front matter. Science in Context, 24(1):f1–f3, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2011:SVIb

[701] Anonymous. SIC volume 24 issue 1 cover and back matter. Science in Context, 24(1):b1–b10, March 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Vetter:2011:ILP

[702] Jeremy Vetter. Introduction: Lay participation in the history of scientific observation. Science in Context, 24(2):127–141, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Brenna:2011:CAF

[703] Brita Brenna. Clergymen abiding in the fields: The making of the naturalist observer in Eighteenth-Century Norwegian natural history. Science in Context, 24(2):143–166, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Werrett:2011:WFE

[704] Simon Werrett. Watching the fireworks: Early modern observation of natural and artificial spectacles. Science in Context, 24(2):167–182, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Hochadel:2011:WEA

[705] Oliver Hochadel. Watching exotic animals next door: ‘scientific’ observations at the zoo (ca. 1870–1910). Science in Context, 24(2):183–214, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Cain:2011:AAE

[706] Victoria Cain. The art of authority: Exhibits, exhibit-makers, and the contest for scientific status in the American Museum of Natural History, 1920–1940. Science in Context, 24(2):215–238, June 2011. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Beckman:2011:CST

[707] Jenny Beckman. Collecting standards: Teaching botanical skills in Sweden, 1850–1950. Science in Context, 24(2):239–258, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Vetter:2011:LOT

[708] Jeremy Vetter. Lay observers, telegraph lines, and Kansas weather: The field network as a mode of knowledge production. Science in Context, 24(2):259–280, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Didier:2011:CRI

[709] Emmanuel Didier. Counting on relief: Industrializing the statistical interviewer during the new deal. Science in Context, 24(2):281–310, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Anonymous:2011:SVIc

[710] Anonymous. SIC volume 24 issue 2 cover and front matter. Science in Context, 24(2):f1–f3, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Anonymous:2011:SVId

[711] Anonymous. SIC volume 24 issue 2 cover and back matter. Science in Context, 24(2):b1–b8, June 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Lay Participation in the History of Scientific Observation.

Wellmann:2011:SC

[712] Janina Wellmann. Science and cinema. Science in Context, 24(3):311– 328, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Canales:2011:DMC

[713] Jimena Canales. Desired machines: Cinema and the world in its own image. Science in Context, 24(3):329–359, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Gaycken:2011:SLM

[714] Oliver Gaycken. ‘The swarming of life’: Moving images, education, and views through the microscope. Science in Context, 24(3):361–380, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Landecker:2011:CDD

[715] Hannah Landecker. Creeping, drinking, dying: The cinematic portal and the microscopic world of the Twentieth–Century cell. Science in Context, 24(3):381–416, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Curtis:2011:TTA

[716] Scott Curtis. ‘Tangible as tissue’: Arnold Gesell, infant behavior, and film analysis. Science in Context, 24(3):417–442, September 2011. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Cartwright:2011:HP

[717] Lisa Cartwright. The hands of the projectionist. Science in Context, 24(3):443–464, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Anonymous:2011:SVIe

[718] Anonymous. SIC volume 24 issue 3 cover and front matter. Science in Context, 24(3):f1–f3, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Anonymous:2011:SVIf

[719] Anonymous. SIC volume 24 issue 3 cover and back matter. Science in Context, 24(3):b1–b6, September 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Cinematography, Seriality, and the Sciences.

Levallois:2011:WWB

[720] Clement Levallois. Why were biological analogies in economics ‘a bad thing’? Edith Penrose’s battles against social Darwinism and McCarthyism. Science in Context, 24(4):465–485, December 2011. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Penders:2011:CEF

[721] Bart Penders and Annemiek P. Nelis. Credibility engineering in the food industry: Linking science, regulation, and marketing in a corporate context. Science in Context, 24(4):487–515, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Gross:2011:MDS

[722] Alan G. Gross. A model for the division of semiotic labor in scientific argument: The interaction of words and images. Science in Context, 24(4):517–544, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lamy:2011:WMD

[723] Jérôme Lamy. What management does to space projects: The Franco-Soviet Project ARCAD 3 in the late 1970s. Science in Context, 24(4): 545–586, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mazliak:2011:WDA

[724] Laurent Mazliak and Glenn Shafer. What does the arrest and release of Emile Borel and his colleagues in 1941 tell us about the German occupation of France? Science in Context, 24(4):587–623, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2011:SDB

[725] Anonymous. Some documents bearing on the arrest and release of Emile Borel. Science in Context, 24(4):625–636, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2011:SVIg

[726] Anonymous. SIC volume 24 issue 4 cover and front matter. Science in Context, 24(4):f1–f3, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2011:SVIh

[727] Anonymous. SIC volume 24 issue 4 cover and back matter. Science in Context, 24(4):b1–b12, December 2011. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Coen:2012:IWD

[728] Deborah R. Coen. Introduction: Witness to disaster: Comparative histories of earthquake science and response. Science in Context, 25(1): 1–15, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Valencius:2012:ANM

[729] Conevery Bolton Valencius. Accounts of the New Madrid earthquakes: Personal narratives across two centuries of North American seismology. Science in Context, 25(1):17–48, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

White:2012:DCG

[730] Paul White. Darwin, concepción, and the geological sublime. Science in Context, 25(1):49–71, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Coen:2012:TSN

[731] Deborah R. Coen. The tongues of seismology in Nineteenth-Century Switzerland. Science in Context, 25(1):73–102, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Orihara:2012:NEG

[732] Minami Orihara and Gregory Clancey. The nature of emergency: The Great Kanto Earthquake and the crisis of reason in Late Imperial Japan. Science in Context, 25(1):103–126, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Fan:2012:CMC

[733] Fa ti Fan. ‘Collective monitoring, collective defense’: Science, earthquakes, and politics in Communist China. Science in Context, 25(1): 127–154, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Anonymous:2012:SVIa

[734] Anonymous. SIC volume 25 issue 1 cover and front matter. Science in Context, 25(1):f1–f3, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Anonymous:2012:SVIb

[735] Anonymous. SIC volume 25 issue 1 cover and back matter. Science in Context, 25(1):b1–b6, March 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Witness to Disaster: Earthquakes and Expertise in Comparative Perspective.

Widom:2012:MRT

[736] Theodore Reed Widom and Dirk Schlimm. Methodological reflections on typologies for numerical notations. Science in Context, 25(2):155– 195, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rowe:2012:ERW

[737] David E. Rowe. Einstein and Relativity: What price fame? Science in Context, 25(2):197–246, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rowbottom:2012:RHB

[738] Darrell P. Rowbottom and R. McNeill Alexander. The role of hypotheses in biomechanical research. Science in Context, 25(2):247–262, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Heukelom:2012:SMA

[739] Floris Heukelom. A sense of mission: The Alfred P. Sloan and Russell Sage Foundations’ behavioral economics program, 1984–1992. Science in Context, 25(2):263–286, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2012:SVIc

[740] Anonymous. SIC volume 25 issue 2 cover and front matter. Science in Context, 25(2):f1–f3, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2012:SVId

[741] Anonymous. SIC volume 25 issue 2 cover and back matter. Science in Context, 25(2):b1–b4, June 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lanzoni:2012:IES

[742] Susan Lanzoni. Introduction: Emotion and the sciences: Varieties of empathy in science, art, and history. Science in Context, 25(3):287–300, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Lanzoni:2012:ETM

[743] Susan Lanzoni. Empathy in translation: Movement and image in the psychological laboratory. Science in Context, 25(3):301–327, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Brain:2012:SPH

[744] Robert Michael Brain. Self-projection: Hugo Münsterberg on empathy and oscillation in cinema spectatorship. Science in Context, 25(3):329– 353, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Gallagher:2012:ESN

[745] Shaun Gallagher. Empathy, simulation, and narrative. Science in Context, 25(3):355–381, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Turner:2012:SWE

[746] Stephen Turner. The strength of weak empathy. Science in Context, 25(3):383–399, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Young:2012:SBM

[747] Allan Young. The social brain and the myth of empathy. Science in Context, 25(3):401–424, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Curtis:2012:EAM

[748] Robin Curtis. Einfühlung and abstraction in the moving image: Historical and contemporary reflections. Science in Context, 25(3):425–446, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Frank:2012:LCT

[749] Adam Frank. Loose coordinations: Theater and thinking in gertrude Stein. Science in Context, 25(3):447–467, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Anonymous:2012:SVIe

[750] Anonymous. SIC volume 25 issue 3 cover and front matter. Science in Context, 25(3):f1–f3, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Anonymous:2012:SVIf

[751] Anonymous. SIC volume 25 issue 3 cover and back matter. Science in Context, 25(3):b1–b10, September 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on The Varieties of Empathy in Science, Art, and History.

Mota:2012:AIB

[752] Bernardo Machado Mota. The astronomical interpretation of Catoptrica. Science in Context, 25(4):469–502, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Blank:2012:JCS

[753] Andreas Blank. Julius Caesar Scaliger on plants, species, and the ordained power of God. Science in Context, 25(4):503–523, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sheppard:2012:BSS

[754] Kathleen L. Sheppard. Between spectacle and science: Margaret Murray and the tomb of the two brothers. Science in Context, 25(4):525–549, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kang:2012:SIL

[755] Hyo Yoon Kang. Science inside law: The making of a new patent class in the international patent classification. Science in Context, 25(4):551–594, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2012:SVIg

[756] Anonymous. SIC volume 25 issue 4 cover and front matter. Science in Context, 25(4):f1–f3, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2012:SVIh

[757] Anonymous. SIC volume 25 issue 4 cover and back matter. Science in Context, 25(4):b1–b10, December 2012. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Corry:2013:YE

[758] Leo Corry, Moritz Epple, Orna Harari, Alexandre Métraux, and Jürgen Renn. Yehuda Elkana (1934–2012). Science in Context, 26(1):1–2, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Almasi:2013:TBS

[759] Gábor Almási. Tycho Brahe and the separation of astronomy from astrology: The making of a new scientific discourse. Science in Context, 26(1):3–30, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anduaga:2013:TNS

[760] Aitor Anduaga. Towards a new sphere of practices and knowledge: The militarization of meteorology in Francoist Spain. Science in Context, 26 (1):31–59, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Rangil:2013:CAE

[761] Teresa Tomás Rangil. Citizen, academic, expert, or international worker? Juggling with identities at UNESCO’s Social Science Department, 1946– 1955. Science in Context, 26(1):61–91, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Karaca:2013:SWSa

[762] Koray Karaca. The strong and weak senses of theory-ladenness of experimentation: Theory-driven versus exploratory experiments in the history of high-energy particle physics. Science in Context, 26(1):93–136, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See erratum [797].

Besel:2013:ACC

[763] Richard D. Besel. Accommodating climate change science: James Hansen and the rhetorical/political emergence of global warming. Science in Context, 26(1):137–152, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Roosth:2013:BCC

[764] Sophia Roosth. Biobricks and crocheted coral: Dispatches from the life sciences in the age of fabrication. Science in Context, 26(1):153–171, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weissman:2013:LDT

[765] Charlotte Weissman. On the legitimacy of the Darwinian theory — August Weismann’s first creed. Science in Context, 26(1):173–179, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Weismann:2013:LDT

[766] August Weismann. On the legitimacy of the Darwinian theory. Science in Context, 26(1):181–201, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2013:SVIa

[767] Anonymous. SIC volume 26 issue 1 cover and front matter. Science in Context, 26(1):f1–f3, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2013:SVIb

[768] Anonymous. SIC volume 26 issue 1 cover and back matter. Science in Context, 26(1):b1–b8, March 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Hoffmann:2013:IKM

[769] Christoph Hoffmann and Barbara Wittmann. Introduction: Knowledge in the making: Drawing and writing as research techniques. Science in Context, 26(2):203–213, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Eddy:2013:SKC

[770] Matthew Daniel Eddy. The shape of knowledge: Children and the visual culture of literacy and numeracy. Science in Context, 26(2):215–245, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Nasim:2013:EGT

[771] Omar W. Nasim. Extending the gaze: The temporality of astronomical paperwork. Science in Context, 26(2):247–277, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Hoffmann:2013:PPW

[772] Christoph Hoffmann. Processes on paper: Writing procedures as nonmaterial research devices. Science in Context, 26(2):279–303, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Krauthausen:2013:WLH

[773] Karin Krauthausen. A writer looking for his writing scene: Paul Valéry’s procedures in his notebooks around 1894. Science in Context, 26(2):305– 343, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Barberousse:2013:WUD

[774] Anouk Barberousse. What is the use of diagrams in theoretical modeling? Science in Context, 26(2):345–362, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Wittmann:2013:OSD

[775] Barbara Wittmann. Outlining species: Drawing as a research technique in contemporary biology. Science in Context, 26(2):363–391, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Anonymous:2013:SVIc

[776] Anonymous. SIC volume 26 issue 2 cover and front matter. Science in Context, 26(2):f1–f3, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Anonymous:2013:SVId

[777] Anonymous. SIC volume 26 issue 2 cover and back matter. Science in Context, 26(2):b1–b8, June 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Knowledge in the Making.

Schirrmacher:2013:ICS

[778] Arne Schirrmacher. Introduction: Communicating science: National approaches in Twentieth–Century Europe. Science in Context, 26(3): 393–404, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Govoni:2013:PWC

[779] Paola Govoni. The power of weak competitors: Women scholars, ‘popular science,’ and the building of a scientific community in Italy, 1860s–1930s. Science in Context, 26(3):405–436, September 2013. CODEN SCCOEW.

ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Bowler:2013:PSM

[780] Peter J. Bowler. Popular science magazines in Interwar Britain: Authors and readerships. Science in Context, 26(3):437–457, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Bensaude-Vincent:2013:PSP

[781] Bernadette Bensaude-Vincent. Popular science and politics in interwar France. Science in Context, 26(3):459–471, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Schirrmacher:2013:PSC

[782] Arne Schirrmacher. Popular science as cultural Dispositif: On the German way of science communication in the Twentieth Century. Science in Context, 26(3):473–508, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Andrews:2013:ESP

[783] James T. Andrews. An evolving scientific public sphere: State science enlightenment, communicative discourse, and public culture from Imperial Russia to Khrushchev’s Soviet times. Science in Context, 26(3): 509–526, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Nieto-Galan:2013:PNM

[784] Agustí Nieto-Galan. From papers to newspapers: Miguel Masriera (1901– 1981) and the role of science popularization under the Franco Regime. Science in Context, 26(3):527–549, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Anonymous:2013:SVIe

[785] Anonymous. SIC volume 26 issue 3 cover and front matter. Science in Context, 26(3):f1–f3, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Anonymous:2013:SVIf

[786] Anonymous. SIC volume 26 issue 3 cover and back matter. Science in Context, 26(3):b1–b10, September 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Communicating Science: National Approaches in Twentieth-Century Europe.

Elazar:2013:AML

[787] Michael Elazar, Moritz Epple, Miriam Greenfield, Orna Harari, and Jürgen Renn. Alexandre Métraux leaves editorship of Science in Context. Science in Context, 26(4):551, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Harman:2013:IRD

[788] Oren Harman. Introduction: ‘right as diverse pathes leden’. Science in Context, 26(4):553–555, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Jablonka:2013:DDB

[789] Eva Jablonka and Marion J. Lamb. Disturbing dogmas: Biologists and the history of biology. Science in Context, 26(4):557–571, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Fisch:2013:HWW

[790] Menachem Fisch. How and why I write history of science. Science in Context, 26(4):573–585, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Lloyd:2013:HAS

[791] Geoffrey Lloyd. The history of ancient science: A personal view. Science in Context, 26(4):587–593, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Oreskes:2013:WP

[792] Naomi Oreskes. Why I am a presentist. Science in Context, 26(4):595– 609, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Smith:2013:WHD

[793] Roger Smith. Why and how do I write the history of science? Science in Context, 26(4):611–625, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Kevles:2013:WHR

[794] Daniel J. Kevles. Why and how: Reflections in an autobiographical key. Science in Context, 26(4):627–638, December 2013. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Rheinberger:2013:MRH

[795] Hans-Jörg Rheinberger. My road to history of science. Science in Context, 26(4):639–648, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Metraux:2013:HSU

[796] Alexandre Métraux. The history of science as unending steeplechase: A dialogue. Science in Context, 26(4):649–664, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Karaca:2013:SWSb

[797] Koray Karaca. The strong and weak senses of theory-ladenness of experimentation: Theory-driven versus exploratory experiments in the history of high-energy particle physics — erratum. Science in Context, 26(4): 665–666, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See [762]. Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Anonymous:2013:SVIg

[798] Anonymous. SIC volume 26 issue 4 cover and front matter. Science in Context, 26(4):f1–f3, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Anonymous:2013:SVIh

[799] Anonymous. SIC volume 26 issue 4 cover and back matter. Science in Context, 26(4):b1–b4, December 2013. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Approaches, Styles, and Narratives: Reflections on Doing History of Science.

Blumenthal:2014:CDC

[800] Geoffrey Blumenthal. Copernicus’s development in context: Politics, astrology, cosmology and a prince–bishopric. Science in Context, 27(1): 1–32, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mozaffari:2014:CSH

[801] S. Mohammad Mozaffari. A case study of how natural phenomena were justified in medieval science: The situation of annular eclipses in medieval astronomy. Science in Context, 27(1):33–47, March 2014. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Simoes:2014:PAM

[802] Ana Simões and Luís Miguel Carolino. The Portuguese astronomer Melo e Simas (1870–1934): Republican ideals and popularization of science. Science in Context, 27(1):49–77, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Esposito:2014:PIR

[803] Maurizio Esposito. Problematic ‘idiosyncrasies’: Rediscovering the historical context of D’Arcy Wentworth Thompson’s science of form. Science in Context, 27(1):79–107, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Pacheco:2014:MMS

[804] José M. Pacheco. Mobility and migration of Spanish mathematicians during the years around the Spanish Civil War and World War II. Science in Context, 27(1):109–141, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Pickersgill:2014:EUA

[805] Martyn Pickersgill. The endurance of uncertainty: Antisociality and ontological anarchy in British psychiatry, 1950–2010. Science in Context, 27(1):143–175, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIa

[806] Anonymous. SIC volume 27 issue 1 cover and front matter. Science in Context, 27(1):f1–f3, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIb

[807] Anonymous. SIC volume 27 issue 1 cover and back matter. Science in Context, 27(1):b1–b8, March 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Asdal:2014:IST

[808] Kristin Asdal and Christoph Gradmann. Introduction: Science, technology, medicine — and the state: The science–state nexus in Scandinavia, 1850–1980. Science in Context, 27(2):177–186, June 2014. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Bauer:2014:AIB

[809] Susanne Bauer. From administrative infrastructure to biomedical resource: Danish population registries, the “Scandinavian Laboratory,” and the “Epidemiologist’s Dream”. Science in Context, 27(2):187–213, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Lie:2014:PSP

[810] Anne Kveim Lie. Producing standards, producing the Nordic region: Antibiotic susceptibility testing, from 1950–1970. Science in Context, 27 (2):215–248, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Widmalm:2014:USI

[811] Sven Widmalm. United in separation: The inventions of gel filtration and the moral economy of research in Swedish biochemistry, ca. 1950– 1970. Science in Context, 27(2):249–274, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Sorlin:2014:CSS

[812] Sverker Sörlin. Circumpolar science: Scandinavian approaches to the Arctic and the North Atlantic, ca. 1920 to 1960. Science in Context, 27 (2):275–305, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Asdal:2014:VMV

[813] Kristin Asdal. Versions of milk and versions of care: The emergence of mother’s milk as an interested object and medicine as a form of dispassionate care. Science in Context, 27(2):307–331, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Druglitro:2014:BTB

[814] Tone Druglitrø and Robert G. W. Kirk. Building transnational bodies: Norway and the international development of laboratory animal science, ca. 1956–1980. Science in Context, 27(2):333–357, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Kjaergaard:2014:HGP

[815] Peter C. Kjærgaard. Inventing Homo gardarensis: Prestige, pressure, and human evolution in interwar Scandinavia. Science in Context, 27 (2):359–383, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Anonymous:2014:SVIc

[816] Anonymous. SIC volume 27 issue 2 cover and front matter. Science in Context, 27(2):f1–f3, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-1980.

Anonymous:2014:SVId

[817] Anonymous. SIC volume 27 issue 2 cover and back matter. Science in Context, 27(2):b1–b8, June 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Science, Technology, Medicine - and the State: The Science-State Nexus in Scandinavia, 1850-

Pollet:2014:IQD

[818] Charlotte-V. Pollet. The influence of Qing Dynasty editorial work on the modern interpretation of mathematical sources: The case of Li Rui’s edition of Li Ye’s mathematical treatises. Science in Context, 27(3): 385–422, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bergwik:2014:FPS

[819] Staffan Bergwik. A fractured position in a stable partnership: Ebba Hult, Gerard De Geer, and Early Twentieth Century Swedish geology. Science in Context, 27(3):423–451, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Duppe:2014:SNE

[820] Till Düppe and E. Roy Weintraub. Siting the new economic science: The Cowles Commission’s activity analysis conference of June 1949. Science in Context, 27(3):453–483, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ankeny:2014:MOM

[821] Rachel A. Ankeny, Sabina Leonelli, Nicole C. Nelson, and Edmund Ramsden. Making organisms model human behavior: Situated models in North-American alcohol research, since 1950. Science in Context, 27(3): 485–509, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Abeysinghe:2014:URW

[822] Sudeepa Abeysinghe. An uncertain risk: The World Health Organization’s account of H1N1. Science in Context, 27(3):511–529, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vanRijswoud:2014:SEC

[823] Erwin van Rijswoud. Shifting expert configurations and the public credibility of science: Boundary work and identity work of hydraulic engineers (1980–2009). Science in Context, 27(3):531–558, September 2014. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:STM

[824] Anonymous. Science, technology, medicine — and the state: The science– state nexus in Scandinavia, 1850–1980 — corrigendum. Science in Context, 27(3):559, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIe

[825] Anonymous. SIC volume 27 issue 3 cover and front matter. Science in Context, 27(3):f1–f3, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIf

[826] Anonymous. SIC volume 27 issue 3 cover and back matter. Science in Context, 27(3):b1–b7, September 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Svetlova:2014:MWM

[827] Ekaterina Svetlova and Vanessa Dirksen. Models at work-models in decision making. Science in Context, 27(4):561–577, December 2014. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Landstrom:2014:VEM

[828] Catharina Landström and Sarah J. Whatmore. Virtually expert: Modes of environmental computer simulation modeling. Science in Context, 27(4):579–603, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wansleben:2014:CFV

[829] Leon Wansleben. Consistent forecasting vs. anchoring of market stories: Two cultures of modeling and model use in a bank. Science in Context, 27(4):605–630, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zeiss:2014:DDM

[830] Ragna Zeiss and Stans van Egmond. Dissolving decision making? Models and their roles in decision-making processes and policy at large. Science in Context, 27(4):631–657, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Blank:2014:NTF

[831] Andreas Blank. Nicolaus Taurellus on forms and elements. Science in Context, 27(4):659–682, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Kistner:2014:DDR

[832] Kelly Kistner. A dictionary without definitions: Romanticist science in the production and presentation of the Grimm Brothers’ German dictionary, 1838–1863. Science in Context, 27(4):683–707, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

DeKock:2014:HHE

[833] Liesbet De Kock. Hermann von Helmholtz’s empirico–transcendentalism reconsidered: Construction and constitution in Helmholtz’s psychology of the object. Science in Context, 27(4):709–744, December 2014. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIg

[834] Anonymous. SIC volume 27 issue 4 cover and front matter. Science in Context, 27(4):f1–f3, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2014:SVIh

[835] Anonymous. SIC volume 27 issue 4 cover and back matter. Science in Context, 27(4):b1–b7, December 2014. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Casper:2015:MEM

[836] Stephen T. Casper. Of means and ends: Mind and brain science in the Twentieth Century. Science in Context, 28(1):1–7, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Miller:2015:DPD

[837] Jason Miller. Dredging and projecting the depths of personality: The thematic apperception test and the narratives of the unconscious. Science in Context, 28(1):9–30, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Garson:2015:BIB

[838] Justin Garson. The birth of information in the brain: Edgar Adrian and the vacuum tube. Science in Context, 28(1):31–52, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Blatter:2015:SPL

[839] Jeremy Blatter. Screening the psychological laboratory: Hugo Münsterberg, psychotechnics, and the cinema, 1892–1916. Science in Context, 28(1):53–76, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Schilling:2015:PMS

[840] Rebecca Schilling and Stephen T. Casper. Of psychometric means: Starke R. Hathaway and the popularization of the Minnesota Multiphasic Personality Inventory. Science in Context, 28(1):77–98, March 2015. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Casey:2015:SEV

[841] Brian P. Casey. The surgical elimination of violence? Conflicting attitudes towards technology and science during the psychosurgery controversy of the 1970s. Science in Context, 28(1):99–129, March 2015. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Scull:2015:CPS

[842] Andrew Scull. Contending professions: Sciences of the brain and mind in the United States, 1850–2013. Science in Context, 28(1):131–161, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Buchanan:2015:ERP

[843] Roderick D. Buchanan. Epilogue: The redux of postmodernity. Science in Context, 28(1):163–170, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Anonymous:2015:SVIa

[844] Anonymous. SIC volume 28 issue 1 cover and front matter. Science in Context, 28(1):f1–f3, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Anonymous:2015:SVIb

[845] Anonymous. SIC volume 28 issue 1 cover and back matter. Science in Context, 28(1):b1–b5, March 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Of Means and Ends: Mind and Brain Science in the Twentieth Century.

Goldin:2015:PTO

[846] Owen Goldin. The Pythagorean table of opposites, symbolic classification, and Aristotle. Science in Context, 28(2):171–193, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Vanderstraeten:2015:DDS

[847] Raf Vanderstraeten and Frederic Vandermoere. Disciplined by the discipline: A social-epistemic fingerprint of the history of science. Science in Context, 28(2):195–214, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Zysiak:2015:MSB

[848] Agata Zysiak. Modernizing science: Between a liberal, social, and socialistic university — the case of Poland and the University of Łód’z (1945–1953). Science in Context, 28(2):215–236, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Arpin:2015:RTJ

[849] Isabelle Arpin and Céline Granjou. The right time for the job? Insights into practices of time in contemporary field sciences. Science in Context, 28(2):237–258, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lachmund:2015:SBO

[850] Jens Lachmund. Strange birds: Ornithology and the advent of the collared dove in post-World War II Germany. Science in Context, 28(2): 259–284, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Fried:2015:OTP

[851] Michael N. Fried and Hans Niels Jahnke. Otto Toeplitz’s 1927 paper on the genetic method in the teaching of mathematics. Science in Context, 28(2):285–295, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Toeplitz:2015:PUC

[852] Otto Toeplitz. The problem of university courses on infinitesimal calculus and their demarcation from infinitesimal calculus in high schools. Science in Context, 28(2):297–310, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). See erratum [899].

Anonymous:2015:SVIc

[853] Anonymous. SIC volume 28 issue 2 cover and front matter. Science in Context, 28(2):f1–f3, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2015:SVId

[854] Anonymous. SIC volume 28 issue 2 cover and back matter. Science in Context, 28(2):b1–b3, June 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Grote:2015:ISH

[855] Mathias Grote and Max Stadler. Introduction: Surface histories. Science in Context, 28(3):311–315, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ramalingam:2015:DPR

[856] Chitra Ramalingam. Dust plate, retina, photograph: Imaging on experimental surfaces in early nineteenth-century physics. Science in Context, 28(3):317–355, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schmidgen:2015:LMH

[857] Henning Schmidgen. Leviathan and the myograph: Hermann Helmholtz’s “second note” on the propagation speed of nervous stimulations. Science in Context, 28(3):357–396, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Espahangizi:2015:TOS

[858] Kijan Espahangizi. From topos to oikos: The standardization of glass containers as epistemic boundaries in modern laboratory research (1850– 1900). Science in Context, 28(3):397–425, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

vonSchwerin:2015:SVB

[859] Alexander von Schwerin. Shaping vulnerable bodies at the thin boundary between environment and organism: Skin, DNA repair, and a genealogy of DNA care strategies. Science in Context, 28(3):427–464, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Berrey:2015:RCH

[860] Marquis Berrey. Reading communities and hippocratism in Hellenistic medicine. Science in Context, 28(3):465–487, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Bjorkman:2015:EGC

[861] Maria Björkman. The emergence of genetic counseling in Sweden: Examples from eugenics and medical genetics. Science in Context, 28(3): 489–513, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2015:SVIe

[862] Anonymous. SIC volume 28 issue 3 cover and front matter. Science in Context, 28(3):f1–f3, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2015:SVIf

[863] Anonymous. SIC volume 28 issue 3 cover and back matter. Science in Context, 28(3):b1–b11, September 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Sangiacomo:2015:ODD

[864] Andrea Sangiacomo. The ontology of determination: From Descartes to Spinoza. Science in Context, 28(4):515–543, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lorenat:2015:PPP

[865] Jemma Lorenat. Polemics in public: Poncelet, Gergonne, Plücker, and the duality controversy. Science in Context, 28(4):545–585, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Ekerholm:2015:KHS

[866] Helena Ekerholm. Keeping a house for science: Sofia Kristensson as matriarch and gatekeeper at Kristineberg Zoological Station as a scientific household, 1877–1889. Science in Context, 28(4):587–611, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Mueller:2015:BDR

[867] Thomas Michael Mueller. The Boussinesq debate: Reversibility, instability, and free will. Science in Context, 28(4):613–635, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Capristo:2015:VFF

[868] Annalisa Capristo. Volterra, Fascism, and France. Science in Context, 28(4):637–674, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Grune-Yanoff:2015:MTD

[869] Till Grüne-Yanoff. Models of temporal discounting 1937–2000: An interdisciplinary exchange between economics and psychology. Science in Context, 28(4):675–713, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2015:SVIg

[870] Anonymous. SIC volume 28 issue 4 cover and front matter. Science in Context, 28(4):f1–f3, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2015:SVIh

[871] Anonymous. SIC volume 28 issue 4 cover and back matter. Science in Context, 28(4):b1–b7, December 2015. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Metraux:2016:SIH

[872] Alexandre Métraux. On some issues of human-animal studies: An introduction. Science in Context, 29(1):1–10, March 2016. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Schmitt:2016:SAR

[873] Stéphane Schmitt. Studies on animals and the rise of comparative anatomy at and around the Parisian Royal Academy of Sciences in the Eighteenth Century. Science in Context, 29(1):11–54, March 2016. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Marinelli:2016:RAT

[874] Lydia Marinelli and Andreas Mayer. The receding animal: Theorizing anxiety and attachment in psychoanalysis from Freud to Imre Hermann.

Science in Context, 29(1):55–76, March 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Winter:2016:CCE

[875] Alison Winter. Cats on the couch: The experimental production of animal neurosis. Science in Context, 29(1):77–105, March 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Benson:2016:NES

[876] Etienne S. Benson. Naming the ethological subject. Science in Context, 29(1):107–128, March 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Anonymous:2016:SVIa

[877] Anonymous. SIC volume 29 issue 1 cover and front matter. Science in Context, 29(1):f1–f3, March 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Anonymous:2016:SVIb

[878] Anonymous. SIC volume 29 issue 1 cover and back matter. Science in Context, 29(1):b1–b7, March 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). Special issue on Animal-Man Relations.

Sialaros:2016:SDG

[879] Michalis Sialaros and Jean Christianidis. Situating the debate on “geometrical algebra” within the framework of premodern algebra. Science in Context, 29(2):129–150, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Blumenthal:2016:CPS

[880] Geoffrey Blumenthal. Copernicus’s publication strategy in the contexts of Imperial and papal censorship and of Warmian diplomatic precedents. Science in Context, 29(2):151–178, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Lee:2016:MTC

[881] Jung Lee. Mutual transformation of colonial and Imperial botanizing? The intimate yet remote collaboration in colonial Korea. Science in Context, 29(2):179–211, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Eriksen:2016:AIE

[882] Anne Eriksen. Advocating inoculation in the eighteenth century: Exemplarity and quantification. Science in Context, 29(2):213–239, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2016:SVIc

[883] Anonymous. SIC volume 29 issue 2 cover and front matter. Science in Context, 29(2):f1–f3, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Anonymous:2016:SVId

[884] Anonymous. SIC volume 29 issue 2 cover and back matter. Science in Context, 29(2):b1–b7, June 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Wagner:2016:WFM

[885] Roi Wagner. Wronski’s foundations of mathematics. Science in Context, 29(3):241–271, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/FB869BE806F705CD4B0C5688BD50F78C.

Le:2016:RNC

[886] François Lê. Reflections on the notion of culture in the history of mathematics: The example of “geometrical equations”. Science in Context, 29(3):273–304, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/product/8AE469D50F9A6C97300E0E9589597AEA.

Bergman:2016:KTP

[887] James Bergman. Knowing their place: The Blue Hill Observatory and the value of local knowledge in an era of synoptic weather forecasting, 1884-1894. Science in Context, 29(3):305–346, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
1A6A37B337A8D6E18B2D2F80C58E97DC.

Garcia:2016:SKP

[888] Mónica García and Stefan Pohl-Valero. Styles of knowledge production in Colombia, 1850–1920. Science in Context, 29(3):347–377, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
AB19F49D658AF246E3262B4DE68AE4A0.

Anonymous:2016:SVIe

[889] Anonymous. SIC volume 29 issue 3 cover and front matter. Science in Context, 29(3):f1–f3, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/BB02354BEE230417DDAFD43C179038E2.

Anonymous:2016:SVIf

[890] Anonymous. SIC volume 29 issue 3 cover and back matter. Science in Context, 29(3):b1–b5, September 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/8EAB6666035B197C7AC28735A5ABC241.

Metraux:2016:EMI

[891] Alexandre Métraux. Ernst Mach invents a new sphygmograph. Science in Context, 29(4):379–407, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/636C7FD19EEFF122E532D61431B7AC8F.

Hoffmann:2016:PVP

[892] Christoph Hoffmann. Paris vs. Prague: A “suspicion of fraud”: Ernst Mach argues over photographs and epistemological prerequisites. Science in Context, 29(4):409–427, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/8E55AA89173208CA715CAE47E5086ABC.

Hoffmann:2016:WIE

[893] Christoph Hoffmann and Alexandre Métraux. Working with instruments: Ernst Mach as material epistemologist, a short introduction. Science in Context, 29(4):429–433, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/FC12CF264600875CFC70A1A54F28A542.

Mach:2016:TS

[894] Ernst Mach. On the theory of the sphygmograph. Science in Context, 29(4):435–440, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/FB992884080A04015423E34E0F9FD86B.

Mach:2016:RSA

[895] Ernst Mach. Remarks on scientific applications of photography. Science in Context, 29(4):441–442, December 2016. CODEN SCCOEW.

ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/62014D196C48F60527D3CD4778B6D5F6.

Mach:2016:PTD

[896] Ludwig Mach. On the principle of temporal diminution in serial photography. Science in Context, 29(4):443–450, December 2016. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/product/CBF03218FD6E442001F6ECC6C2F85090.

Perez:2016:MWE

[897] Nuria Valverde Pérez. Meanings of waves: Electroencephalography and society in Mexico City, 1940–1950. Science in Context, 29(4):451–472, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
2F37CA81187A9B0FED06167AF36B7CDB.

Radtka:2016:HPA

[898] Catherine Radtka. How place and audience matter: Perspectives on mathematics plural identities from late 1950s French and English middle school textbooks. Science in Context, 29(4):473–521, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
E1C84898A99850E9196DB9ACE0BC922F.

Toeplitz:2016:PUC

[899] Otto Toeplitz. The problem of university courses on infinitesimal calculus and their demarcation from infinitesimal calculus in high schools — Erratum. Science in Context, 29(4):523, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
087FCAEE1F2082AFE083CE99ECD005A4. See [852].

Anonymous:2016:SVIg

[900] Anonymous. SIC volume 29 issue 4 cover and front matter. Science in Context, 29(4):f1–f3, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/A4BDB09EC7EAB4D7DAC2AB37243FF1A4.

Anonymous:2016:SVIh

[901] Anonymous. SIC volume 29 issue 4 cover and back matter. Science in Context, 29(4):b1–b5, December 2016. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/81A23885D5C1D8ADCD01C858A6132940.

Mozaffari:2017:HBP

[902] S. Mohammad Mozaffari. Holding or breaking with Ptolemy’s generalization: Considerations about the motion of the planetary apsidal lines in medieval Islamic astronomy. Science in Context, 30(1):1– 32, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
CB8327AAC2AE90885D75B06AC5D89053.

Liu:2017:RPM

[903] Chien-Ling Liu. Relocating Pastorian medicine: Accommodation and acclimatization of Pastorian practices against smallpox at the Pasteur Institute of Chengdu, China, 1908–1927. Science in Context, 30(1):33– 59, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
87260B9400788C45B7AD9AD1750C2FBB.

Krasnodebski:2017:CSF

[904] Marcin Krasnodebski. Can science feed on a crisis? Expectations, the Pine Institute, and the decline of the French resin industry. Science in Context, 30(1):61–87, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/94D28F18B87556CDCAA25B571D49CC97.

Anaya-Munoz:2017:FUH

[905] Víctor Hugo Anaya-Muñoz, Vivette García-Deister, and Edna Suárez-Díaz. Flattening and unpacking human genetic variation in Mexico, postwar to present. Science in Context, 30(1):89–112, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https:

//www.cambridge.org/core/product/24809F13F204E4C9025BDA6AD3F44167.

Anonymous:2017:SVIa

[906] Anonymous. SIC volume 30 issue 1 cover and front matter. Science in Context, 30(1):f1–f3, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/A670CE607D9FD6E15F137A3677318F76.

Anonymous:2017:SVIb

[907] Anonymous. SIC volume 30 issue 1 cover and back matter. Science in Context, 30(1):b1–b4, March 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/D208831A78BBB68F2514C369F4F5936A.

Ratzon:2017:FJA

[908] Eshbal Ratzon. The first Jewish astronomers: Lunar theory and reconstruction of a Dead Sea scroll. Science in Context, 30(2):113– 139, June 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
00F88339D2C03CC48D3D9562AF3F1340.

Oaks:2017:ICR

[909] Jeffrey A. Oaks. Irrational “coefficients” in Renaissance algebra. Science in Context, 30(2):141–172, June 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/38391278E789631131A9F9BFC4DC8B0B.

Schaffer:2017:OMP

[910] Simon Schaffer. Oriental metrology and the politics of antiquity in nineteenth-century survey sciences. Science in Context, 30(2):173– 212, June 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
9E92049F31B702B2C58A5B809752B3F9.

Anonymous:2017:SVIc

[911] Anonymous. SIC volume 30 issue 2 cover and front matter. Science in Context, 30(2):f1–f3, June 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/4CF63B1098A15DD303738F34F385700C.

Anonymous:2017:SVId

[912] Anonymous. SIC volume 30 issue 2 cover and back matter. Science in Context, 30(2):b1–b3, June 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/35D5A2C436E72FA43AE961D5B5D63DD7.

Sela:2017:IPO

[913] Ori Sela. Introduction: Paving the old-new way from Qing to China. Science in Context, 30(3):213–217, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/E5BE5D76C5D3CCDB4D0A79F49FAD93E3.

Hanson:2017:VGD

[914] Marta Hanson. Visualizing the geography of the diseases of China: Western disease maps from analytical tools to tools of empire, sovereignty, and public health propaganda, 1878–1929. Science in Context, 30(3):219–280, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
EC1BFFB27B2EA0B5524DC171FB6C55F9.

Brokaw:2017:DON

[915] Cynthia Brokaw. The dance of “old” and “new” in Chinese print culture, 1860s–1955. Science in Context, 30(3):281–324, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/product/
A018D803C1D6151E6449F58CE9985A59.

Sela:2017:FHW

[916] Ori Sela. “To feel at home in the wonderful world of modern science”: New Chinese historiography and Qing intellectual history. Science in Context, 30(3):325–358, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/8ED7CC364CB9ABFFDDE66BF320AF4DE6.

Judge:2017:SCC

[917] Joan Judge. Science for the Chinese common reader? Myriad treasures and new knowledge at the turn of the Twentieth Century. Science in Context, 30(3):359–383, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/product/C776740AEF6DC6BDEE5F48AE584C7361.

Anonymous:2017:SVIe

[918] Anonymous. SIC volume 30 issue 3 cover and front matter. Science in Context, 30(3):f1–f3, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/F362A17E5FDB3C07F5928BFED5EDD32C.

Anonymous:2017:SVIf

[919] Anonymous. SIC volume 30 issue 3 cover and back matter. Science in Context, 30(3):b1–b3, September 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/product/22459093974EA45E8AAE14705B01D7E6.

Guzzardi:2017:RBF

[920] Luca Guzzardi. Ruggiero Boscovich and “the forces existing in nature”. Science in Context, 30(4):385–422, December 2017. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
ruggiero-boscovich-and-the-forces-existing-in-nature/5E54A09C66AB5E7081DF931D4A40A173.

Quick:2017:DPP

[921] Tom Quick. Disciplining physiological psychology: Cinematographs as epistemic devices in the work of Henri Bergson and Charles Scott Sherrington. Science in Context, 30(4):423–474, December 2017. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
disciplining-physiological-psychology-cinematographs-as-epistemic-
devices-in-the-work-of-henri-bergson-and-charles-scott-sherrington/
EF967716D83DBFCAC33C66F3F8E23180.

Martin:2017:PAA

[922] Joseph D. Martin. Prestige asymmetry in American physics: Aspirations, applications, and the purloined letter effect. Science in Context, 30(4):475–506, December 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
prestige-asymmetry-in-american-physics-aspirations-applications-
and-the-purloined-letter-effect/7421B49270A1E216DCE37E08E95F4B22.

Anonymous:2017:SVIg

[923] Anonymous. SIC volume 30 issue 4 cover and front matter. Science in Context, 30(4):f1–f3, December 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-30-
issue-4-cover-and-front-matter/A4D65A4530BD5AB04BDC72CDDC700122.

Anonymous:2017:SVIh

[924] Anonymous. SIC volume 30 issue 4 cover and back matter. Science in Context, 30(4):b1–b4, December 2017. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-30-
issue-4-cover-and-back-matter/10CE35B47B0F530F2E6B3FCEA9073AB2.

Vidal:2018:IPS

[925] Fernando Vidal. Introduction: From “the popularization of science through film” to “the public understanding of science”. Science in Context, 31(1):1–14, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/introduction-from-
the-popularization-of-science-through-film-to-the-public-understanding-
of-science/0D4A4414C0BE83E8ABB6B4BB923EA1E5.

Louson:2018:TSS

[926] Eleanor Louson. Taking spectacle seriously: Wildlife film and the legacy of natural history display. Science in Context, 31(1):15–38, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/taking-spectacle-seriously-wildlife-film-
and-the-legacy-of-natural-history-display/B52FF1E16D671B635E3B7CED0105CA3C.

Mellor:2018:CEA

[927] Felicity Mellor. Configuring epistemic authority: The significance of film style in documentaries about science. Science in Context, 31(1): 39–59, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/configuring-epistemic-authority-
the-significance-of-film-style-in-documentaries-about-science/
7A858C685B69C6CA36393845F2073424.

Tabernero:2018:CNM

[928] Carlos Tabernero. The changing nature of modernization discourses in documentary films. Science in Context, 31(1):61–83, March 2018. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/changing-nature-of-modernization-discourses-in-documentary-
films/F5F47CFDFDBCB9413C80CE7B9C60CF47.

Kirby:2018:HPP

[929] David A. Kirby. Harnessing the persuasive power of narrative: Science, storytelling, and movie censorship, 1930–1968. Science in Context, 31(1): 85–106, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/harnessing-the-persuasive-power-
of-narrative-science-storytelling-and-movie-censorship-19301968/
37FDFA939151C5D4255ABC4FE5BCE047.

Ostherr:2018:SAE

[930] Kirsten Ostherr. The shifting aesthetics of expertise in the sharing economy of scientific medicine. Science in Context, 31(1):107–127, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/shifting-aesthetics-of-expertise-in-the-
sharing-economy-of-scientific-medicine/80C584CA30E1C930E1EABA1730646E3C.

Vidal:2018:AAF

[931] Fernando Vidal. Accuracy, authenticity, fidelity: Aesthetic realism, the “deficit model,” and the public understanding of science. Science in Context, 31(1):129–153, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
accuracy-authenticity-fidelity-aesthetic-realism-the-deficit-
model-and-the-public-understanding-of-science/0DC36D7B508937C740A61FBDFE614D5E.

Anonymous:2018:SVIa

[932] Anonymous. SIC volume 31 issue 1 cover and front matter. Science in Context, 31(1):f1–f3, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-1-cover-and-front-matter/27326D93FDADE6F7096EA1641440E71B.

Anonymous:2018:SVIb

[933] Anonymous. SIC volume 31 issue 1 cover and back matter. Science in Context, 31(1):b1–b4, March 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-1-cover-and-back-matter/40DE1E405F8B45470BD2A47ED3AD471B.

Frumer:2018:JRC

[934] Yulia Frumer. Japanese reverse compasses: Grounding cognition in history and society. Science in Context, 31(2):155–187, June 2018. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/japanese-reverse-compasses-grounding-cognition-in-history-
and-society/927F435A7E67BB339B0653896C42EFF4.

Goddu:2018:BCH

[935] André Goddu. Birkenmajer’s Copernicus: Historical context, original insights, and contributions to current debates. Science in Context, 31(2): 189–222, June 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/birkenmajers-copernicus-historical-
context-original-insights-and-contributions-to-current-debates/
C88B8091A07958EEA3A50CB21399A942.

dePaz:2018:JMJ

[936] María de Paz. From jurisprudence to mechanics: Jacobi, Reech, and Poincaré on convention. Science in Context, 31(2):223–250, June 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/from-jurisprudence-to-mechanics-
jacobi-reech-and-poincare-on-convention/D76A99F689F7C8923D55CFC05988D1FA.

Anonymous:2018:SVIc

[937] Anonymous. SIC volume 31 issue 2 cover and front matter. Science in Context, 31(2):f1–f3, June 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-2-cover-and-front-matter/E27EFBD7E514BCA32E4A0D872D7D1208.

Anonymous:2018:SVId

[938] Anonymous. SIC volume 31 issue 2 cover and back matter. Science in Context, 31(2):b1–b4, June 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-2-cover-and-back-matter/3EFBC6FBDB4070EC92A4F73D0CE8BEA9.

Katzir:2018:IPT

[939] Shaul Katzir. Introduction: Physics, technology, and technics during the interwar period. Science in Context, 31(3):251–261, September 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/introduction-physics-technology-
and-technics-during-the-interwar-period/FDA514C8C4CC9B07BF6B254B4F0764C4.

Staley:2018:IPM

[940] Richard Staley. The interwar period as a machine age: Mechanics, the machine, mechanisms, and the market in discourse. Science in Context, 31(3):263–292, September 2018. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
interwar-period-as-a-machine-age-mechanics-the-machine-mechanisms-
and-the-market-in-discourse/0FB582D16E51AEC2598812A280DDAF55.

Ito:2018:ETE

[941] Kenji Ito. “Electron theory” and the emergence of atomic physics in Japan. Science in Context, 31(3):293–320, September 2018. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
electron-theory-and-the-emergence-of-atomic-physics-in-japan/
9339A13C96EEC95A848D72F2B23084E5.

Katzir:2018:SIP

[942] Shaul Katzir. The shaping of interwar physics by technology: The case of piezoelectricity. Science in Context, 31(3):321–350, September 2018. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/shaping-of-interwar-physics-by-technology-the-case-
of-piezoelectricity/F911CA4E5025917D6BDCB31D486BE77A.

Navarro:2018:HUF

[943] Jaume Navarro. How to use a fundamental discovery in physics: The early days of electron diffraction. Science in Context, 31(3):351–379, September 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/how-to-use-a-fundamental-discovery-
in-physics-the-early-days-of-electron-diffraction/C7A7CEBF9387943E335DC62E77299F1E.

Eckert:2018:TRB

[944] Michael Eckert. Turbulence research in the 1920s and 1930s between mathematics, physics, and engineering. Science in Context, 31(3):381– 404, September 2018. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/turbulence-research-in-
the-1920s-and-1930s-between-mathematics-physics-and-engineering/
3C50916CA865D0F67BB111FBC6F95D07.

Anonymous:2018:SVIe

[945] Anonymous. SIC volume 31 issue 3 cover and front matter. Science in Context, 31(3):f1–f3, September 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-3-cover-and-front-matter/18FD7A39410C764F5FCADAA3C45DE53C.

Anonymous:2018:SVIf

[946] Anonymous. SIC volume 31 issue 3 cover and back matter. Science in Context, 31(3):b1–b4, September 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-3-cover-and-back-matter/D7A8B826D0126AC1CEE442C52852322F.

Katayama:2018:API

[947] Errol G. Katayama. Aristotle’s Physics II 1 and cultivated plants. Science in Context, 31(4):405–419, December 2018. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
aristotles-physics-ii-1-and-cultivated-plants/E2A13E9BED99761DDF7F6E00340EEE07.

Sangiacomo:2018:SCA

[948] Andrea Sangiacomo. Samuel Clarke on agent causation, voluntarism, and occasionalism. Science in Context, 31(4):421–456, December 2018. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/samuel-clarke-on-agent-
causation-voluntarism-and-occasionalism/00AE01A79015E487DDB667474574F929.

Vanderstraeten:2018:WFS

[949] Raf Vanderstraeten. Who had faith in sociology? scholarly and ideological divergences in Belgium around 1900. Science in Context, 31(4):457–475, December 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/who-had-faith-
in-sociology-scholarly-and-ideological-divergences-in-belgium-
around-1900/003D1F6F2CF8497C2D0062109F83B075.

Grondal:2018:EAR

[950] Hedvig Gröndal. The emergence of antimicrobial resistance as a public matter of concern: A Swedish history of a “transformative event”. Science in Context, 31(4):477–500, December 2018. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
emergence-of-antimicrobial-resistance-as-a-public-matter-of-
concern-a-swedish-history-of-a-transformative-event/B9D962C8DE915EB0A38BB4162F8604F2.

Brennen:2018:SNP

[951] J. Scott Brennen. The SENSE of nuclear physics: New frontiers, media, and collaborations. Science in Context, 31(4):501–520, December 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/sense-of-nuclear-physics-new-frontiers-media-
and-collaborations/8D39E7574D3582AEE062A426AE848BAD.

Anonymous:2018:SVIg

[952] Anonymous. SIC volume 31 issue 4 cover and front matter. Science in Context, 31(4):f1–f3, December 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-4-cover-and-front-matter/668E0B6099240E4D81BA7CD56447B2CC.

Anonymous:2018:SVIh

[953] Anonymous. SIC volume 31 issue 4 cover and back matter. Science in Context, 31(4):b1–b3, December 2018. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-31-
issue-4-cover-and-back-matter/A026700D9C42A17208ACCF0A5BD57CBB.

Davidovitch:2019:SMP

[954] Nadav Davidovitch and Rakefet Zalashik. Scientific medicine and the politics of public health: Minorities in interwar Eastern Europe. Science in Context, 32(1):1–4, March 2019. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
scientific-medicine-and-the-politics-of-public-health-minorities-
in-interwar-eastern-europe/E06991F9BF59C8771DEB9BFD48A2AED9.

Mogilner:2019:ARC

[955] Marina Mogilner. ARA relief campaign in the Volga region, Jewish anthropometric statistics, and the scientific promise of integration. Science in Context, 32(1):5–24, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.

cambridge.org/core/journals/science-in-context/article/ara-
relief-campaign-in-the-volga-region-jewish-anthropometric-statistics-
and-the-scientific-promise-of-integration/EAF851CAF6AFE200A71FE09B94B6C3A0.

Tuszewicki:2019:GTS

[956] Marek Tuszewicki. Giving Tshuve to the sick: correspondence columns of the Yiddish medical press in Poland. Science in Context, 32(1):25–41, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/giving-tshuve-
to-the-sick-correspondence-columns-of-the-yiddish-medical-press-
in-poland/E72A83CEA9BCF218DBD4B6097820F185.

Kreuder-Sonnen:2019:ESB

[957] Katharina Kreuder-Sonnen. Epidemiological state-building in interwar Poland: discourses and paper technologies. Science in Context, 32(1): 43–65, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/epidemiological-statebuilding-in-
interwar-poland-discourses-and-paper-technologies/6C5D68882D1843DEF01E859979D8E7EF.

Rudling:2019:ERA

[958] Per Anders Rudling. Eugenics and racial anthropology in the Ukrainian radical nationalist tradition. Science in Context, 32(1):67–91, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/eugenics-and-racial-anthropology-in-the-
ukrainian-radical-nationalist-tradition/16E526E005BCAD7A8B34D15EE3D243DA.

Zalashik:2019:CPJ

[959] Rakefet Zalashik and Nadav Davidovitch. The course of professionalization: Jewish nursing in Poland in the interwar period. Science in Context, 32(1):93–109, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
course-of-professionalization-jewish-nursing-in-poland-in-the-
interwar-period/01C649A838CCFBC42C34ABB102D8E4DC.

Anonymous:2019:SVIa

[960] Anonymous. SIC volume 32 issue 1 cover and front matter. Science in Context, 32(1):f1–f3, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-1-cover-and-front-matter/B0FA676282ECD621DEFD48F0D78842DA.

Anonymous:2019:SVIb

[961] Anonymous. SIC volume 32 issue 1 cover and back matter. Science in Context, 32(1):b1–b4, March 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-1-cover-and-back-matter/161561CD66CABC4D87ECE22C8389612C.

Dietz:2019:YMH

[962] Feike Dietz and Sven Dupré. Youthful minds and hands: Learning practical knowledge in early modern Europe. Science in Context, 32(2): 113–118, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/youthful-minds-and-hands-learning-
practical-knowledge-in-early-modern-europe/25B1594445A76912785023054C66A0EC.

Oosterhoff:2019:ARU

[963] Richard J. Oosterhoff. Apprenticeship in the Renaissance University: Student authorship and craft knowledge. Science in Context, 32(2):119–136, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/apprenticeship-in-
the-renaissance-university-student-authorship-and-craft-knowledge/
D05402020DA1214C278492E760D7D411.

Barry:2019:EPSa

[964] Jonathan Barry. Educating physicians in seventeenth-century England. Science in Context, 32(2):137–154, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/educating-physicians-in-seventeenthcentury-england/
A689F030E705C748E9D2382689E18C5F. See addendum [976].

Wallis:2019:BAS

[965] Patrick Wallis. Between apprenticeship and skill: Acquiring knowledge outside the Academy in Early Modern England. Science in Context, 32 (2):155–170, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/between-apprenticeship-
and-skill-acquiring-knowledge-outside-the-academy-in-early-
modern-england/FE0E11A30FECA8CD92CB2AD362066310.

Wandel:2019:MP

[966] Lee Palmer Wandel. Maps for a prince. Science in Context, 32(2):171–192, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/maps-for-a-prince/
34B540D306A3FB89CA0D0E3EA7933312.

Dietz:2019:MEE

[967] Feike Dietz. Mediated education in early modern travel stories: How travel stories contribute to children’s empirical learning. Science in Context, 32(2):193–212, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
mediated-education-in-early-modern-travel-stories-how-travel-
stories-contribute-to-childrens-empirical-learning/78498CF8CB961E90D5933D305AEC2461.

Stronks:2019:CYK

[968] Els Stronks. Curiosity, youth, and knowledge in the visual and textual culture of the Dutch Republic. Science in Context, 32(2):213–236, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/curiosity-youth-and-knowledge-in-the-visual-
and-textual-culture-of-the-dutch-republic/4566309B7F14EDD5BD931A62DDE36872.

Anonymous:2019:SVIc

[969] Anonymous. SIC volume 32 issue 2 cover and front matter. Science in Context, 32(2):f1–f3, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-2-cover-and-front-matter/CE2BE930DC4414175AFC072AB9ECD394.

Anonymous:2019:SVId

[970] Anonymous. SIC volume 32 issue 2 cover and back matter. Science in Context, 32(2):b1–b3, June 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-2-cover-and-back-matter/F071D9F06A4829E57954F493F6704933.

Hyun:2019:BPS

[971] Jaehwan Hyun. Blood purity and scientific independence: blood science and postcolonial struggles in Korea, 1926–1975. Science in Context, 32(3):239–260, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
blood-purity-and-scientific-independence-blood-science-and-
postcolonial-struggles-in-korea-19261975/AED85577C6404CF632D1EE3BAD8D949A.

Message:2019:DDE

[972] Reuben Message. “The disadvantages of a defective education”: identity, experiment and persuasion in the natural history of the salmon and Parr controversy, c. 1825–1850. Science in Context, 32(3):261–284, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/disadvantages-
of-a-defective-education-identity-experiment-and-persuasion-
in-the-natural-history-of-the-salmon-and-parr-controversy-c-
18251850/77A521B335B4E55087EEFF893FDE2DDE.

Norrgren:2019:HEA

[973] Hilde Norrgrén. Hans Egede (1686–1758) and the alchemical tradition in Denmark–Norway. Science in Context, 32(3):285–307, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/hans-egede-16861758-and-the-alchemical-tradition-
in-denmarknorway/7870777F0ED519EA45ECE7B910274D55.

Duschinsky:2019:AAB

[974] Robbie Duschinsky. Attachment and the archive: barriers and facilitators to the use of historical sociology as complementary developmental science. Science in Context, 32(3):309–326, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/attachment-and-the-archive-barriers-and-
facilitators-to-the-use-of-historical-sociology-as-complementary-
developmental-science/8087070DFE379149FD77B64E43E56766.

Koch:2019:UTE

[975] Ulrich Koch. The uses of trauma in experiment: Traumatic stress and the history of experimental neurosis, c. 1925–1975. Science in Context, 32(3):327–351, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/uses-of-trauma-in-
experiment-traumatic-stress-and-the-history-of-experimental-
neurosis-c-19251975/D54F4D7B9AAC43BA0B3DFAAD58599235.

Barry:2019:EPSb

[976] Jonathan Barry. Educating physicians in seventeenth-century England — addendum. Science in Context, 32(3):353, September 2019. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/educating-physicians-in-seventeenthcentury-england-
addendum/C60033BF71BBAC61DE88AC1A8E8AC449. See [964].

Anonymous:2019:SVIe

[977] Anonymous. SIC volume 32 issue 3 cover and front matter. Science in Context, 32(3):f1–f3, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-3-cover-and-front-matter/FC5E4B32847ED04DA6B447813FAA5ECB.

Anonymous:2019:SVIf

[978] Anonymous. SIC volume 32 issue 3 cover and back matter. Science in Context, 32(3):b1–b3, September 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-3-cover-and-back-matter/D6613214C4228A0D9E576E3020F1B3F5.

Kapeller:2019:ICP

[979] Jakob Kapeller and Daniel Meyer. Introduction: change and persistence in contemporary economics. Science in Context, 32(4):357–360, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/introduction-change-and-persistence-in-contemporary-
economics/CD639B41D68030A6632416D309991844.

Aistleitner:2019:CPE

[980] Matthias Aistleitner, Jakob Kapeller, and Stefan Steinerberger. Citation patterns in economics and beyond. Science in Context, 32 (4):361–380, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/citation-patterns-
in-economics-and-beyond/C0B9173D478C3B7767BC1826F555A422.

Glotzl:2019:SDC

[981] Florentin Glötzl and Ernest Aigner. Six dimensions of concentration in economics: Evidence from a large-scale data set. Science in Context, 32(4):381–410, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/six-dimensions-of-
concentration-in-economics-evidence-from-a-largescale-data-
set/896C6A7390B3C8128BCC2A9EDFF259F8.

Corsi:2019:EGC

[982] Marcella Corsi, Carlo D’Ippoliti, and Giulia Zacchia. On the evolution of the glass ceiling in Italian academia: the case of economics. Science in Context, 32(4):411–430, December 2019. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
on-the-evolution-of-the-glass-ceiling-in-italian-academia-the-
case-of-economics/FA1ECE5B84D47247AD1B5F21683A928C.

Winzler:2019:SEE

[983] Tim Winzler. The superiority of economics and the economics of externalism — a sketch. Science in Context, 32(4):431–447, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/superiority-of-economics-and-the-economics-
of-externalism-a-sketch/468E3097BD1416C53C42005510B40C30.

Jacob:2019:CAD

[984] Ulrike Jacob and Oliver A. Brust. Confronting the anomaly: directions in (German) economic research after the crisis. Science in Context, 32(4):449–471, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
confronting-the-anomaly-directions-in-german-economic-research-
after-the-crisis/90EC8251E7DC14CB8FED5E5D0E3B48F4.

Anonymous:2019:SVIh

[985] Anonymous. SIC volume 32 issue 4 cover and front matter. Science in Context, 32(4):f1–f3, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-4-cover-and-front-matter/5EA4FA2DD579C26AE083807F7E9426DA.

Anonymous:2019:SVIi

[986] Anonymous. SIC volume 32 issue 4 cover and back matter. Science in Context, 32(4):b1–b3, December 2019. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-32-
issue-4-cover-and-back-matter/2A56B369CEEF9A78B1AE5DA1D4AB820B.

Pecere:2020:RID

[987] Paolo Pecere. Reconsidering the ignorabimus: du Bois-Reymond and the hard problem of consciousness. Science in Context, 33 (1):1–18, March 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/reconsidering-the-
ignorabimus-du-boisreymond-and-the-hard-problem-of-consciousness/
8C19E49A253DD5011482AD4252FF563D.

Beiermann:2020:PPL

[988] Lea Beiermann and Elisabeth Wesseling. Physiology and philhellenism in the late nineteenth century: The self-fashioning of Emil du Bois-Reymond. Science in Context, 33(1):19–35, March 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664

(electronic).  URL https://www.cambridge.org/core/journals/
science-in-context/article/physiology-and-philhellenism-in-
the-late-nineteenth-century-the-selffashioning-of-emil-du-boisreymond/
41EDA6E1FC05565B0A2EEE01C0158BE5.

Pollet:2020:PSP

[989] Charlotte-V. Pollet. The procedure of the Section of Pieces of Areas in Li Ye and Yang Hui’s works: genealogy of diagrams and equations. Science in Context, 33(1):37–63, March 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
procedure-of-the-section-of-pieces-of-areas-in-li-ye-and-yang-
huis-works-genealogy-of-diagrams-and-equations/E4EEF062F0798A058451BA669312DDE9.

Chen:2020:PRP

[990] Jiang-Ping Jeff Chen. Practices of reasoning: persuasion and refutation in a seventeenth-century Chinese mathematical treatise of “linear algebra”. Science in Context, 33(1):65–93, March 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
practices-of-reasoning-persuasion-and-refutation-in-a-seventeenthcentury-
chinese-mathematical-treatise-of-linear-algebra/86453DFB85F88F88598143DF6E5A2294.

Anonymous:2020:SVIa

[991] Anonymous. SIC volume 33 issue 1 cover and front matter. Science in Context, 33(1):f1–f3, March 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-1-cover-and-front-matter/CF9189F49EBE33C7FAC14C360C6BEF73.

Anonymous:2020:SVIb

[992] Anonymous. SIC volume 33 issue 1 cover and back matter. Science in Context, 33(1):b1–b2, March 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-1-cover-and-back-matter/BFD401BAA9B363BCD8A22AE23BEB5732.

Boncourt:2020:PPD

[993] Thibaud Boncourt and Paulo Ravecca. Power, politics, and the development of political science in the Americas. Science in Context, 33 (2):95–100, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/power-politics-and-the-
development-of-political-science-in-the-americas/0084C703803354B44273A6CB7D0FEDAC.

Hauptmann:2020:WTS

[994] Emily Hauptmann. Why they shared: recovering early arguments for sharing social scientific data. Science in Context, 33(2):101–119, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/why-they-shared-recovering-early-arguments-
for-sharing-social-scientific-data/5EEDC91BAD33452832837181D7582613.

See erratum [1028].

Zerndt:2020:HCS

[995] Emily Zerndt. Historicizing the comparative survey of freedom: tracing the social trajectory of an influential indicator. Science in Context, 33(2):121–144, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/historicizing-the-
comparative-survey-of-freedom-tracing-the-social-trajectory-
of-an-influential-indicator/D40ECE92677A77F2F0329FCF3DAB9D51.

Rodrigues:2020:BPS

[996] Lidiane Soares Rodrigues. Brazilian political scientists and the Cold War: Soviet hearts, North-American minds (1966–1988). Science in Context, 33(2):145–169, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
brazilian-political-scientists-and-the-cold-war-soviet-hearts-
northamerican-minds-19661988/7899A75D56FB372C06763E51E10DF5DB.

Ravecca:2020:DTF

[997] Paulo Ravecca. Dictatorship, transition, and the forging of political science in Uruguay. Science in Context, 33(2):171–193, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/dictatorship-transition-and-the-forging-
of-political-science-in-uruguay/A0DE99309EE9BB14E3EEC5EB59F157CD.

Anonymous:2020:SVIc

[998] Anonymous. SIC volume 33 issue 2 cover and front matter. Science in Context, 33(2):f1–f3, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-2-cover-and-front-matter/E9CE3C9881E827BDE7B7F5DFCA67149B.

Anonymous:2020:SVId

[999] Anonymous. SIC volume 33 issue 2 cover and back matter. Science in Context, 33(2):b1–b2, June 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-2-cover-and-back-matter/353E00702E96AF77DCD8720AA79977BC.

Fraser:2020:RIC

[1000] Jennifer Fraser. Rendering Inuit cancer “visible”: Geography, pathology, and nosology in Arctic cancer research. Science in Context, 33(3): 195–225, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/rendering-inuit-cancer-
visible-geography-pathology-and-nosology-in-arctic-cancer-research/
0DD98B57486CD5CB60ECB28B93A4872D.

tenHagen:2020:LVG

[1001] Sjang L. ten Hagen. The Local versus the Global in the history of relativity: The case of Belgium. Science in Context, 33(3):227–250, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/local-versus-the-global-in-the-history-of-
relativity-the-case-of-belgium/FE3225A6006BBBADECDAE7519CB451C7.

Rossini:2020:CCM

[1002] Paolo Rossini. Changing conceptions of mathematics and infinity in giordano Bruno’s vernacular and Latin works. Science in Context, 33(3):251–271, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
changing-conceptions-of-mathematics-and-infinity-in-giordano-
brunos-vernacular-and-latin-works/98590F74983CED3EEBB33E3F506B5D1D.

Plutniak:2020:EPP

[1003] Sébastien Plutniak. The effects of publishing processes on scientific thought: Typography and typology in prehistoric archaeology (1950s–1990s). Science in Context, 33(3):273–297, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/effects-of-publishing-processes-
on-scientific-thought-typography-and-typology-in-prehistoric-
archaeology-1950s1990s/72D1ADC25C9511CC4A333DEFE68BDD2E.

Lloyd:2020:TTB

[1004] Stephanie Lloyd and Alexandre Larivée. Time, trauma, and the brain: How suicide came to have no significant precipitating event. Science in Context, 33(3):299–327, September 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/time-trauma-and-the-brain-how-suicide-came-to-have-
no-significant-precipitating-event/8AFC1CB5697B36D6FFEA6CEA8ADB41FB.

Henry:2020:NSG

[1005] John Henry. Newton, the sensorium of God, and the cause of gravity. Science in Context, 33(3):329–351, September 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
newton-the-sensorium-of-god-and-the-cause-of-gravity/B291C61C50623FEC5F4100CA69C0CBD2.

Anonymous:2020:SVIe

[1006] Anonymous. SIC volume 33 issue 3 cover and front matter. Science in Context, 33(3):f1–f3, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-3-cover-and-front-matter/0E11A8D7F54C820DFD3F143006038BFC.

Anonymous:2020:SVIf

[1007] Anonymous. SIC volume 33 issue 3 cover and back matter. Science in Context, 33(3):b1–b2, September 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-3-cover-and-back-matter/652D08FEF0278AE725067DBB030C0879.

Anonymous:2020:MMG

[1008] Anonymous. In memory of Miriam Greenfield. Science in Context, 33(4): 353, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/in-memory-of-miriam-greenfield/10412C7A333AF34297EDF3ADB247454F.

Podgorny:2020:PCB

[1009] Irina Podgorny and Daniel Gethmann. “Please, come in.” Being a charlatan, or the question of trustworthy knowledge. Science in Context, 33(4): 355–361, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/please-come-in-being-a-
charlatan-or-the-question-of-trustworthy-knowledge/59679B7EF1AD1ED553E5B3D4F90F0BCF.

Vermeir:2020:CEI

[1010] Koen Vermeir. Charlatan epistemology: As illustrated by a study of wonder-working in the late seventeenth-century Dutch Republic.

Science in Context, 33(4):363–384, December 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
charlatan-epistemology-as-illustrated-by-a-study-of-wonderworking-
in-the-late-seventeenthcentury-dutch-republic/8B64A35DBB9461FC5AC9BE5BCC87AFB2.

Guillemain:2020:DPH

[1011] Hervé Guillemain. A democratic program for healing: The Raspail domestic medicine method in 1840s France. Science in Context, 33(4): 385–403, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/democratic-program-for-
healing-the-raspail-domestic-medicine-method-in-1840s-france/
DA53EDF8A766682D2E61694FB04DC386.

Nieto-Galan:2020:UCG

[1012] Agustí Nieto-Galan. Useful charlatans: Giovanni Succi and Stefano Merlatti’s fasting contest in Paris, 1886. Science in Context, 33 (4):405–422, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/useful-charlatans-
giovanni-succi-and-stefano-merlattis-fasting-contest-in-paris-
1886/BF11A690B9F79591266854E41575A6BA.

Sevilla:2020:CCC

[1013] Elisa Sevilla and Ana Sevilla. The “Controversial Cundurango Cure”: Medical professionalization and the global circulation of drugs. Science in Context, 33(4):423–440, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/controversial-
cundurango-cure-medical-professionalization-and-the-global-
circulation-of-drugs/336127CD5229059B06C21B725AD540D4.

deCeglia:2020:BER

[1014] Francesco Paolo de Ceglia and Lorenzo Leporiere. Becoming Eusapia: The rise of the “Diva of Scientists”. Science in Context, 33(4):441–471, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/becoming-eusapia-the-rise-of-the-
diva-of-scientists/D73C08D019A4FFF6729FB5A1B3479013.

Gethmann:2020:LCT

[1015] Daniel Gethmann. Levels of communication: The talking horse experiments. Science in Context, 33(4):473–490, December 2020. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
levels-of-communication-the-talking-horse-experiments/0E0846F5FAC4054F60D4EAA54590A316.

Richard:2020:BLP

[1016] Nathalie Richard. Between learned and popular culture: a world of syncretism and acculturation. Science in Context, 33(4):491–495, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/between-learned-and-popular-culture-a-world-
of-syncretism-and-acculturation/96EFB4CD0960DFCEC7639B143AB3FAA9.

Anonymous:2020:SVIg

[1017] Anonymous. SIC volume 33 issue 4 cover and front matter. Science in Context, 33(4):f1–f3, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-4-cover-and-front-matter/98FDFB5BA8FEE0818731B6B1C3817581.

Anonymous:2020:SVIh

[1018] Anonymous. SIC volume 33 issue 4 cover and back matter. Science in Context, 33(4):b1–b2, December 2020. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-33-
issue-4-cover-and-back-matter/CFD004AD2EA52B35A9D289715724534C.

Kochan:2021:ANT

[1019] Jeff Kochan. Animism and natural teleology from Avicenna to Boyle. Science in Context, 34(1):1–23, March 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
animism-and-natural-teleology-from-avicenna-to-boyle/1066C279A854DFD688104EBFBA71EA25.

Ducheyne:2021:UOT

[1020] Steffen Ducheyne. Using one’s talents in honor of God: Lambert ten Kate (1674–1731) and Isaac Newton’s natural philosophy. Science in Context, 34(1):25–50, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
using-ones-talents-in-honor-of-god-lambert-ten-kate-16741731-
and-isaac-newtons-natural-philosophy/8DD64733666B3D587AB86278F6BB4261.

Bar-On:2021:TNP

[1021] Kati Kish Bar-On. Towards a new philosophical perspective on Hermann Weyl’s turn to intuitionism. Science in Context, 34(1):51–68, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/towards-a-new-philosophical-perspective-
on-hermann-weyls-turn-to-intuitionism/90D2CA6E090FA8A100CD5B1A6BAC06CE.

Lipphardt:2021:RVP

[1022] Veronika Lipphardt, Gudrun A. Rappold, and Mihai Surdu. Representing vulnerable populations in genetic studies: The case of the Roma. Science in Context, 34(1):69–100, March 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
representing-vulnerable-populations-in-genetic-studies-the-
case-of-the-roma/0B4C1709436BE9891F9F12A214E93601.

Koolage:2021:IAS

[1023] W. John Koolage, Lauren M. Williams, and Morgen L. Barroso. An infrastructural account of scientific objectivity for legal contexts and bloodstain pattern analysis. Science in Context, 34(1):101–119, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/an-infrastructural-account-of-scientific-
objectivity-for-legal-contexts-and-bloodstain-pattern-analysis/
558850A5B9770772F9717C3228883C3D.

Winkler:2021:ICP

[1024] Sean Winkler. Individuals and collectives in the philosophy of Boris Hessen: an introduction. Science in Context, 34(1):121–136, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/individuals-and-collectives-in-the-philosophy-
of-boris-hessen-an-introduction/2A720B1D74E12FC8A2D465646D0D2771.

Hessen:2021:MST

[1025] Boris M. Hessen. Marian Smoluchowski (on the tenth anniversary of his death). Science in Context, 34(1):137–141, March 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
marian-smoluchowski-on-the-tenth-anniversary-of-his-death/41070AD541CEEC6C663A22355E680D3A.

Hessen:2021:CTA

[1026] Boris M. Hessen and Vasiliy P. Egorshin. On Comrade Timiryazev’s attitude towards contemporary science. Science in Context, 34(1):143– 153, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/on-comrade-timiryazevs-attitude-
towards-contemporary-science/FC00EC1ED31891D5CF9D53F324AF53CE.

Hessen:2021:MMM

[1027] Boris M. Hessen. Mechanical materialism and modern physics. Science in Context, 34(1):155–186, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
mechanical-materialism-and-modern-physics/23CB5622A8842091310B84D797DA1668.

Hauptmann:2021:WTS

[1028] Emily Hauptmann. Why they shared: recovering early arguments for sharing social scientific data — erratum. Science in Context, 34(1): 187, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/why-they-shared-recovering-early-
arguments-for-sharing-social-scientific-data-erratum/9F495C56D8FE96B336646ECBDAA8AE19.

See [994].

Anonymous:2021:SVIa

[1029] Anonymous. SIC volume 34 issue 1 cover and front matter. Science in Context, 34(1):f1–f3, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-1-cover-and-front-matter/2CAF8D2647233A1764EC349B28B66382.

Anonymous:2021:SVIb

[1030] Anonymous. SIC volume 34 issue 1 cover and back matter. Science in Context, 34(1):b1–b2, March 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-1-cover-and-back-matter/5707E0B107A0669910C1698989AF5EAB.

Navarro:2021:HSO

[1031] Jaume Navarro. How scientific objects end. Science in Context, 34 (2):189–193, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/how-scientific-objects-
end/59AB470CF538FCEE3371020475B5DECD.

Arabatzis:2021:DSO

[1032] Theodore Arabatzis. Do scientific objects have a life (which may end)? Science in Context, 34(2):195–208, June 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
do-scientific-objects-have-a-life-which-may-end/5F27A50206F2A1A7A58F05C4ABC33E31.

Navarro:2021:KOO

[1033] Jaume Navarro. Killed by its own obituaries: Explaining the demise of the ether. Science in Context, 34(2):209–225, June 2021. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/killed-by-its-own-obituaries-explaining-the-demise-
of-the-ether/A4FF25A9BB5D0D8221277068A48A193A.

Chang:2021:DUC

[1034] Hasok Chang. Dead or “undead”? The curious and untidy history of Volta’s concept of “contact potential”. Science in Context, 34(2): 227–247, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/dead-or-undead-the-curious-and-untidy-
history-of-voltas-concept-of-contact-potential/200ED0A6FFC0430711B9BC5E4C08CD53.

Belteki:2021:ELH

[1035] Daniel Belteki. At the ends of the line: How the Airy Transit Circle was gradually overshadowed by the Greenwich Prime Meridian. Science in Context, 34(2):249–264, June 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
at-the-ends-of-the-line-how-the-airy-transit-circle-was-gradually-
overshadowed-by-the-greenwich-prime-meridian/DFF4D0405D8EE2A0FA0BA8050F4E899B.

Ion:2021:ACM

[1036] Alexandra Ion. Anatomy collections as “modern ruins”: The nostalgia of lonely specimens. Science in Context, 34(2):265–279, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/anatomy-collections-as-modern-ruins-the-
nostalgia-of-lonely-specimens/EC9B6A94320F7A3A535806EE233EAD5B.

Guerra:2021:TPC

[1037] Corinna Guerra. A terrifying poison or a cheap fertilizer? The life and death of Mount Vesuvius ash. Science in Context, 34 (2):281–296, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/terrifying-poison-
or-a-cheap-fertilizer-the-life-and-death-of-mount-vesuvius-
ash/FB6CBB7EC527FBE85F9A776C23ADE788.

Paskins:2021:DDD

[1038] Mat Paskins. Dyeing off: On the deaths of dyestuffs as scientific objects. Science in Context, 34(2):297–311, June 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
dyeing-off-on-the-deaths-of-dyestuffs-as-scientific-objects/
B8900B3CBB7C019C25A6A94CD374F7FA.

Anonymous:2021:SVIc

[1039] Anonymous. SIC volume 34 issue 2 cover and front matter. Science in Context, 34(2):f1–f3, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-2-cover-and-front-matter/5270A91D165BF17A3D991171F2661A5A.

Anonymous:2021:SVId

[1040] Anonymous. SIC volume 34 issue 2 cover and back matter. Science in Context, 34(2):b1–b2, June 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-2-cover-and-back-matter/CD2D7785E333741E60AF92B4A643FD71.

Mazliak:2021:BPN

[1041] Laurent Mazliak. Belgium and probability in the nineteenth century: The case of Paul Mansion. Science in Context, 34(3):313–340, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/belgium-and-probability-in-the-nineteenth-
century-the-case-of-paul-mansion/DE277BA28C7BA1B76B28B311A3A8644A.

Szamosi:2021:MDI

[1042] Barna Szamosi. Medical decisions influenced by eugenics: Hungarian gynecological practices during the 1910s. Science in Context, 34(3):341– 355, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/medical-decisions-influenced-by-
eugenics-hungarian-gynecological-practices-during-the-1910s/
A05F41EFF5D0260C6FE601BA6B9C816A.

Anta:2021:IMP

[1043] Javier Anta. Information, meaning and physics: The intellectual evolution of the English School of Information Theory during 1946–1956. Science in Context, 34(3):357–373, September 2021. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://

www.cambridge.org/core/journals/science-in-context/article/
information-meaning-and-physics-the-intellectual-evolution-
of-the-english-school-of-information-theory-during-19461956/
0ADDEE8EFD570F9AFEFFC6F84DAF5731.

Naylor:2021:BSF

[1044] Robert L. Naylor. The Bryson synthesis: The forging of climate change narratives during the World Food Crisis. Science in Context, 34(3):375– 391, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/bryson-synthesis-the-forging-of-
climate-change-narratives-during-the-world-food-crisis/B8AEFC18E35AB1D294915A12E6E90678.

Anzola:2021:TPG

[1045] David Anzola. The theory–practice gap in the evaluation of agentbased social simulations. Science in Context, 34(3):393–410, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (elec-

tronic). URL https://www.cambridge.org/core/journals/science-
in-context/article/theorypractice-gap-in-the-evaluation-of-
agentbased-social-simulations/F538908F0C0AA62D67281AE15886A698.

Anonymous:2021:SVIe

[1046] Anonymous. SIC volume 34 issue 3 cover and front matter. Science in Context, 34(3):f1–f3, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-3-cover-and-front-matter/7654275EC042F2786FCC7BD783527DCE.

Anonymous:2021:SVIf

[1047] Anonymous. SIC volume 34 issue 3 cover and back matter. Science in Context, 34(3):b1–b2, September 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-3-cover-and-back-matter/775E7E624BB644B7226073458880921F.

Schlicht:2021:DD

[1048] Laurens Schlicht, Sophie Ledebur, and Anna Echterhölter. Data at the doorstep. Science in Context, 34(4):411–421, December 2021. CO-DEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL

https://www.cambridge.org/core/journals/science-in-context/
article/data-at-the-doorstep/C0256CEFB2258EFBECC3A948CFF0B84B.

teHeesen:2021:IIP

[1049] Anke te Heesen. Interview and interior: Procedures of narrative surveys around 1900. Science in Context, 34(4):423–438, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/interview-and-interior-
procedures-of-narrative-surveys-around-1900/4FF1271F228DF2A12C358F220F30532E.

vonOertzen:2021:TFM

[1050] Christine von Oertzen. True to form: Media and data technologies of self-inscription. Science in Context, 34(4):439–458, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/true-to-form-media-and-
data-technologies-of-selfinscription/61C1F69BA159B18E0567DECD4D236890.

Ledebur:2021:EUC

[1051] Sophie Ledebur. Evidence of undercounting: Collecting data on mental illness in Germany (c. 1825–1925). Science in Context, 34(4):459–478, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/evidence-of-undercounting-collecting-
data-on-mental-illness-in-germany-c-18251925/DD9953B3EB0454C53A2D704ADAAA1C81.

Schlicht:2021:WDC

[1052] Laurens Schlicht. The world of deviance in the classroom: Psychological experiments on schoolchildren in Weimar Germany. Science in Context, 34(4):479–499, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.org/

core/journals/science-in-context/article/world-of-deviance-
in-the-classroom-psychological-experiments-on-schoolchildren-
in-weimar-germany/5FA244559666A12AC6E828193A416F4F.

Soudan:2021:CDS

[1053] Gabrielle Soudan, David Philippy, and Harro Maas. Crossing the doorsteps for social reform: The social crusades of Florence Kelley and Ellen Richards. Science in Context, 34(4):501–525, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print),

1474-0664 (electronic). URL https://www.cambridge.org/core/
journals/science-in-context/article/crossing-the-doorsteps-
for-social-reform-the-social-crusades-of-florence-kelley-and-
ellen-richards/414DEA65D69E20AE6253E6461A7ACC93.

Echterholter:2021:FEC

[1054] Anna Echterhölter. Formative encounters: Colonial data collection on land and law in German Micronesia. Science in Context, 34(4):527–552, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-

0664 (electronic). URL https://www.cambridge.org/core/journals/
science-in-context/article/formative-encounters-colonial-data-
collection-on-land-and-law-in-german-micronesia/32F34AEC186EDF08DC385A29FEE9F2A8.

Anonymous:2021:SVIg

[1055] Anonymous. SIC volume 34 issue 4 cover and front matter. Science in Context, 34(4):f1–f3, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-4-cover-and-front-matter/7279F1CE77F621D748FCA203BE31DA3E.

Anonymous:2021:SVIh

[1056] Anonymous. SIC volume 34 issue 4 cover and back matter. Science in Context, 34(4):b1–b2, December 2021. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic). URL https://www.cambridge.

org/core/journals/science-in-context/article/sic-volume-34-
issue-4-cover-and-back-matter/9A6F5DD2B04468E5CEF862D4449BF47F.

Hammer:1958:PST

[1057] Franz Hammer. Nicht Pascal sondern der Tübinger Professor Wilhelm Schickard erfand die Rechenmaschine!. (German) [Not Pascal, but the Tübingen professor William Schickard, invented the calculator!]. Büromarkt, 20(??):1023–1025, ???? 1958. ISSN 0007-3148.

Seck:1978:WSA

[1058] Friedrich Seck, editor. Wilhelm Schickard, 1592–1635: Astronom, Geograph, Orientalist, Erfinder der Rechenmaschine. (German) [Wilhelm Schickard, 1592–1635: Astronomer, Geographer, Orientalist, Inventor of the calculator], volume 25 of Contubernium. Mohr, Tübingen, West Germany, 1978. ISBN 3-16-939772-9. 422 pp. LCCN QB36.S312 W54.

Kistermann:1995:RWS

[1059] Friedrich Wilhelm Kistermann. Die Rechentechnik um 1600 und Wilhelm Schickards Rechenmaschine. (German) [The calculating technique of 1600 and Wilhelm Schickard’s calculator]. In Seck [1063], pages 241– 272. ISBN 3-7995-3235-8. ISSN 0340-6857. LCCN ???? DM 76.00, sfr 76.00, S 600.00.

Lehmann:1995:SLE

[1060] N. Joachim Lehmann. Schickard und Leibniz als Erfinder von rechenmaschinen. (German) [Schickard and Leibniz, the inventors of calculators]. In Seck [1063], pages 273–286. ISBN 3-7995-3235-8. ISSN 0340-6857. LCCN ???? DM 76.00, sfr 76.00, S 600.00.

Kistermann:2001:CHU

[1061] Friedrich W. Kistermann. Calculators: How to use the Schickard calculator. IEEE Annals of the History of Computing, 23(1):80–85, January/March 2001. CODEN IAHCEX. ISSN 1058-6180 (print),

1934-1547 (electronic). URL http://dlib.computer.org/an/books/
an2001/pdf/a1080.pdf.

Seck:1981:WWS

[1062] Friedrich Seck, editor. Wissenschaftsgeschichte um Wilhelm Schickard: Vorträge bei dem Symposion der Universität Tübingen im 500. Jahr ihres Bestehens am 24.und 25. Juni 1977. (German) [History of Science and William Schickard Presentations at the Symposium of the University of Tübingen in 500 Years of existence on 24th and 25th June 1977], volume 26 of Contubernium. J. C. B. Mohr, Tübingen, West Germany, 1981. ISBN 3-16-444151-7. LCCN QB29.

Seck:1995:GWS

[1063] Friedrich Seck, editor. Zum 400. Geburtstag von Wilhelm Schickard: Zweites Tübinger Schickard-Symposion, 25. bis 27. Juni 1992. (German) [On the 400th Birthday of Wilhelm Schickard: Second Tubingen Schickard Symposium, 25–27 June 1992], volume 41 of Contubernium. Thorbecke, Sigmaringen, Germany, 1995. ISBN 3-7995-3235-8. ISSN 0340-6857. LCCN ???? DM 76.00, sfr 76.00, S 600.00.